### v1

# UnsafeBench: Benchmarking Image Safety Classifiers on Real-World and AI-Generated Images

Abstract
Image safety classifiers play an important role in identifying and mitigating the spread of unsafe images online (e.g., images including violence, hateful rhetoric, etc.). At the same time, with the advent of text-to-image models and increasing concerns about the safety of AI models, developers are increasingly relying on image safety classifiers to safeguard their models. Yet, the performance of current image safety classifiers remains unknown for real-world and AI-generated images. To bridge this research gap, in this work, we propose UnsafeBench, a benchmarking framework that evaluates the effectiveness and robustness of image safety classifiers. First, we curate a large dataset of 10K real-world and AI-generated images that are annotated as safe or unsafe based on a set of 11 unsafe categories of images (sexual, violent, hateful, etc.). Then, we evaluate the effectiveness and robustness of five popular image safety classifiers, as well as three classifiers that are powered by general-purpose visual language models. Our assessment indicates that existing image safety classifiers are not comprehensive and effective enough in mitigating the multifaceted problem of unsafe images. Also, we find that classifiers trained only on real-world images tend to have degraded performance when applied to AI-generated images. Motivated by these findings, we design and implement a comprehensive image moderation tool called PerspectiveVision, which effectively identifies 11 categories of real-world and AI-generated unsafe images. The best PerspectiveVision model achieves an overall F1-Score of 0.810 on six evaluation datasets, which is comparable with closed-source and expensive state-of-the-art models like GPT-4V. UnsafeBench and PerspectiveVision can aid the research community in better understanding the landscape of image safety classification in the era of generative AI. [^1] Our project website is at https://unsafebench.github.io.
Disclaimer. This paper contains disturbing and unsafe images. We only blur/censor Not-Safe-for-Work (NSFW) imagery. Nevertheless, reader discretion is advised.
Introduction
Unsafe images that include inappropriate content, such as violence, self-harm, and hate, are prevalent across Web communities like Reddit [20] and 4chan [1]. Their presence poses significant challenges to communities and society at large; they can reinforce stereotypes [29, 37, 70], incite hate and violence [46, 39, 55], and trigger self-harm behaviors [54]. To combat this longstanding problem, online platforms rely on image safety classifiers and human moderators to identify and remove unsafe images from the Web. Image safety classifiers like Q16 [63] and Not-Suitable-For-Work (NSFW) detector [16] are trained on real-world unsafe images and are used for detecting real-world unsafe images that are shared on the Web. For instance, LAION-AI [9], a data provider for machine learning models and datasets scrapped from the Web, applies these two image safety classifiers to report unsafe real-world images in their datasets, like LAION-400M [66] and LAION-5B [65]. Overall, both the Web and publicly available datasets include a non-negligible amount of real-world unsafe images that need to be moderated.
Parallel to real-world unsafe images, recently emerged text-to-image models like Stable Diffusion [60] and Midjourney [12] are also capable of producing unsafe images. According to a recent study [56], if intentionally misled by malicious users, text-to-image models have a 16%-51% probability of generating unsafe content like sexual, disturbing, violent content, etc. A notable real-world incident in January 2024 is the viral spread of AI-generated sexually explicit images of Taylor Swift [3], which further highlights the risks of these text-to-image models. To mitigate these risks, AI practitioners, again, rely on image safety classifiers to identify unsafe images during the generation process and block generated unsafe images before presenting them to end-users. For example, Stable Diffusion implements a safety filter [21] that defines 20 sensitive concepts, such as “nude,” “sex,” and “porn,” to prevent the generation of unsafe images. Additionally, extensive research [62, 38, 76, 31] applies Q16 [63] and NudeNet [17] to evaluate the safety of text-to-image models and propose new safety mechanisms.
Despite the wide use of these image safety classifiers, their performance lacks a thorough evaluation across both real-world and AI-generated unsafe images. It remains unclear if they are truly capable of detecting a broad range of unsafe content and whether their performance can be generalized to AI-generated images, given that most image safety classifiers are trained on real-world image datasets. Furthermore, the recent development of large visual language models (VLMs), such as LLaVA [48] and GPT-4V [5], provides new possibilities in this field. These models are general-purpose visual language models that have undergone rigorous safety alignment, which empowers them to understand unsafe content portrayed in images. Given the current state, it is still unknown if these advanced VLMs will supersede existing image safety classifiers with better performance.
[…]
Based on the insights drawn from UnsafeBench, we introduce PerspectiveVision, a set of models that provide fine-grained classification of unsafe images based on the 11 categories of unsafe images. We train and evaluate PerspectiveVision on our annotated dataset and further assess its generalizability across six external datasets containing both real-world and AI-generated images.
Main Findings. Our study yields the following main findings:
  • Overall, GPT-4V is the top-performing model in identifying unsafe content, with an F1-Score between 0.423 and 0.847 (depending on the category of unsafe content). However, due to its closed-source and commercial nature, widely applying GPT-4V on image safety classification tasks is impractical. Currently, there is a lack of an open-source classifier that can comprehensively and effectively identify the wide spectrum of unsafe images.
  • There is an imbalance in moderation across different types of unsafe images. The Sexual and Shocking categories are more effectively detected, with an average F1-Score close to 0.8. However, the categories of Hate, Harassment, and Self-Harm are not effectively detected, with an average F1-Score below 0.6. For example, certain hateful symbols, such as Neo-Nazi symbols in tattoos, evade detection by GPT-4V.
  • Conventional classifiers trained only on real-world images, e.g., NSFW_Detector and NudeNet, experience performance degradation on AI-generated images, with F1-Scores decreasing from 0.650-0.698 to 0.596-0.603. We find that AI-generated images often have unique characteristics, such as artistic representations of unsafe content and the composition of unsafe images in a grid layout. Through case studies, we verify that these characteristics can indeed affect the classifier’s predictions, thereby contributing to performance degradation on AI-generated images.
  • Leveraging large pre-trained foundation models such as VLMs and CLIP presents higher robustness compared to training a small classifier from scratch. Under perturbed images, VLM-based classifiers present the highest Robust Accuracy (RA) above 0.608. Among conventional classifiers, those that utilize CLIP as an image feature extractor, such as Q16, MultiHeaded, SD_Filter, and NSFW_Detector, demonstrate higher robustness compared to those trained from scratch, e.g., NudeNet. Minimal perturbation is sufficient to deceive NudeNet, with an average RA of only 0.293.
  • Our evaluation shows that PerspectiveVision achieves comparable performance to GPT-4V in identifying unsafe images. The best F1-Score is 0.859 on the UnsafeBench test set and 0.810 on six external evaluation datasets, which include morally negative, nudity, hateful, self-harm, and violent images from either real-world or AI-generated sources. As the top-performing baseline, GPT-4V achieves an F1-Score of 0.818 on these external datasets.
Contributions. Our work makes three important contributions. First, we contribute to the research community with a comprehensive image dataset for image safety research. The dataset consists of 10K real-world and AI-generated images, covering unsafe content from 11 categories, with each image being meticulously annotated by three authors. It serves as a foundation dataset for future research in this field. Second, we thoroughly assess the effectiveness and robustness of current image safety classifiers and compare them with newly emerged VLMs. Our analysis highlights the varying performance of these classifiers across different unsafe categories, aiding these models to learn the vulnerabilities in identifying specific unsafe content, such as tattooed Neo-Nazi symbols. Furthermore, our analysis shows that AI-generated unsafe images pose challenges to the existing classifiers, especially those trained on real-world images, largely due to the distinct characteristics of AI-generated images. These findings highlight the importance for platform moderators and model developers to include AI-generated images in their training data to enhance models’ generalizability. Finally, inspired by Google’s Perspective API in the NLP domain, we contribute to the community with PerspectiveVision, which identifies a wide range of unsafe images with comparable performances to GPT-4V. PerspectiveVision is made available as an open-source tool, establishing a baseline for detecting general unsafe images. To aid future research, we are committed to responsibly sharing and continuously updating our annotated dataset and PerspectiveVision models. We call for joint and ongoing efforts from the research community to build better content moderation tools.
Ethical Considerations. We curate our dataset based on public databases from LAION-5B and the Lexica website. This process does not involve user deanonymization, therefore, our work is not considered human subjects research by our Institutional Review Boards (IRB). During image annotation, we rely on three authors instead of crowdsourcing workers to avoid exposing unsafe content to third parties. We have also given careful consideration to sharing our dataset responsibly. For example, to prevent misuse, we will make the dataset available on Hugging Face, but only upon request and for research purposes.
Background
Unsafe Image Taxonomy
[…]
Existing Image Safety Classifiers
We introduce five popular image safety classifiers: Q16 [63], MultiHeaded [56], SD_Filter [58], NSFW_Detector [16], and NudeNet [17]. Table 1 provides an overview of their backbone models, training paradigms, and training datasets.
Table 1: Overview of existing image safety classifiers. “D” denotes the dataset. “R” and “A” represent real-world and AI-generated images, respectively.
Classifier Backbone Training Paradigm Training Dataset
Q16 CLIP Prompt Learning SMID [32] (R)
MultiHeaded CLIP Linear Probing MultiHeaded-D [56] (A)
SD_Filter CLIP - -
NSFW_Detector CLIP Linear Probing LAION-400M Subset (R)
NudeNet Xception Fully Supervised NudeNet-D [18] (R)
Q16 [63]. Q16 is a binary image classifier that predicts a given image as morally positive or negative. [^3] https://github.com/ml-research/Q16. It first extracts the image embedding with the CLIP [57] image encoder and calculates the cosine similarity between the image embedding and two soft prompts (optimized text embeddings), one representing the positive class and the other being the negative class. Q16 then assigns the image to the more relevant class. To obtain the optimal soft prompts, the Q16 developer starts with two normal prompts, e.g., “This image is positive/negative,” and iteratively updates their text embeddings to maximize the cosine similarity values on a labeled image dataset, SMID [32]. The Socio-Moral Image Database (SMID) includes 2,941 morally positive and negative images, covering concepts including harm, inequality, degradation, deception, etc. Q16 has been widely used in evaluating the safety/appropriateness of images [63, 65, 62, 47, 76, 38, 31, 69, 30], including identifying unsafe images sourced from LAION-5B [65] and text-to-image models [62, 47, 76, 38, 51], as well as testing the reliability of safety mechanism built in text-to-image models [31, 69, 30].
MultiHeaded Safety Classifier (MultiHeaded) [56]. This classifier is also a CLIP-based model with five linear classification heads, detecting sexually explicit, violent, shocking, hateful, and political images. [^4] https://github.com/YitingQu/unsafe-diffusion. The classifier is trained on 800 labeled images generated by text-to-image models like Stable Diffusion [56]. It has been employed to estimate the cleanliness (percentage of unsafe images) of large public datasets [56] such as LAION-400M [66], LAION-2B [7], LAION-aesthetics [64], and YFCC15M [67].
Stable Diffusion Filter (SD_Filter) [58]. This filter is built in Stable Diffusion to prevent generating explicit images. [^5] https://huggingface.co/CompVis/stable-diffusion-safety-checker. It relies on the CLIP image encoder to assess image embeddings, then measures the distance between these embeddings and text embeddings of 20 sensitive concepts [26], such as “sexual,” “nude,” and “sex.” During the image generation process, the filter compares the embedding distance between the generated image with a set of pre-defined thresholds. If the image is too close to any sensitive concepts, the filter will predict it as unsafe and block it from users.
NSFW_Detector [16]. The NSFW_Detector is also a classification model with the CLIP image encoder as the backbone and a multilayer perceptron (MLP) as the classification head. [^6] https://github.com/LAION-AI/CLIP-based-NSFW-Detector. It mainly detects sexual and nudity images. LAION-AI [9] uses this model to assess the NSFW score for images in their large-scale datasets including LAION-400M [66] and LAION-5B [8].
NudeNet [17]. NudeNet is a lightweight nudity detection tool and consists of a detector that detects the sexually explicit areas in images and a classifier that directly classifies them as explicit images or not. In this study, we use the binary NudeNet classification model. [^7] https://pypi.org/project/nudenet/. It is a classifier with Xception as the backbone, trained on roughly 160K automatically-labeled images [18]. It has been extensively applied in detecting sexually explicit images  [62, 71, 51, 38, 76] and is often combined with Q16 for identifying general unsafe content [62, 38, 51, 76].
Large Visual Language Models (VLMs)
Large visual language models have achieved extraordinary capabilities in understanding visual and text content. Given an image and a text instruction, these models can read the image and generate responses following the instruction. Recent studies [41, 59] show that VLMs can be used to detect user-generated unsafe images [41] and hateful memes [59].
LLaVA [48]. LLaVA is an open-source visual language model that can answer questions based on the user’s provided image and prompt. It comprehends the image with the CLIP image encoder and understands the user’s prompts with a large language model, Vicuna [27]. LLaVA is trained on the LAION-CC-SBU dataset and instruction dataset generated by GPT4V [48]. We use the llava-v1.5-7b checkpoint. [^8] https://huggingface.co/liuhaotian/llava-v1.5-7b.
InstructBLIP [33]. InstructBLIP is also an open-source VLM. It is an enhanced version of the BLIP-2 model, extended with the capability to follow instructions related to visual content. InstructBLIP is trained on various datasets, including the same instruction dataset generated by GPT4V [33]. We adopt the instructblip-vicuna-7b checkpoint. [^9] https://huggingface.co/Salesforce/instructblip-vicuna-7b.
GPT-4V [5]. GPT-4V is a multimodal version of the GPT-4 architecture, specifically tailored for visual understanding and analysis. Building on the language processing capabilities of GPT-4, GPT-4V integrates enhanced image recognition and interpretation features. In this study, we use the gpt-4-vision-preview checkpoint.
Overview of UnsafeBench
We build UnsafeBench, a benchmarking framework that comprehensively evaluates the performance of image safety classifiers. We show an overview of UnsafeBench in Figure 1. Due to the absence of a comprehensive labeled dataset in the image safety domain, we first construct the UnsafeBench dataset, which contains 10K images with human annotations. We then collect the existing image safety classifiers, as conventional classifiers, and three VLMs capable of classifying unsafe images, as VLM-based classifiers. We identify the range of unsafe content covered by these classifiers and align them with our unsafe image taxonomy, i.e., 11 unsafe categories. Finally, we evaluate the effectiveness and robustness of these classifiers; effectiveness measures how accurately the classifiers can identify unsafe images, while robustness reflects their ability to maintain accuracy when faced with perturbed images.
Figure 1: Overview of UnsafeBench.
Dataset Construction
Image Collection. We rely on open large-scale multimodal datasets, LAION-5B [65] and Lexica [10], to collect real-world and AI-generated unsafe images. LAION-5B is currently the largest public image-text dataset in the world [8], containing 5.85 billion image-text pairs collected from webpages [65]. We regard LAION-5B as the source to collect real-world images. [^10] It is possible that webpages also contain AI-generated images. However, LAION-5B was collected in March, 2022 [8], while text-to-image generative models like Stable Diffusion and Midjourney gained popularity after July 2022 [25, 13]. Thus, we believe the majority of the images in LAION-5B are real-world images. Lexica [10] is an AI-generated image gallery that allows users to create and disclose images generated by models like Stable Diffusion [60]. Both datasets function as a search engine and enable users to find the most relevant images based on a textual description. Inspired by this, we use unsafe keywords to query these datasets and collect potentially unsafe images.
To collect unsafe keywords, we split the definitions of 11 unsafe categories (see Section 2) as initial unsafe keywords. Additionally, we employ a large language model, Vicuna-33b, [^11] https://huggingface.co/lmsys/vicuna-33b-v1.3. to augment the definition of each unsafe category and provide 10 specific examples. We show the generated examples in Table 8 in the Appendix. Combining the initial unsafe keywords and those generated with Vicuna-33b, we obtain 158 unsafe keywords, covering 11 unsafe categories (each containing 12-16 keywords). We then query LAION-5B and Lexica using these keywords and collect the most relevant images. Specifically, for each query, Lexica returns 50 images by default. LAION-5B, on the other hand, returns a flexible number of image URLs. However, some URLs are repetitive or inaccessible. To ensure a balanced number of images from both datasets, we set the number of returned images from LAION-5B to 100. After removing duplicates, we collect a total of 12,932 potentially unsafe images, comprising 5,815 images from LAION-5B and 7,117 from Lexica.
Image Annotation. We perform a human annotation to determine if these collected images are truly unsafe. Three authors of this paper serve as annotators and manually annotate them independently. We require these annotators to first read the definition from each unsafe category as the criterion in determining whether an image is safe, unsafe, or not applicable (N/A). Note that N/A mainly represents noise images that cannot be identified as safe or unsafe, e.g., a blurry image or one with unidentifiable texts. The annotation process undergoes two rounds. In the first round, two annotators independently assign a label of safe, unsafe, or N/A to each image. For images where two annotators disagree, we further introduce the third annotator to provide additional labels. The final label for each image is then determined based on the majority vote among these labels. In the second round, annotators revisit images annotated as unsafe and determine the unsafe category. If an image displays a mix of unsafe elements, we choose the predominant category it violates. Through two rounds of annotation, we ensure that each image is labeled as safe, unsafe, or N/A, and if it is unsafe, it can be classified into a specific unsafe category.
To evaluate the reliability of our annotation, we calculate the Fleiss’ Kappa score [36, 35], a statistical measure for assessing the agreement among annotators. The score is 0.47, indicating a moderate level of agreement among annotators [36, 35]. We understand such a Fleiss’ Kappa score is not perfect, as the annotators’ perceptions of unsafe images may still vary under the same definition of our unsafe image taxonomy. That explains why it is necessary to introduce the third annotator in the first round to improve the reliability of our annotation.
Table 2: Statistics of the annotated images.
Dataset # All # Safe # Unsafe # N/A Fleiss’ Kappa
LAION-5B 5,815 3,228 1,832 755 0.49
Lexica 7,117 2,870 2,216 2,031 0.46
All 12,932 6,098 4,048 2,786 0.47
Table 3: Aligning the unsafe content covered by image safety classifiers with 11 unsafe categories.
Classifier Hate Harassment Violence Self-Harm Sexual Shocking Illegal Activity Deception Political Health Spam
Conventional Classifiers Q16
MultiHeaded
SD_Filter
NSFW_Detector
NudeNet
VLM-Based Classifiers LLaVA
InstructBLIP
GPT-4V
Dataset Statistics. We show the statistics of our annotated images in Table 2. Overall, we annotate 12,932 images, including 6,098 safe images, 4,048 unsafe images, and 2,786 N/A images. For the following experiments, we exclude the N/A images and use the remaining images (safe + unsafe), collectively referred to as the UnsafeBench dataset.
To the best of our knowledge, the UnsafeBench dataset is currently the most comprehensive annotated dataset in the image safety domain. It consists of 10,146 labeled images, encompassing 11 unsafe categories and two sources (real-world and AI-generated). Each unsafe category contains 769-1,151 images. We display more statistical details in Table 9 and image examples in Figure 10 in the Appendix. In this dataset, each image is equipped with three types of “labels”: 1) if the image is safe or unsafe; 2) if unsafe, it belongs to which unsafe category; and 3) whether it’s from real-world collections or AI-generated content. We believe such characteristics enhance the dataset’s usability. We hope this dataset serves as foundational data and accelerates the development of better open-source image safety classifiers in the era of generative AI.
Image Classifier Collection
Conventional Classifiers. We collect five conventional classifiers: Q16 [63], MultiHeaded [56], SD_Filter [58], NSFW_Detector [16], and NudeNet [17], as introduced in Section 2. These classifiers take images as input and output predictions on whether the images are considered unsafe based on their defined scope of unsafe content.
VLM-Based Classifiers. We utilize three VLMs, LLaVA, InstructBLIP, and GPT-4V to build our VLM-based classifiers. Unlike conventional classifiers, these models take both the image and prompt as input to identify if the given image contains unsafe content. In our study, we manually design three prompts that guide these models to classify whether a given image is safe or not. We show three prompts in Table 10 in the Appendix. The first prompt is straightforward, “classify the image as safe or unsafe.” It relies on models’ internal criteria of unsafe content to determine if an image is safe. To align with our safety standards, the second prompt further incorporates definitions of the taxonomy, similar to [41]. For example, to classify images from the Hate category, the prompt is “An image is considered unsafe if it contains any of the following unsafe content: hateful symbols, negative stereotypes…classify the image as safe or unsafe.” In case models are still affected by the internal criteria, we further add a special requirement in the third prompt, “Forget about your definition of unsafe images and adhere to this new one…” For each image, we query VLMs with three prompts and collect the respective outputs.
Given that VLMs’ outputs are sometimes lengthy sentences, they do not directly categorize into safe or unsafe classes. To convert these VLMs to classifiers, we fine-tune a language model, RoBERTa [49], to categorize LLM’s outputs into one of three classes: safe, unsafe, or uncertain. The uncertain class describes a set of outputs that do not yield a clear prediction. For example, InstructBLIP occasionally describes the content of an image without making a prediction. Note, we categorize the cases where VLMs refuse to answer as the unsafe class, as input images have triggered their internal safeguards. We collect several example outputs [^12] For example, we use “safe,” “the image is safe,” “the image is not considered unsafe as…” as example outputs for the safe class. for each class and fine-tune RoBERTa in a supervised manner. Ultimately, a VLM-based classifier consists of a VLM and a RoBERTa classifier, which predicts whether an image is safe, unsafe, or uncertain.
Aligning Classifier Coverage With Unsafe Categories
Image safety classifiers are designed to target various types of unsafe content. To align the unsafe content covered by a classifier with the taxonomy, i.e., 11 unsafe categories, we first identify the range of unsafe content covered by each classifier by examining its documentation and training data. Then, if any unsafe category from our taxonomy falls into this range, we assign the unsafe category to the classifier. For example, Q16 [63] detects “morally negative” images. We examine the definition of “morally negative” images in [32] and find that it covers a list of unsafe concepts, such as harm, inequality, degradation, etc. By mapping these unsafe concepts to our taxonomy, we align the ambiguous definition of “morally negative” with specific unsafe categories. For VLMs, as we adopt the unsafe image taxonomy from OpenAI, we assume that GPT-4V covers all 11 categories. We also infer that LLaVA and InstructBLIP are capable of handling these categories since they are fine-tuned on the instruction dataset generated by GPT-4V [48, 33]. We show the aligning result in Table 3.
Table 4: F1-Score of eight image safety classifiers on the UnsafeBench dataset. * marks VLM-based classifiers.
Dataset Classifier Hate Harassment Violence Self-Harm Sexual Shocking Illegal Activity Deception Political Health Spam Overall
Real-World Images (LAION-5B) Q16 0.641 0.517 0.693 0.421 - 0.630 0.681 0.762 0.271 0.144 - -
MultiHeaded 0.320 - 0.247 0.000 0.692 0.644 - - 0.209 - - -
SD_Filter - - - - 0.833 - - - - - - -
NSFW_Detector - 0.517 - - 0.783 - - - - - - -
NudeNet - - - - 0.650 - - - - - - -
LLaVA* 0.227 0.167 0.413 0.451 0.674 0.714 0.498 0.376 0.116 0.333 0.059 0.426
InstructBLIP* 0.351 0.394 0.606 0.275 0.796 0.467 0.722 0.653 0.444 0.379 0.380 0.520
GPT-4V* 0.556 0.706 0.774 0.557 0.866 0.724 0.897 0.827 0.605 0.405 0.718 0.712
AI-Generated Images (Lexica) Q16 0.336 0.416 0.612 0.521 - 0.836 0.497 0.384 0.597 0.615 - -
MultiHeaded 0.225 - 0.533 - 0.815 0.780 - - 0.744 - - -
SD_Filter - - - - 0.727 - - - - - - -
NSFW_Detector - 0.259 - - 0.678 - - - - - - -
NudeNet - - - - 0.596 - - - - - - -
LLaVA* 0.169 0.224 0.632 0.663 0.866 0.826 0.534 0.179 0.340 0.665 0.045 0.634
InstructBLIP* 0.178 0.332 0.629 0.383 0.757 0.795 0.665 0.302 0.785 0.663 0.519 0.589
GPT-4V* 0.254 0.635 0.712 0.613 0.827 0.875 0.701 0.422 0.909 0.621 0.171 0.707
Overall Q16 0.533 0.475 0.648 0.483 - 0.784 0.571 0.652 0.482 0.498 - -
MultiHeaded 0.292 - 0.426 - 0.757 0.749 - - 0.600 - - -
SD_Filter - - - - 0.785 - - - - - - -
NSFW_Detector - 0.449 - - 0.738 - - - - - - -
NudeNet - - - - 0.624 - - - - - - -
LLaVA* 0.210 0.189 0.550 0.590 0.780 0.800 0.519 0.332 0.266 0.575 0.054 0.549
InstructBLIP* 0.270 0.368 0.615 0.333 0.777 0.697 0.687 0.506 0.660 0.545 0.490 0.559
GPT-4V* 0.423 0.681 0.738 0.590 0.847 0.839 0.780 0.673 0.780 0.492 0.537 0.709
Evaluation Methodology
Effectiveness. We evaluate the effectiveness of eight image safety classifiers (five conventional classifiers and three VLM-based classifiers). For conventional classifiers, we feed images from each unsafe category and obtain binary predictions (safe/unsafe). Regarding VLM-based classifiers, we input both images with the designed prompts to gather outputs, which are then classified into safe/unsafe/uncertain classes by the fine-tuned RoBERTa. Uncertain predictions account for about 0.1%. However, since conventional classifiers directly predict images as safe or unsafe, to ensure a fair comparison, we randomly assign a prediction of safe or unsafe to VLM’s uncertain predictions. Recall that we design three prompts for VLMs, resulting in three separate predictions for each image. The final prediction is determined by a majority vote.
Evaluation Metric for Effectiveness. With the predictions in place, we calculate the F1-Score to evaluate the effectiveness of image safety classifiers. We choose F1-Score as it accounts for both false positives and false negatives, providing a balanced measure of the classifier’s Precision and Recall.
Robustness. Robustness reflects the classifier’s ability to maintain its performance beyond training examples. We evaluate the robustness of classifiers using images perturbed with both random noise (Gaussian noise) and adversarial noise (i.e., adversarial examples) [52, 50]. Adversarial examples ($x_{adv}$) are original inputs ($x$) with optimized perturbations ($\Delta x$), which maximizes the loss ($L(\theta,x+\Delta x,y)$) and fools the model into making incorrect predictions. The p-norm of the perturbation vector is limited by $\epsilon$ to ensure the perturbation is imperceptible.
$\begin{gathered}x_{adv}:=\underset{\Delta x}{\text{argmax}}\,L(\theta,x+\Delta x,y),\\ \quad\left\|\Delta x\right\|_{p}<\epsilon.\end{gathered}$ (1)
We craft adversarial examples using three gradient-based adversarial algorithms, FGSM [40], PGD [50], and DeepFool [53], with details introduced in Section A.1 in the Appendix. These algorithms optimize perturbations using gradients of the classifiers’ loss with respect to the input image. For conventional classifiers, we directly optimize the perturbations using three types of adversarial algorithms. However, for VLM-based classifiers (VLM+RoBERTa), the standard optimization approach is not applicable because we lose continuous gradient flow after VLMs generate word tokens. To address this challenge, we approximate the gradients by instructing the VLMs to generate a target output ($y_{tar}$), which is contrary to the image label. For example, if the VLM initially classifies a safe image with the output “This image is safe,” we compute the gradients when the target output is “unsafe.” Different from conventional classifiers, where we maximize the loss between the classifier’s prediction and the original label ($y$), here, we minimize the loss of a VLM between its prediction and the defined target output ($y_{tar}$). We craft adversarial examples with FGSM and PGD for open-source VLM-based classifiers (LLaVA and InstructBLIP). Note, we exclude DeepFool as it requires the gradients of logits, which are not directly applicable in this context.
$\begin{gathered}x_{adv}:=\underset{\Delta x}{\text{argmin}}\,L(\theta,x+\Delta x,y_{tar}),\\ \quad\left\|\Delta x\right\|_{p}<\epsilon.\end{gathered}$ (2)
Evaluation Metric for Robustness. We calculate the Robust Accuracy (RA) to evaluate the robustness of image safety classifiers. RA is the percentage of perturbed images that have been correctly predicted by classifiers out of all perturbed images. Here, we generate perturbed images only for those that are correctly predicted by classifiers during the effectiveness evaluation. Furthermore, we calculate the Mean Square Error (MSE) between original images and adversarial examples to reveal the magnitude of perturbations required to mislead classifiers into making incorrect predictions.
Evaluation Results
Effectiveness
Table 4 shows the effectiveness of eight image classifiers on the UnsafeBench dataset. This table is consistent with the aligning result shown in Table 3. If a classifier cannot identify a specific unsafe category according to Table 3, we fill the corresponding cell with a “-.”
Classifier-Level. The top-performing image safety classifier is the commercial VLM-based classifier, GPT-4V. It achieves the highest overall F1-Score, surpassing LLaVA and InstructBLIP, as well as other conventional classifiers in most unsafe categories, except for Hate and Deception. GPT-4V shows exceptional effectiveness in detecting Sexual (0.847), Shocking (0.839), Political (0.780), Illegal Activity (0.780), and Violence content (0.738). Among conventional classifiers, Q16 stands out by identifying the broadest spectrum of unsafe content, with an F1-Score from 0.475 to 0.784. However, Q16 does not support the detection of Sexual content and requires improvements in Harassment, Self-Harm, and Political categories. In contrast, MultiHeaded, SD_Filter, NSFW_Detector, and NudeNet can detect sexual images with an F1-Score of 0.624-0.785.
Although GPT-4V shows superior performance on our annotated dataset, the wide application is constrained by its commercial nature, mainly due to the financial cost and slow inference speed. It costs approximately \$20 and 50 minutes to classify 1K images. Additionally, we observe that zero-shot VLMs’ performances significantly rely on prompt designing. For example, GPT-4V’s overall F1-Score achieves 0.68 and 0.71 when we classify images with the second and third prompts (shown in Table 10 in the Appendix). However, the score drops to 0.61 on the first prompt. To obtain reliable results from VLMs like GPT-4V, we typically query the VLM with the same image using multiple prompts and determine the label based on a majority vote. This approach further increases the cost, making GPT-4V impractical for application on large-scale image datasets.
Current research [62, 38, 51, 76] prefers smaller classifiers over larger VLMs due to their faster inference speed. A common practice is combining Q16 and a classifier that detects sexually explicit content like NudeNet to detect generally unsafe images [62, 38, 51, 76], e.g., regarding the image is unsafe if either Q16 or NudeNet classifies it as unsafe. According to Table 4, this strategy can indeed cover many unsafe categories with a relatively good performance. We calculate the overall F1-Score of Q16 combined with NudeNet across the unsafe categories they can cover, and the score is 0.665. This performance needs to be further enhanced.
Different Unsafe Categories. To understand which unsafe categories are more effectively detected, we count the number of classifiers capable of identifying each category and record the average F1-Score in Figure 2. We find that the Sexual and Shocking categories have higher average F1-Scores, close to 0.8, compared to other unsafe categories. They are also covered by more classifiers, with 5-7 classifiers capable of identifying such content. In contrast, the remaining unsafe categories, particularly Spam, Hate, Harassment, and Self-Harm, have lower average F1-Scores (below 0.5). This discrepancy highlights the imbalanced moderation across different types of unsafe content.
Figure 2: Average F1-Score and number of classifiers for each unsafe category.
Among 11 unsafe categories, we particularly focus on the Hate category, given that hateful images can proliferate across Web communities as memes and are often used in coordinated hate campaigns [75, 39]. For instance, the anti-Semitic symbol, Happy Merchant [6], widely spreads on platforms like 4chan and mainly targets the Jewish community. Due to the harmful nature, OpenAI ranks the Hate category as the highest priority in its content policy [19]. However, among the five classifiers that can detect hateful content, the highest F1-Score is 0.533, achieved by Q16. The second highest is 0.423, achieved by GPT-4V. To understand the underlying reason behind such low F1-Scores, we zoom into the misclassified examples from the Hate category for both Q16 and GPT-4V. We calculate the false positive and false negative rates for both classifiers. For Q16, the false positive rate is 0.166 and the false negative rate is 0.306. In comparison, for GPT-4V, the false positive rate is 0.425, and the false negative rate is 0.144. This result suggests that both classifiers have a considerable likelihood to misclassify hateful images: either failing to identify truly hateful images or incorrectly predicting safe images as hateful. For model providers and platform moderators, the false negative rate may be more important as it indicates the likelihood of hateful images evading the detection of models’ safeguards. In our manual review of common false negatives, we observe that 58 images containing hateful content are misclassified as safe by both Q16 and GPT-4V. Notably, some Neo-Nazi and anti-Semitic symbols, such as the tattooed swastika [15] and the Happy Merchant meme [6], manage to evade detection. We show these examples in Figure 12 in the Appendix.
Figure 3: Average F1-Score of classifiers on real-world and AI-generated images.
(a) False Negatives (Misclassify Unsafe as Safe)
(b) False Positives (Misclassify Safe as Unsafe)
Figure 4: Image clusters from the Sexual category that are misclassified by SD_Filter, NSFW_Detector, and NudeNet. We annotate each central image with its cluster ID and cluster size. We blur sexual images for censoring purposes.
Real-World Versus AI-Generated. Real-world images from LAION-5B and AI-generated ones from Lexica have different distributions (see 11(a) in the Appendix). To compare the performance of tested classifiers on two groups of images, we calculate the average F1-Score of each classifier on real-world and AI-generated images in Figure 3. We find that SD_Filter, NSFW_Detector, NudeNet, and GPT-4V have better performance in identifying real-world unsafe images compared to AI-generated ones. For example, the F1-Score of SD_Filter decreases from 0.833 on real-world images to 0.727 on AI-generated images. Meanwhile, MultiHeaded, LLaVA, and InstructBLIP have higher F1-Scores on AI-generated images than real-world ones. The different behaviors of classifiers on two sets of images are multifaceted. A key factor is the distribution shift between training and testing images, including nuanced differences in semantics, image styles, etc. After examining the training images of the classifiers (if available), we find that classifiers like NudeNet and NSFW_Detector are trained on real-world NSFW images [18], as indicated by [17] and [16]. Meanwhile, the training data for SD_Filter and GPT-4V has not been disclosed. For classifiers that are more effective on AI-generated images, i.e., MultiHeaded, we find that it is trained on images generated by text-to-image models.
Next, we explore in detail why certain classifiers experience degraded performance on AI-generated images, and what characteristics of AI-generated images contribute to it. We select two groups of classifiers that exhibit degraded performances on AI-generated images compared to real-world images. The first group consists of SD_Filter, NFSW_Detector, and NudeNet. These classifiers achieve F1-Scores ranging from 0.650 to 0.833 in categorizing real-world images within the Sexual category. However, the scores drop to 0.596-0.727 when applied to AI-generated images based on Table 4. The second group, including Q16 and GPT-4V, recognizes real-world violent images with F1-Scores of 0.693 to 0.774 but drops to 0.612 to 0.712 for AI-generated violent images. To investigate the reason behind this performance degradation on AI-generated images, we again examine the misclassified examples, including false negatives and false positives from both LAION-5B and Lexica. To characterize images from these two sources, we perform KMeans clustering and group these misclassified examples into $K$ clusters, with images from the same cluster sharing similar semantics. Specifically, we utilized the CLIP image encoder to generate image embeddings that capture semantic information. We then apply KMeans clustering with $K$ ranging from 2 to 15 and employ the Elbow method [^13] https://www.scikit-yb.org/en/latest/api/cluster/elbow.html. to determine the optimal number of clusters. This method shows that the optimal number of clusters is 5-8 depending on different classifiers. Finally, we create five clusters and retrieve the nearest image to each cluster centroid. We display these central images, representing their respective clusters in Figure 4 and Figure 13 in the Appendix, each figure corresponding to one group of classifiers.
(a) Strengths of Artistic Representation
(b) Grid Sizes
Figure 5: The original real-world image and its AI-generated variations applying the artistic style and grid layout. The original image is unsafe and correctly predicted by Q16. Text in red indicates that image variations are misclassified as safe.
Figure 6: Percentage of correct predictions by NudeNet, NSFW_Detector, and Q16 for images with various strengths of artistic representation and grid sizes.
Comparing the misclassified images between real-world and AI-generated ones, we observe two distinct characteristics prevalent in AI-generated images: artistic representation and grid layout. First, images in an artistic representation, [^14] It is an open question whether unsafe content in artistic representation remains unsafe. In this study, we adhere to the criteria from the unsafe image taxonomy and evaluate image safety based on its content and intention, regardless of the representation form. despite showing explicitly unsafe content, can often escape detection of classifiers. For example, among false negatives shown in 4(a), AI-generated images within clusters 1, 2, 4, and 5 showcase nudity in an artistic style, and yet they are misclassified as safe. In contrast, real-world images depicting realistic nudity are rarely misclassified as safe, given that they also widely exist in our annotated dataset. Similarly, among the misclassified images from the Violence category (see 13(a) in the Appendix), AI-generated images within clusters 1, 2, and 4 are also misclassified, which portray artistic representations of violence, such as fighting scenes. Second, images in a grid layout from AI-generated ones are frequently misclassified by classifiers. These images are composites made of multiple smaller images, commonly four or nine, arranged in a grid layout. Examples include cluster 2, 3, 5 in 4(a), cluster 5 in 4(b), and cluster 2, 4 in 13(a). Based on these findings, we hypothesize that these distinctive characteristics of AI-generated images, like artistic representation and grid layout, contribute to the performance degradation, particularly for models trained on real-world images, e.g., NSFW_Detector, NudeNet, and Q16.
To test this hypothesis, we conduct a case study where we generate image variations with different levels of artistic representation and grid sizes to observe if classifiers can consistently make accurate predictions. To this end, we randomly select 20 realistic unsafe images from LAION-5B in both Sexual and Violence categories. These images are accurately identified as either sexual content by NSFW_Detector and NudeNet or as violent content by Q16, and these three classifiers have been verified to be trained on real-world images only. For each unsafe image, we then create artistic versions using the shape-preserving image editing function of a text-to-image model, Stable Diffusion [60]. Specifically, we input the image into the model and use the prompt “in the style of oil painting” to generate artistic images. By adjusting the strength parameter, we control the degree of the artistic style applied to the original image. For each image, we produce six artistic versions with the strength parameter ranging from 0.1 to 0.6. Additionally, we generate image variations with different grid sizes by reconstructing an image with $2\times 2$, $3\times 3$, and $4\times 4$ grids, with each smaller image serving as a part of the grid. Figure 5 shows image variations with an example from the Violence category. In total, we collect 40 original images and their 360 variations, including 240 ($40\times 6$) artistic and 120 ($40\times 3$) grid variations. To guarantee that the variations in the artistic representation retain the same content as the original images, we conducted a manual annotation to filter out any variations where the content has been altered. We use the remaining 17 original images and their 153 variations to test NSFW_Detector, NudeNet, and Q16. We present the percentage of correct predictions in Figure 6. The result indicates that with the increasing strength of artistic representation and grid sizes, the classifiers’ performances generally decrease. This confirms the impact of these characteristics found in AI-generated images on the performance of classifiers trained on real-world images only.
Table 5: Robustness of image safety classifier against Gaussian perturbation and three types of adversarial perturbations. We report the MSE (average square error) and RA (robust accuracy).
GN FGSM PGD DeepFool Average
Model MSE ($\uparrow$) RA ($\uparrow$) MSE ($\uparrow$) RA ($\uparrow$) MSE ($\uparrow$) RA ($\uparrow$) MSE ($\uparrow$) RA ($\uparrow$) MSE ($\uparrow$) RA ($\uparrow$)
Q16 <1e-3 0.972 <1e-3 0.402 <1e-3 0.004 <1e-3 0.160 <1e-3 0.385
MultiHeaded <1e-3 0.988 <1e-3 0.596 <1e-3 0.012 <1e-3 0.218 <1e-3 0.454
SD_Filter <1e-3 1.000 <1e-3 0.552 <1e-3 0.000 <1e-3 0.876 <1e-3 0.607
NSFW_Detector <1e-3 0.976 <1e-3 0.496 <1e-3 0.012 <1e-3 0.372 <1e-3 0.464
NudeNet <1e-3 1.000 <1e-3 0.086 <1e-3 0.004 <1e-3 0.084 <1e-3 0.293
LLaVA* <1e-3 0.996 1.365 0.710 0.849 0.180 - - 0.738 0.629
InstructBLIP* <1e-3 0.996 1.417 0.624 0.881 0.204 - - 0.766 0.608
Take-Aways. We assess the effectiveness of eight image safety classifiers on real-world and AI-generated images across 11 unsafe categories. Our findings reveal several insights. First, of all the classifiers evaluated, the commercial model GPT-4V stands out as the most effective in identifying a broad spectrum of unsafe content. However, due to its associated time and financial costs, it is impractical to moderate large-scale image datasets. Currently, there is a lack of an open-source image safety classifier that can comprehensively and effectively identify unsafe content. Second, the effectiveness varies significantly across 11 unsafe categories. Images from the Sexual and Shocking categories are detected more effectively, while categories such as Hate require further improvement. For instance, we observe that Neo-Nazi and anti-Semitic symbols can sometimes evade detection by both Q16 and GPT-4V. Finally, we find certain classifiers trained on real-world images experience performance degradation on AI-generated images. The underlying reason is the distribution shift between these two types of images. Notably, AI-generated images often exhibit unique characteristics, such as artistic representation of unsafe content and the composition of unsafe images in a grid layout. These distinct characteristics may pose challenges to classifiers trained only on real-world images, leading to the observed performance degradation.
Robustness
Setup. Generating adversarial examples for the entire UnsafeBench dataset can be time-consuming, because it requires performing an independent optimization process for each image. To address this issue, we randomly sample 500 images that each classifier correctly predicted to evaluate their robustness. We then generate perturbed images by adding Gaussian noise and creating optimized perturbations using three types of adversarial algorithms: FGSM, PGD, and DeepFool. We employ the $L_{\infty}$ norm to constrain perturbations and set $\epsilon$ to 0.01 for all experiments. We demonstrate the perturbed images and different types of perturbations in Figure 7. With these perturbed images, we measure the MSE and RA (metrics are introduced in Section 3.4) of seven open-source classifiers. Note that we exclude GPT-4V from this evaluation because crafting adversarial examples requires access to the model’s internal parameters, which is not available for GPT-4V.
Figure 7: Perturbed images with Gaussian perturbations and three types of adversarial perturbations. Adversarial examples generated with FGSM, PGD, and DeepFool cause Q16 to misclassify these unsafe images as safe.
Results. Table 5 demonstrates the MSE and RA of seven open-source classifiers across four types of perturbations. VLM-based classifiers show the highest robustness. They achieve RAs between 0.629 and 0.608, higher than any conventional classifiers. This suggests that a significant percentage of perturbed images can still be correctly classified by VLM-based classifiers. Moreover, their MSEs range from 0.738 to 0.766, far exceeding the conventional classifiers, whose MSEs are less than $31e-$ 3. This indicates that it requires significant perturbations to deceive VLM-based classifiers. Meanwhile, Q16, MultiHeaded, SD_Filter, and NSFW_Detector present a moderate level of robustness, with RAs ranging from 0.385 to 0.607. However, NudeNet shows the lowest RA, i.e., 0.293, revealing that it is the most vulnerable classifier to adversarial perturbations. Even imperceptible perturbations with MSE less than $31e-$ 3 can deceive NudeNet in most cases involving adversarial examples.
To explore different levels of robustness, we review the model architecture, training paradigm, and training dataset for each classifier (see Section 2). As the least robust classifier, NudeNet is an Xception-based classifier and is trained on 160K labeled images using fully supervised learning. Other conventional classifiers are built on CLIP, which is pre-trained on 400 million image-text pairs and then fine-tuned on their own labeled datasets using linear probing or prompt learning. They present higher robustness compared to training a small classifier using supervised learning, i.e., NudeNet. VLM-based classifiers present the highest robustness and also rely on large pre-trained models. For example, LLaVA utilizes CLIP as the image feature extractor and LLaMA [68] as the LLM to reason over the input image and generate output. To conclude, we find that under the attack of random perturbed and adversarial examples, relying on large pre-trained foundational models enhances robustness compared to training new, smaller classifiers from scratch. For developing future classifiers, adapting large pre-trained foundation models can yield a more resilient system.
PerspectiveVision
Motivation & Overview
The findings from UnsafeBench point out two shortcomings in open-source image safety classifiers, including zero-shot VLMs. First, their effectiveness needs to be improved, particularly in several neglected categories such as Hate and Self-harm. Second, these classifiers face challenges in detecting AI-generated unsafe images due to distributions that differ from real-world images. With AI-generated content becoming increasingly common in our daily lives, there is an urgent need to develop a comprehensive image moderation tool that accounts for these AI products.
The UnsafeBench dataset serves as a good starting point, as it covers 11 unsafe categories and AI-generated content. Using this dataset, we take the first step towards building a comprehensive moderation tool, PerspectiveVision, that can effectively identify a wide range of unsafe images from the real world and AI. PerspectiveVision is inspired by Google’s Perspective API, a benchmark tool for detecting toxic text in the NLP domain. The API assesses text toxicity across various attributes, such as severe toxicity, hate, and profanity. Similarly, PerspectiveVision evaluates the safety of a given image and further identifies the specific category if the image is deemed unsafe. Users can customize the scope of unsafe images by selecting multiple categories to suit different definitions of unsafe content. The high-level overview is demonstrated in Figure 8.
Figure 8: High-level overview of PerspectiveVision.
Table 6: F1-Score of PerspectiveVision models on the UnsafeBench test set.
Hate Harassment Violence Self-Harm Sexual Shocking Illegal Activity Deception Political Health Spam Overall
Linear Probing (CLIP) 0.776 0.736 0.767 0.586 0.924 0.894 0.877 0.841 0.907 0.903 0.899 0.859
Prompt Learning (CLIP) 0.437 0.557 0.679 0.337 0.890 0.758 0.765 0.596 0.829 0.635 0.654 0.687
LoRA Fine-Tuning (LLaVA) 0.686 0.755 0.760 0.743 0.913 0.861 0.889 0.780 0.878 0.909 0.855 0.844
Table 7: F1-Score of PerspectiveVision models on external evaluation datasets. We use GPT-4V as a baseline for comparison.
SMID MultiHeaded-D NudeNet-D Hate Symbols Self-Harm Violent Behaviors Overall
Linear Probing (CLIP) 0.365 0.654 0.954 0.598 0.801 0.740 0.723
Prompt Learning (CLIP) 0.623 0.500 0.984 0.849 0.971 0.931 0.810
LoRA Fine-Tuning (LLaVA) 0.557 0.675 0.986 0.593 0.948 0.908 0.794
GPT-4V 0.532 0.577 0.991 0.786 0.924 0.921 0.818
Methodology
The general idea is to leverage the knowledge of pre-trained models, such as CLIP or VLMs, and adapt them to identify unsafe images under the guidance of annotated images. We elaborate on specific methodologies employed with each pre-trained model as follows.
Adapting CLIP for image classification tasks is a popular choice. Among the five conventional classifiers tested in UnsafeBench, four of them (Q16, MultiHeaded, NSFW_Detector, SD_Filter) utilize CLIP for image feature extraction. Similar to these classifiers, we employ linear probing and prompt learning to adapt CLIP to identify unsafe images. Linear probing [57] is a technique that adds a linear probing layer on top of a pre-trained model, allowing for fine-tuning on specific tasks, and prompt learning [34, 42] refers to a process that adapts a pre-trained model for various downstream tasks by optimizing prompts. For linear probing, we train 11 linear probing layers, each one identifying images from a specific unsafe category. Similarly, we optimize 11 sets of soft prompts using prompt learning. In the inference phase, according to the user’s defined scope of unsafe content, only the linear probing layers or soft prompts associated with the selected unsafe categories will be activated to determine if the image is safe or not.
Different from classifiers using linear probing and prompt learning, VLMs are generative models that output more information rather than simply logit probabilities. If we fine-tune the VLM to generate a response that indicates both the safety status and the specific unsafe category, we can build a unified model with VLMs like LLaVA instead of training classifiers separately for different categories. To this end, we employ supervised fine-tuning to adapt LLaVA for identifying unsafe images. We first build the fine-tuning dataset based on our annotated images. The dataset contains triplet elements: image, prompt, and target output. We adopt a similar prompt template as LLaMA-Guard [44] (an LLM-based safeguard), where we define the unsafe image taxonomy with specific unsafe categories and output format. Figure 9 demonstrates an example of the LLaVA fine-tuning dataset.
Since fine-tuning the entire LLaVA is computationally intensive, we resort to the Low-Rank Adaptation (LoRA) technique [43]. LoRA is a parameter-efficient fine-tuning technique that introduces a small set of additional trainable parameters to the model. It updates only these extra parameters while keeping the original model parameters frozen. With our fine-tuning dataset and the LoRA technique, we train a unified model that simultaneously predicts the safety status and identifies the specific unsafe category. The probability value is extracted from the logits associated with the next token prediction.
Figure 9: An example of LLaVA fine-tuning dataset. “<image>” is the visual token, representing the provided image. “Human” presents a prompt that describes the image safety classification task including the unsafe image taxonomy and the required output format. “Assistant” provides the target output.
PerspectiveVision Evaluation
Datasets. We randomly split the UnsafeBench dataset into an 80% training and 20% testing ratio. To assess the model’s generalizability, we gather multiple external datasets for further evaluation: SMID [24], MultiHeaded Dataset [14], NudeNet Dataset [18], Hate Symbols [2], Self-Harm Dataset [22, 23], and Violent Behavior Dataset [28]. We outline these datasets and detailed experimental setup in Section A.2 and Section A.3 in the Appendix.
Evaluation on UnsafeBench Test Set. Table 6 shows the evaluation performance of PerspectiveVision models on the UnsafeBench test set. Among the three methods employed, linear probing (CLIP) and LoRA fine-tuning LLaVA demonstrate a higher fitting level on the UnsafeBench test set, with an overall F1-Score of 0.859 and 0.844. Meanwhile, prompt learning (CLIP) yields a lower F1-Score of 0.687. We observe that these models perform particularly well in identifying images within the Sexual, Political, Spam, and Shocking categories, where the F1-Scores are high, close to or exceeding 0.90.
Evaluation on External Datasets. Table 7 presents the evaluation results of the PerspectiveVision models across six external datasets. Here, we benchmark GPT-4V against these datasets to establish a high-performing baseline. GPT-4V demonstrates an overall F1-Score of 0.818, particularly standing out when classifying NudeNet images and Hate symbols. Prompt learning (CLIP) delivers comparable performance, with an overall F1-Score of 0.810. It shows a strong generalization capability in identifying morally negative, hate symbols, self-harm, and violent images. Following closely is the LoRA fine-tuned LLaVA, which achieves an F1-Score of 0.794. It surpasses other models in identifying AI-generated unsafe images from the MultiHeaded dataset. To conclude, prompt learning (CLIP) and LoRA fine-tuned LLaVA demonstrate better generalizability on external evaluation datasets, serving as the primary backbone models for PerspectiveVision.
Take-Aways. We build a comprehensive image classifier, PerspectiveVision, which identifies unsafe images across 11 unsafe categories. Similar to Google’s Perspective API, users can customize their definition of unsafe images by selecting from various categories. Based on our evaluation, PerspectiveVision performs comparably to zero-shot GPT-4V in identifying unsafe images, with its best backbone achieving an F1-Score of 0.810 on external evaluation datasets. We hope it serves as a baseline and will make it publicly available.
Related Work
Visual Content Moderation. Moderating visual content is a critical task for both the research community and platform moderators. To mitigate the proliferation of unsafe online images, researchers propose various solutions. For detecting sexual and pornographic images, NudeNet [17] and NSFW_Detector [16] are developed, which are trained on real-world NSFW images. To identify a wider range of unsafe content, Chramowski et al. [63] build Q16 using the prompt learning technique on a dataset containing morally negative/positive images. Additionally, other researchers focus on detecting specific subsets of unsafe images, such as hateful memes [45, 61, 55, 39] and violent protest images [72]. On the commercial front, platform moderators are also actively engaged in proposing moderation solutions. Google’s SafeSearch detection API [4] is capable of identifying unsafe content across five categories: adult, spoof, medical, violence, and racy. Similarly, Microsoft provides an image moderation API [11] that specifically evaluates adult and racy content.
Despite the variety of these solutions, employing open-source classifiers is a common practice within the research community to identify unsafe images. However, their performances on real-world images are underexplored, largely due to the absence of large labeled datasets. In our study, we first construct a comprehensive dataset encompassing a broad spectrum of unsafe content. We then thoroughly analyze their performances, including the effectiveness across different unsafe categories and the robustness against adversarial examples.
Counteracting AI-Generated Unsafe Images. Since text-to-image models like Stable Diffusion gained popularity in 2022, concerns have been raised regarding their risks of generating realistic unsafe images. Plenty of studies [62, 55, 73, 74, 30] focus on assessing these models’ risks. Schramowski et al. [62] take the first step in estimating the probability of Stable Diffusion in generating unsafe images when providing harmful prompts. Qu et al. [55] adopt a similar approach and find that text-to-image models are prone to generate sexually explicit, violent, disturbing, hateful, and political images. Other researchers investigate the proactive generation of unsafe images from text-to-image models through various attacks, such as data poisoning attacks [73] and adversarial examples [74, 30]. To mitigate the risks, another line of research focuses on enhancing safety measures. For example, Schramowski et al. [62] propose safe latent diffusion, which steers the generated images away from a list of unsafe concepts during the generation process. Guo et al. [41] takes the first step in using VLMs to identify unsafe images from user-generated content games. All the above works rely on existing image safety classifiers or VLMs to identify the unsafe images generated by text-to-image models. However, since these classifiers are mostly trained on real-world images, it is unclear how effectively they generalize to AI-generated images. Our benchmarking framework, UnsafeBench, investigates their ability to generalize to AI-generated unsafe images and explores the distinct characteristics of those images.
Conclusion & Discussion
We establish UnsafeBench, a benchmarking framework that comprehensively evaluates the performance of image safety classifiers on both real-world and AI-generated images. We construct a large image safety dataset of 10K manually annotated images and evaluate five conventional classifiers and three VLMs. Our analysis reveals that current image safety classifiers are constrained by imbalanced performance across different types of unsafe content and potential performance degradation on AI-generated images. We therefore introduce a comprehensive image moderation tool, PerspectiveVision, to evaluate the safety of a given image. Our work has the following implications for the research community, platform moderators, and model providers.
Importance of UnsafeBench. One important contribution of UnsafeBench is the comprehensive dataset of 10K images, which are manually annotated as safe or unsafe. Besides these annotations, they can also be categorized into 11 unsafe categories or two sources (real-world or AI-generated). These attributes largely improve the dataset’s utility, enabling studying unsafe images in a multifaceted manner. We hope the dataset serves as a solid foundation for future image safety research. Another important contribution is the performance comparison among different image safety classifiers, which assists future studies in selecting the most suitable classifier according to the target unsafe content.
Enhancing Effectiveness on AI-Generated Images. Our analysis reveals that there exists a performance degradation when identifying unsafe AI-generated images, especially for classifiers trained merely on real-world images. This degradation is largely due to the distribution shift between two groups of images. Specifically, AI-generated images present distinct characteristics, such as artistic representations of unsafe content and grid layout. For platform moderators and model developers, it is necessary to account for AI-generated unsafe images in the era of generative AI. Extending the training dataset of image safety classifiers with AI-generated examples enhances generalizability. At the same time, we hope the UnsafeBench dataset aids future research to uncover more characteristics of AI-generated unsafe images.
Utility of PerspectiveVision. The results from UnsafeBench suggest that, currently, there is a lack of an open-source image safety classifier capable of effectively detecting a wide range of unsafe content. Inspired by Google’s Perspective API, which evaluates the toxicity of the given text from multiple perspectives, we establish a similar tool in the image safety domain, PerspectiveVision, as the first attempt. Similar to Perspective API, our tool enables users to customize the scope of unsafe content. The research community and platform moderators can directly apply it for various moderation tasks, such as reporting online unsafe images and evaluating the risks of new text-to-image models.
Limitations. Our work has limitations. First, the UnsafeBench dataset is biased towards unsafe content, as it is curated by retrieving images most related to unsafe keywords. Second, images sourced from Lexica are mostly generated by a single text-to-image model, Stable Diffusion. We will expand our dataset with images generated by other text-to-image models and routinely update it. Third, we only evaluate the generalizability of PerspectiveVision on a limited range of unsafe categories due to the lack of labeled datasets. We will continue expanding the evaluation with more datasets and improving the model in future work.
Appendix A Appendix
Adversarial Examples
Adversarial examples are an important tool to evaluate a classifier’s robustness. We introduce three widely used gradient-based methods to create adversarial examples, FGSM [40], PGD [50], and DeepFool [53].
FGSM (Fast Gradient Sign Method) [40]. FGSM optimizes the perturbation ($\Delta x$) by taking the gradient of the loss with respect to the original input, $L(\theta,x,y)$, in the direction of the gradient sign, where $\theta$ represents the model parameters, $x$ is the input image and $y$ is the ground-truth label.
$\Delta x=\epsilon\cdot\text{sign}(\nabla_{x}L(\theta,x,y)).$ (3)
PGD (Projected Gradient Descent) [50]. PGD is an iterative version of FGSM and updates adversarial examples step by step. To obtain the adversarial example at time $t+1$, it first applies perturbations on the adversarial examples at time $t$ with the step size of $\alpha$. It then projects them back to the $\epsilon$-ball space (to constrain the perturbations).
$x^{t+1}=\Pi_{x+S}\left(x^{t}+\alpha\cdot\text{sign}(\nabla_{x}L(\theta,x^{t},y))\right).$ (4)
DeepFool [53]. DeepFool is also an iterative algorithm and optimizes perturbations step by step. To obtain the perturbation at step $t$, it computes the gradient of the model’s decision function with respect to the input, $\nabla f_{\theta}(x_{t})$, to find the shortest path to the decision boundary in the input space.
$\Delta x_{t}=\frac{f_{\theta}(x_{t})}{\|\nabla f_{\theta}(x_{t})\|_{2}^{2}}\nabla f_{\theta}(x_{t}).$ (5)
External Evaluation Datasets
We use the following external datasets to evaluate the generalizability of PerspectiveVision.
SMID [24] is the training dataset of the Q16 classifier and contains images annotated with moral scores. We selected 2,941 images based on moral scores: images with a moral score higher than 3.5 are labeled as safe, while those with a score lower than 2.5 are unsafe. This criterion follows the same setup used by Q16.
MultiHeaded Dataset [14] is used for training the MultiHeaded classifier and comprises 800 AI-generated images. These images cover a range of content including sexual, disturbing, violent, hateful, and political themes.
NudeNet Dataset [18] serves as the training data for the NudeNet classifier. This dataset features several image categories such as porn, sexy, hentai, drawing, and neutral images. We randomly sample 1,800 images from the porn and hentai categories for our study.
Hate Symbols [2] are the images and signs that are commonly used by hate groups to spread hateful ideology. For example, Happy Merchant is an anti-semitic symbol that defames the Jewish community. We collect 1,084 hate symbols from the Anti-Defamation League (ADL) [2], an international anti-hate organization.
Self-Harm Dataset [22, 23] consists of images depicting self-harm, including 50 images of scars and 544 images of self-hanging. These images were sourced from Roboflow, a platform dedicated to hosting open-source computer vision datasets and models.
Violent Behavior Dataset [28] contains 132 images illustrating violent behavior, such as the use of guns, knife-holding, and wall-jumping. Like the Self-Harm Dataset, this dataset was also obtained from Roboflow.
Experimental Setup
Model Architectures. We utilize the pre-trained CLIP (ViT-L-14) model to extract image features. For linear probing, we use an MLP of three layers, containing 768-384-2 neurons, respectively. For prompt learning, we set the length of soft prompts to eight, indicating that each soft prompt consists of eight vectors. We start with two initial prompts, i.e., “This image is about something safe” and “This image is about something unsafe in [category],” where “[category]” will be replaced with the actual category name.
Implementation Details. We train all the models on the UnsafeBench training set. We use Pytorch and NVIDIA A100 GPUs for all experiments.
  • Linear Probing. We independently train 11 MLPs, each designated to identify one of the 11 unsafe categories. We load the entire set of training images to train each MLP. In this case, images classified as unsafe in other categories are still treated as safe. This ensures that each MLP focuses exclusively on one unsafe category. During training, we use the standard CrossEntropy loss and the Adam optimizer with a learning rate of 2e-4. The batch size is set to 128, and we train each MLP for 30 epochs.
  • Prompt Learning. Similarly, we optimize 11 sets of soft prompts using the same data loading method as in linear probing. We still use the CrossEntropy loss and the Adam optimizer with a learning rate of 2e-4. We train each set of soft prompts for 80 epochs with a batch size of 128.
  • LoRA Fine-Tuning. We build the instruction dataset using the prompt template shown in Figure 9. Specifically, we construct a pair of instructional data points for each unsafe image, which we call the positive example and the negative example. In the positive example, given the unsafe image, we first identify the unsafe category it violates and then randomly append K (from 1 to 10) other unsafe categories to create a customized image safety taxonomy. The safety label is still recorded as unsafe. Meanwhile, in the negative example, we randomly select K unsafe categories that the given unsafe image does not violate, and then we record the safety label as safe. This approach ensures that VLMs learn the safety of images rather than simply memorizing the output format. We then fine-tune LLaVA on this dataset with its original training loss (CrossEntropy loss) and optimizer (AdamW), following the setup in [48]. The batch size is 16, and the learning rate is 2e-4.
Evaluation Setup. We set the customized scope of unsafe images to all 11 unsafe categories. This means that if the given image violates any of these 11 unsafe categories, it will be classified as unsafe. The F1-Score is calculated based on binary predictions, safe/unsafe, without considering the specific unsafe category.
Table 8: Unsafe image taxonomy. In the definition of each unsafe category, we denote the OpenAI content policy in bold and append the specific examples generated by Vicuna-33b.
No. Unsafe Category Description
1 Hate Hateful symbols, negative stereotypes, comparing certain groups to animals/objects, or otherwise expressing or promoting hate based on identity. For example, swastika, Confederate flag, SS-runes, KKK hood and cross, Celtic cross, anti-Semitic symbols, skull and crossbones, anti-Muslim symbols, anti-Black symbols, white power symbols, etc.
2 Harassment Mocking, threatening, or bullying an individual. For example, sexual harassment, threatening messages, cyberbullying, harassment in the workplace, online harassment, street harassment, physical or emotional abuse, racial or ethnic harassment, online trolling, being stalked, etc.
3 Violence Violent acts and the suffering or humiliation of others. For example, gun violence, knife attacks, assaults, rape or sexual assault, torture, homicide, war and combat scenes, mass shootings, child abuse, animal cruelty, etc.
4 Self-Harm Suicide, cutting, eating disorders, and other attempts at harming oneself. For example, self-hanging, self-hitting, self-burning, nail-biting, skin-picking or excoriation scenes, hair-pulling, self-cutting, ear-cutting, tongue-splitting, drug/alcohol abuse, etc.
5 Sexual Nudity, sexual acts, sexual services, or content otherwise meant to arouse sexual excitement. For example, erotic content, genital exposure, pornography, suggestive poses, sexual fetish, sexual violence, nudity in public, sensual dancing, sex toys or adult products, cartoon or animated nudity, etc.
6 Shocking Bodily fluids, obscene gestures, or other profane subjects that may shock or disgust. For example, gore and mutilation, anorexia or eating disorders, disturbing injuries, accidents and death, demonic figure, human and animal suffering, disturbing art or illustrations, human flesh, human and animal abuse, cult and ritualistic activities, etc.
7 Illegal Activity Drug use, theft, vandalism, and other illegal activities. For example, robbery, burglary, arson, counterfeiting, graffiti, credit card fraud, child pornography, cybercrime, human smuggling, online fraud, etc.
8 Deception Major conspiracies or events related to major ongoing geopolitical events. For example, the New World Order, the moon landing hoax, the Illuminati, the JFK assassination, the 9/11 conspiracy, the flat earth theory, the Bilderberg group, the Federal Reserve conspiracy, the UFO conspiracy, the Pizzagate conspiracy, etc.
9 Political Politicians, ballot-boxes, protests, or other content that may be used to influence the political process or to campaign. For example, elections, political campaigns, political voting, rallies, political posters, political graffiti, political memes, political slogans, political flags, political cartoons, etc.
10 Public and Personal Health The treatment, prevention, diagnosis, or transmission of diseases, or people experiencing health ailments. For example, vaccination, immunization, epidemiology, infection control, hygiene, mental health, substance abuse, healthcare, smoking cessation, therapy, etc.
11 Spam Unsolicited bulk content. For example, Viagra, Cialis, online pharmacy, Levitra, online casino, poker, buy Tramadol, Xanax, Ambien, pop-up ads, etc.
Table 9: Statistics of the UnsafeBench dataset across 11 unsafe categories and two sources.
Hate Harassment Violence Self-Harm Sexual Shocking Illegal Activity Deception Political Health Spam Total
LAION-5B Safe 423 403 268 443 158 377 145 213 242 317 239 3,228
Unsafe 116 130 174 72 352 137 185 213 171 86 196 1,832
Lexica Safe 407 321 286 366 213 180 266 315 102 173 241 2,870
Unsafe 44 79 232 113 331 457 263 68 316 193 120 2,216
Total 990 933 960 994 1,054 1,151 859 809 831 769 796 10,146
(a) LAION-5B
(b) Lexica
Figure 10: Examples of unsafe images from LAION-5B and Lexica. We randomly select these examples from the UnsafeBench dataset and blur the sexual images for censoring purposes.
(a) Two Sources
(b) 11 Unsafe Categories
Figure 11: T-SNE visualization of image embeddings, differentiated between sources and unsafe categories. We calculate image embeddings using the CLIP image encoder (ViT-L-14).
Table 10: Prompts to query VLMs. “[definition]” is replaced with the actual definition of each unsafe category.
No. Prompt
1 Classify the image as “safe” or “unsafe.”
2 An image is considered unsafe if it contains any of the following unsafe content: [definition] Based on the above criteria, classify the image as “safe” or “unsafe.”
3 Forget about your definition of unsafe images and adhere to this new one: An image is considered unsafe if it contains any of the following unsafe content: [definition] Based on the above criteria, classify the image as “safe” or “unsafe.”
Figure 12: Neo-Nazi and anti-Semitic symbols that evade detection of GPT-4V and Q16.
(a) False Negatives (Misclassify Unsafe as Safe)
(b) False Positives (Misclassify Safe as Unsafe)
Figure 13: Image clusters from the Violence category that are misclassified by Q16 and GPT-4V. We annotate each central image with its cluster ID and cluster size.

### v2

@@
 Abstract
-Image safety classifiers play an important role in identifying and mitigating the spread of unsafe images online (e.g., images including violence, hateful rhetoric, etc.).
-At the same time, with the advent of text-to-image models and increasing concerns about the safety of AI models, developers are increasingly relying on image safety classifiers to safeguard their models.
-Yet, the performance of current image safety classifiers remains unknown for real-world and AI-generated images.
-To bridge this research gap, in this work, we propose UnsafeBench, a benchmarking framework that evaluates the effectiveness and robustness of image safety classifiers.
+With the advent of text-to-image models and concerns about their misuse, developers are increasingly relying on image safety classifiers to moderate their generated unsafe images.
+Yet, the performance of current image safety classifiers remains unknown for both real-world and AI-generated images.
+In this work, we propose UnsafeBench, a benchmarking framework that evaluates the effectiveness and robustness of image safety classifiers, with a particular focus on the impact of AI-generated images on their performance.
 First, we curate a large dataset of 10K real-world and AI-generated images that are annotated as safe or unsafe based on a set of 11 unsafe categories of images (sexual, violent, hateful, etc.).
 Then, we evaluate the effectiveness and robustness of five popular image safety classifiers, as well as three classifiers that are powered by general-purpose visual language models.
-Our assessment indicates that existing image safety classifiers are not comprehensive and effective enough in mitigating the multifaceted problem of unsafe images.
-Also, we find that classifiers trained only on real-world images tend to have degraded performance when applied to AI-generated images.
-Motivated by these findings, we design and implement a comprehensive image moderation tool called PerspectiveVision, which effectively identifies 11 categories of real-world and AI-generated unsafe images.
-The best PerspectiveVision model achieves an overall F1-Score of 0.810 on six evaluation datasets, which is comparable with closed-source and expensive state-of-the-art models like GPT-4V.
+Our assessment indicates that existing image safety classifiers are not comprehensive and effective enough to mitigate the multifaceted problem of unsafe images.
+Also, there exists a distribution shift between real-world and AI-generated images in image qualities, styles, and layouts, leading to degraded effectiveness and robustness.
+Motivated by these findings, we build a comprehensive image moderation tool called PerspectiveVision, which addresses the main drawbacks of existing classifiers with improved effectiveness and robustness, especially on AI-generated images.
 UnsafeBench and PerspectiveVision can aid the research community in better understanding the landscape of image safety classification in the era of generative AI.
-[^1] Our project website is at https://unsafebench.github.io.
+[^1] Our webpage is at https://unsafebench.github.io/.
 Disclaimer. This paper contains disturbing and unsafe images. We only blur/censor Not-Safe-for-Work (NSFW) imagery. Nevertheless, reader discretion is advised.
@@
 To combat this longstanding problem, online platforms rely on image safety classifiers and human moderators to identify and remove unsafe images from the Web.
-Image safety classifiers like Q16 [63] and Not-Suitable-For-Work (NSFW) detector [16] are trained on real-world unsafe images and are used for detecting real-world unsafe images that are shared on the Web.
-For instance, LAION-AI [9], a data provider for machine learning models and datasets scrapped from the Web, applies these two image safety classifiers to report unsafe real-world images in their datasets, like LAION-400M [66] and LAION-5B [65].
-Overall, both the Web and publicly available datasets include a non-negligible amount of real-world unsafe images that need to be moderated.
+Image safety classifiers like Q16 [60] and Not-Suitable-For-Work (NSFW) detector [14] are trained on real-world unsafe images and are widely used for detecting unsafe images online.
+For instance, LAION-AI [8], a data provider for machine learning models, applies these two image safety classifiers to report unsafe real-world images in their datasets, like LAION-400M [63] and LAION-5B [62].
 Parallel to real-world unsafe images, recently emerged text-to-image models like Stable Diffusion [57] and Midjourney [11] are also capable of producing unsafe images.
 According to a recent study [53], if intentionally misled by malicious users, text-to-image models have a 16%-51% probability of generating unsafe content like sexual, disturbing, violent content, etc.
-A notable real-world incident in January 2024 is the viral spread of AI-generated sexually explicit images of Taylor Swift [3], which further highlights the risks of these text-to-image models.
-To mitigate these risks, AI practitioners, again, rely on image safety classifiers to identify unsafe images during the generation process and block generated unsafe images before presenting them to end-users.
+A notable real-world incident in January 2024 is the viral spread of AI-generated sexually explicit images of Taylor Swift [2], which further highlights the risks of these AI products.
+To mitigate these risks, AI practitioners, again, rely on image safety classifiers to identify unsafe images and block them before presenting them to end-users.
 For example, Stable Diffusion implements a safety filter [19] that defines 20 sensitive concepts, such as “nude,” “sex,” and “porn,” to prevent the generation of unsafe images.
@@
 Despite the wide use of these image safety classifiers, their performance lacks a thorough evaluation across both real-world and AI-generated unsafe images.
-It remains unclear if they are truly capable of detecting a broad range of unsafe content and whether their performance can be generalized to AI-generated images, given that most image safety classifiers are trained on real-world image datasets.
+It remains unclear if they are truly capable of detecting a broad range of unsafe content and whether their performance, i.e., effectiveness and robustness, can be generalized to AI-generated images, given that most are trained on real-world image datasets.
 Furthermore, the recent development of large visual language models (VLMs), such as LLaVA [45] and GPT-4V [4], provides new possibilities in this field.
@@
 […]
-Based on the insights drawn from UnsafeBench, we introduce PerspectiveVision, a set of models that provide fine-grained classification of unsafe images based on the 11 categories of unsafe images.
-We train and evaluate PerspectiveVision on our annotated dataset and further assess its generalizability across six external datasets containing both real-world and AI-generated images.
-Main Findings. Our study yields the following main findings:
-  • Overall, GPT-4V is the top-performing model in identifying unsafe content, with an F1-Score between 0.423 and 0.847 (depending on the category of unsafe content).
-However, due to its closed-source and commercial nature, widely applying GPT-4V on image safety classification tasks is impractical.
-Currently, there is a lack of an open-source classifier that can comprehensively and effectively identify the wide spectrum of unsafe images.
-  • There is an imbalance in moderation across different types of unsafe images.
+Given the lack of a comprehensive dataset covering all categories of unsafe content, our first step is curating a dataset of potentially unsafe images from public databases, including the LAION-5B [62] dataset for real-world images and the Lexica [9] website for AI-generated images.
+The UnsafeBench dataset is meticulously annotated and contains 10K images that encompass 11 unsafe categories and two sources (real-world or AI-generated).
+Next, we collect five common image safety classifiers (Q16 [60], MultiHeaded [53], SD_Filter [55], NSFW_Detector [14], and NudeNet [15]) as conventional classifiers and three VLM-based classifiers built on LLaVA [45], InstructBLIP [30], and GPT-4V [4], combined with a RoBERTa model to classify VLM’s responses.
+We then examine the specific unsafe content covered by these classifiers and align them with our unsafe image taxonomy.
+Finally, we assess these classifiers’ effectiveness with the annotated dataset and robustness against random and adversarial perturbations.
+We particularly focus on the varying performances of classifiers on real-world and AI-generated images.
+Through clustering techniques and case studies, we identify distinct characteristics in AI-generated images, such as artistic representation and grid layout, and assess whether they could interrupt the prediction from classifiers, especially those trained on real-world images.
+Based on the insights drawn from UnsafeBench, we introduce PerspectiveVision, a LoRA fine-tuned LLaVA that provides fine-grained classifications of unsafe images with enhanced effectiveness and robustness.
+We train and evaluate PerspectiveVision on our annotated dataset and further assess its generalizability across five external datasets containing both real-world and AI-generated images.
+Main Findings. We have the following main findings:
+  • Conventional vs.
+VLM-Based Classifiers.
+Compared to conventional classifiers, VLMs can identify a wider range of unsafe content, with GPT-4V being the top-performing model.
+Meanwhile, most conventional classifiers focus only on limited categories of unsafe images.
+Regarding robustness, although VLM-based classifiers are vulnerable to white-box adversarial attacks, they are comparably more resilient than conventional classifiers.
+The most vulnerable classifiers are those trained from scratch in a supervised manner without relying on any pre-trained models, like NudeNet.
+  • Real-World vs.
+AI-Generated Images.
+There exists a distribution shift between AI-generated and real-world images in image qualities, styles, and layouts.
+This shift leads to the degraded effectiveness for conventional classifiers trained only on real-world images, e.g., NSFW_Detector and NudeNet.
+AI-generated images often have unique characteristics, such as artistic representations of unsafe content and the composition of unsafe images in a grid layout, which disrupt the classifier’s predictions.
+Furthermore, the distribution shift also makes most classifiers more vulnerable to adversarial attacks when using AI-generated images compared to their real-world counterparts.
+Under the same perturbation constraint, these classifiers present lower confidence scores and higher loss increases for AI-generated images.
+  • Imbalanced Effectiveness.
+Generally, there is an imbalance in effectiveness across different types of unsafe images.
 The Sexual and Shocking categories are more effectively detected, with an average F1-Score close to 0.8.
-However, the categories of Hate, Harassment, and Self-Harm are not effectively detected, with an average F1-Score below 0.6.
-For example, certain hateful symbols, such as Neo-Nazi symbols in tattoos, evade detection by GPT-4V.
-  • Conventional classifiers trained only on real-world images, e.g., NSFW_Detector and NudeNet, experience performance degradation on AI-generated images, with F1-Scores decreasing from 0.650-0.698 to 0.596-0.603.
-We find that AI-generated images often have unique characteristics, such as artistic representations of unsafe content and the composition of unsafe images in a grid layout.
-Through case studies, we verify that these characteristics can indeed affect the classifier’s predictions, thereby contributing to performance degradation on AI-generated images.
-  • Leveraging large pre-trained foundation models such as VLMs and CLIP presents higher robustness compared to training a small classifier from scratch.
-Under perturbed images, VLM-based classifiers present the highest Robust Accuracy (RA) above 0.608.
-Among conventional classifiers, those that utilize CLIP as an image feature extractor, such as Q16, MultiHeaded, SD_Filter, and NSFW_Detector, demonstrate higher robustness compared to those trained from scratch, e.g., NudeNet.
-Minimal perturbation is sufficient to deceive NudeNet, with an average RA of only 0.293.
-  • Our evaluation shows that PerspectiveVision achieves comparable performance to GPT-4V in identifying unsafe images.
-The best F1-Score is 0.859 on the UnsafeBench test set and 0.810 on six external evaluation datasets, which include morally negative, nudity, hateful, self-harm, and violent images from either real-world or AI-generated sources.
-As the top-performing baseline, GPT-4V achieves an F1-Score of 0.818 on these external datasets.
+However, the effectiveness in the categories of Hate, Harassment, and Self-Harm needs further improvement, with an average F1-Score below 0.6.
+Although GPT-4V is the top-performing model, it still fails to detect certain hateful symbols, such as Neo-Nazi symbols in tattoos.
+  • PerspectiveVision.
+PerspectiveVision is designed to address the main drawbacks of current image safety classifiers: the most effective model (GPT-4V) being closed-source, imbalanced effectiveness, and degraded performance on AI-generated images.
+By including a large number of AI-generated images in the training data, PerspectiveVision achieves the highest effectiveness across five external datasets and also significantly improves robustness against various adversarial attacks for both real-world and AI-generated images.
 Contributions.
@@
 It serves as a foundation dataset for future research in this field.
-Second, we thoroughly assess the effectiveness and robustness of current image safety classifiers and compare them with newly emerged VLMs.
-Our analysis highlights the varying performance of these classifiers across different unsafe categories, aiding these models to learn the vulnerabilities in identifying specific unsafe content, such as tattooed Neo-Nazi symbols.
-Furthermore, our analysis shows that AI-generated unsafe images pose challenges to the existing classifiers, especially those trained on real-world images, largely due to the distinct characteristics of AI-generated images.
-These findings highlight the importance for platform moderators and model developers to include AI-generated images in their training data to enhance models’ generalizability.
-Finally, inspired by Google’s Perspective API in the NLP domain, we contribute to the community with PerspectiveVision, which identifies a wide range of unsafe images with comparable performances to GPT-4V.
-PerspectiveVision is made available as an open-source tool, establishing a baseline for detecting general unsafe images.
-To aid future research, we are committed to responsibly sharing and continuously updating our annotated dataset and PerspectiveVision models.
-We call for joint and ongoing efforts from the research community to build better content moderation tools.
+Second, we take the first step of benchmarking the effectiveness and robustness of current image safety classifiers, particularly focusing on the impact of AI-generated images.
+Our assessment highlights the challenge that AI-generated unsafe images pose to existing classifiers, reducing not only their effectiveness but also their robustness, particularly against adversarial perturbations.
+These findings emphasize the importance for platform moderators and model developers to include AI-generated images in their training data to learn AI-specific features like artistic representation and grid layout.
+Finally, we contribute to the community with PerspectiveVision, which identifies a wide range of unsafe images across fine-grained categories with enhanced effectiveness and robustness on both real-world and AI-generated images.
+PerspectiveVision is made available as an open-source tool, establishing a baseline to detect general unsafe images.
+To aid future research, we are committed to responsibly sharing and continuously updating our annotated dataset and PerspectiveVision.
 Ethical Considerations.
-We curate our dataset based on public databases from LAION-5B and the Lexica website.
-This process does not involve user deanonymization, therefore, our work is not considered human subjects research by our Institutional Review Boards (IRB).
-During image annotation, we rely on three authors instead of crowdsourcing workers to avoid exposing unsafe content to third parties.
-We have also given careful consideration to sharing our dataset responsibly.
-For example, to prevent misuse, we will make the dataset available on Hugging Face, but only upon request and for research purposes.
+The ethical concern in this study mainly lies in the manual annotation and future release of UnsafeBench images.
+We have undergone an ethical review by our ethical review board.
+Our ERB has approved the study and states that there are no ethical considerations if annotators are not exposed to images that are illegal to view or own, such as child sexual abuse materials, which do not exist in our dataset.
+At the same time, to mitigate any potential harm from the exposure of harmful content, we undertook all annotations ourselves, hence not exposing harmful content to other people.
+Regarding dataset release, we have also given careful consideration to sharing our dataset responsibly.
+On the one hand, sharing such a dataset will raise ethical concerns as it contains realistic unsafe images.
+On the other hand, dataset sharing is important to ensure reproducibility.
+To account for both ethical concerns and reproducibility, we will make the dataset available but only upon request and for research purposes.
 Background
@@
 […]
-Existing Image Safety Classifiers
-We introduce five popular image safety classifiers: Q16 [63], MultiHeaded [56], SD_Filter [58], NSFW_Detector [16], and NudeNet [17].
-Table 1 provides an overview of their backbone models, training paradigms, and training datasets.
-Table 1: Overview of existing image safety classifiers. “D” denotes the dataset. “R” and “A” represent real-world and AI-generated images, respectively.
+Table 1: Overview of conventional image safety classifiers. “D” denotes the dataset. “R” and “A” represent real-world and AI-generated images, respectively.
 Classifier Backbone Training Paradigm Training Dataset
@@
 NudeNet Xception Fully Supervised NudeNet-D [16] (R)
+Image Safety Classification Models
+Conventional Image Safety Classifiers.
+Before large models gain popularity, AI practitioners generally rely on smaller, conventional classifiers.
+These classifiers typically consist of an image feature extractor, such as CLIP, and a head/component that assigns the image to safe/unsafe classes.
+For example, Q16 [60] is a widely used  [60, 62, 59, 44, 73, 35, 28, 66, 27] binary image classifier that predicts a given image as morally positive or negative.
+After extracting image features with CLIP, it compares the image feature with two soft prompts (optimized text embeddings), one representing the positive class and the other being the negative class.
+In this paper, we examine five popular image safety classifiers: Q16 [60], MultiHeaded [53], SD_Filter [55], NSFW_Detector [14], and NudeNet [15].
+We summarize the backbone, training paradigm, and training dataset in Table 1 and elaborate on these details in Section A.1 in the Appendix.
+Large Visual Language Models.
+Large visual language models have achieved extraordinary capabilities in understanding visual and text content.
+Given an image and a text instruction, these models can read the image and generate responses following the instruction.
+Recent studies [38, 56] show that VLMs can be used to detect user-generated unsafe images [38] and hateful memes [56].
+We test both commercial and open-source VLMs including GPT-4V [30], LLaVA (7B) [45], and InstructBLIP (7B) [30].
+The specific checkpoints we used are provided in Section A.2 in the Appendix.
+Overview of UnsafeBench
+We build UnsafeBench, a benchmarking framework that comprehensively evaluates the performance of image safety classifiers.
+We show an overview of UnsafeBench in Figure 1.
+Due to the absence of a comprehensive labeled dataset in the image safety domain, we first construct the UnsafeBench dataset, which contains 10K images with human annotations.
+We then collect the existing image safety classifiers, as conventional classifiers, and three VLMs capable of classifying unsafe images, as VLM-based classifiers.
+We identify the range of unsafe content covered by these classifiers and align them with our unsafe image taxonomy, i.e., 11 unsafe categories.
+Finally, we evaluate the effectiveness and robustness of these classifiers; effectiveness measures how accurately the classifiers can identify unsafe images, while robustness reflects their ability to maintain accuracy against perturbations.
+Figure 1: Overview of UnsafeBench.
+Dataset Construction
+Image Collection.
+We rely on open large-scale multimodal datasets, LAION-5B [62] and Lexica [9], to collect real-world and AI-generated unsafe images.
+LAION-5B is currently the largest public image-text dataset in the world [7], containing 5.85 billion image-text pairs collected from webpages [62].
+We regard LAION-5B as the source to collect real-world images.
+Lexica [9] is an AI-generated image gallery that allows users to create and disclose images generated by models like Stable Diffusion [57].
+Both datasets serve as a search engine and enable users to find the most relevant images based on a textual description.
+Inspired by this, we use unsafe keywords to query these datasets and collect potentially unsafe images.
+To collect unsafe keywords, we split the definitions of 11 unsafe categories (see Section 2) as initial unsafe keywords.
+Additionally, we employ a large language model, Vicuna-33b, [^3] https://huggingface.co/lmsys/vicuna-33b-v1.3.
+to augment the definition of each unsafe category and provide 10 specific examples.
+We show the generated examples in Table 8 in the Appendix.
+Combining the initial unsafe keywords and those generated with Vicuna-33b, we obtain 158 unsafe keywords, covering 11 unsafe categories (each containing 12-16 keywords).
+We then query LAION-5B and Lexica using these keywords and collect the most relevant images.
+After removing duplicates, we collect a total of 12,932 potentially unsafe images, comprising 5,815 images from LAION-5B and 7,117 from Lexica.
+Image Annotation.
+We perform a human annotation to determine if these collected images are truly unsafe.
+Three authors of this paper serve as annotators and manually annotate them independently.
+We require these annotators to first read the definition from each unsafe category as the criterion in determining whether an image is safe, unsafe, or not applicable (N/A).
+Note that N/A mainly represents noise images that cannot be identified as safe or unsafe, e.g., a blurry image or one with unidentifiable texts.
+The annotation process undergoes two rounds.
+In the first round, two annotators independently assign a label of safe, unsafe, or N/A to each image.
+For images where two annotators disagree, we further introduce the third annotator to provide additional labels.
+The final label for each image is then determined based on the majority vote among these labels.
+In the second round, annotators revisit images annotated as unsafe and determine the unsafe category.
+If an image displays a mix of unsafe elements, we choose the predominant category it violates.
+Through two rounds of annotation, we ensure that each image is labeled as safe, unsafe, or N/A, and if it is unsafe, it can be classified into a specific unsafe category.
+To evaluate the reliability of our annotation, we calculate the Fleiss’ Kappa score [33, 32], a statistical measure for assessing the agreement among annotators.
+The score is 0.47, indicating a moderate level of agreement among annotators [33, 32].
+Table 2: Statistics of the annotated images.
+Dataset # All # Safe # Unsafe # N/A Fleiss’ Kappa
+LAION-5B 5,815 3,228 1,832 755 0.49
+Lexica 7,117 2,870 2,216 2,031 0.46
+All 12,932 6,098 4,048 2,786 0.47
+Table 3: Aligning the unsafe content covered by image safety classifiers with 11 unsafe categories.
+Classifier Hate Harassment Violence Self-Harm Sexual Shocking Illegal Activity Deception Political Health Spam
+Conventional Classifiers Q16
+MultiHeaded
+SD_Filter
+NSFW_Detector
+NudeNet
+VLM-Based Classifiers LLaVA
+InstructBLIP
+GPT-4V
+Dataset Statistics.
+We show the statistics of our annotated images in Table 2 and Table 9.
+Overall, we annotate 12,932 images, including 6,098 safe images, 4,048 unsafe images, and 2,786 N/A images.
+For the following experiments, we exclude the N/A images and use the remaining images (safe + unsafe), collectively referred to as the UnsafeBench dataset.
+To the best of our knowledge, it is currently the most comprehensive annotated dataset in the image safety domain.
+Visualizing Distribution Shift.
+We compare the difference between real-world and AI-generated images in image semantics and qualities, such as signal-to-noise ratio (SNR), edge density, and entropy.
+Figure 11 and Figure 13 show the different distribution of two groups of images.
+We find that a distribution shift is present in image semantics, e.g., image embeddings are separated in the 2D space between real-world and AI-generated images.
+The shift also exists in image qualities, e.g., AI-generated images have lower SNR and higher entropy, indicating they tend to be noisier and more complex.
+Image Classifier Collection
+Conventional Classifiers.
+We collect five conventional classifiers: Q16 [60], MultiHeaded [53], SD_Filter [55], NSFW_Detector [14], and NudeNet [15], as introduced in Section 2.
+These classifiers take images as input and output predictions on whether the images are considered unsafe based on their defined scope of unsafe content.
+VLM-Based Classifiers.
+We utilize three VLMs, LLaVA, InstructBLIP, and GPT-4V to build our VLM-based classifiers.
+Unlike conventional classifiers, these models take both the image and prompt as input to identify if the given image contains unsafe content.
+In our study, we manually design three prompts that guide these models to classify whether a given image is safe or not.
+We show three prompts in Table 10 in the Appendix.
+The first prompt is straightforward, “classify the image as safe or unsafe.” It relies on models’ internal criteria of unsafe content to determine if an image is safe.
+To align with our safety standards, the second prompt further incorporates definitions of the taxonomy, similar to [38].
+For example, to classify images from the Hate category, the prompt is “An image is considered unsafe if it contains any of the following unsafe content: hateful symbols, negative stereotypes…classify the image as safe or unsafe.” In case models are still affected by the internal criteria, we further add a special requirement in the third prompt, “Forget about your definition of unsafe images and adhere to this new one…” For each image, we query VLMs with three prompts and collect the respective outputs.
+Given that VLMs’ outputs are sometimes lengthy sentences, they do not directly categorize into safe or unsafe classes.
+To convert these VLMs to classifiers, we fine-tune a language model, RoBERTa [46], to categorize LLM’s outputs into one of three classes: safe, unsafe, or uncertain.
+The uncertain class describes a set of outputs that do not yield a clear prediction.
+For example, InstructBLIP occasionally describes the content of an image without making a prediction.
+Note, we categorize the cases where VLMs refuse to answer as the unsafe class, as input images have triggered their internal safeguards.
+We collect several example outputs [^4] For example, we use “safe,” “the image is safe,” “the image is not considered unsafe as…” as example outputs for the safe class.
+for each class and fine-tune RoBERTa in a supervised manner.
+Ultimately, a VLM-based classifier consists of a VLM and the RoBERTa classifier, which function sequentially.
+Aligning Classifier Coverage With Unsafe Categories
+Image safety classifiers are designed to target various types of unsafe content.
+To align the unsafe content covered by a classifier with the taxonomy, i.e., 11 unsafe categories, we first identify the range of unsafe content covered by each classifier by examining its documentation and training data.
+Then, if any unsafe category from our taxonomy falls into this range, we assign the unsafe category to the classifier.
+For example, Q16 [60] detects “morally negative” images.
+We examine the definition of “morally negative” content in [29] and find that it covers a list of unsafe concepts, such as harm, inequality, degradation, etc.
+By mapping these unsafe concepts to our taxonomy, we align the ambiguous definition of “morally negative” with specific unsafe categories.
+For VLMs, as we adopt the unsafe image taxonomy from OpenAI, we assume that GPT-4V covers all 11 categories.
+We also infer that LLaVA and InstructBLIP are capable of handling these categories since they are fine-tuned on the instruction dataset generated by GPT-4V [45, 30].
+We show the aligning result in Table 3.
+Evaluation Methodology
+Effectiveness.
+We evaluate the effectiveness of eight image safety classifiers (five conventional classifiers and three VLM-based classifiers).
+For conventional classifiers, we feed images from each unsafe category and obtain binary predictions (safe/unsafe).
+Regarding VLM-based classifiers, we input both images with the designed prompts to gather outputs, which are then classified into safe/unsafe/uncertain classes by the fine-tuned RoBERTa.
+Uncertain predictions account for about 0.1%.
+However, since conventional classifiers directly predict images as safe or unsafe, to ensure a fair comparison, we randomly assign a prediction of safe or unsafe to VLM’s uncertain predictions.
+Recall that we design three prompts for VLMs, resulting in three separate predictions for each image.
+The final prediction is determined by a majority vote.
+Evaluation Metric for Effectiveness.
+With the predictions in place, we calculate the F1-Score to evaluate the effectiveness of image safety classifiers.
+We choose F1-Score as it accounts for both false positives and false negatives, providing a balanced measure of the classifier’s Precision and Recall.
+Robustness.
+We evaluate the robustness of classifiers using images perturbed with both random noise (Gaussian noise) and adversarial noise (i.e., adversarial examples) [49, 47].
+Adversarial examples ($x_{adv}$) are original inputs ($x$) with optimized perturbations ($\Delta x$), which maximizes the loss ($L(\theta,x+\Delta x,y)$) of a model ($\theta$) and fools the model into making incorrect predictions.
+The p-norm of the perturbation vector is limited by a small $\epsilon$ to ensure the perturbation is imperceptible.
+$\begin{gathered}x_{adv}:=\underset{\Delta x}{\text{argmax}}\,L(\theta,x+\Delta x,y),\\ \quad\left\|\Delta x\right\|_{p}<\epsilon.\end{gathered}$ (1)
+We create adversarial examples using three gradient-based adversarial algorithms, FGSM [37], PGD [47], and DeepFool [50], with details introduced in Section A.3 in the Appendix.
+These algorithms optimize perturbations using gradients of the classifiers’ loss with respect to the input image.
+For conventional classifiers, we directly use three types of adversarial algorithms to solve the Equation 1 and obtain the optimized perturbations.
+However, for VLMs, directly solving Equation 1 does not necessarily create adversarial examples that lead to opposite predictions.
+Unlike binary classifiers, drifting the VLM away from its original predictions, e.g., “the image is safe,” could result in unexpected outputs that are unrelated to the image safety classification task.
+To address this, we transform the untargeted attack in Equation 1 into targeted attacks (targeting the opposite class), which are equivalent in binary classification settings.
+For example, if the VLM initially classifies an image as safe, we then optimize the perturbation such that the output moves toward the “unsafe” direction by setting the target output as “unsafe.” Therefore, instead of maximizing the loss between the classifier’s prediction and the original label ($y$), here, we minimize the loss of a VLM between its prediction and the defined target output ($y_{tar}$), as shown in Equation 2.
+By solving the equation, we update optimized perturbations until the RoBERTa classifier classifies the VLM output as the opposite class from the image label.
+Using this strategy, we create adversarial examples using FGSM, PGD, and DeepFool on open-source VLM-based classifiers (LLaVA and InstructBLIP).
+Note that we exclude GPT-4V from this evaluation because crafting adversarial examples requires model gradients, which is not available for GPT-4V.
+$\begin{gathered}x_{adv}:=\underset{\Delta x}{\text{argmin}}\,L(\theta,x+\Delta x,y_{tar}),\\ \quad\left\|\Delta x\right\|_{p}<\epsilon.\end{gathered}$ (2)
+To maintain the same perturbation budget for both conventional classifiers and VLMs, we use the $L_{\infty}$ norm, set $\epsilon$ to 0.01, and also limit the number of optimization iterations to a maximum of 100.
+We demonstrate the perturbed images and different types of perturbations in Figure 2.
+Figure 2: Perturbed images with Gaussian perturbations and three types of adversarial perturbations bounded by $L_{\infty}$ norm ($\epsilon=0.01$).
+Evaluation Metric for Robustness.
+We calculate the Robust Accuracy (RA) to evaluate the robustness of image safety classifiers.
+RA is the percentage of perturbed images that have been correctly predicted by classifiers out of all perturbed images, which is also equal to 1 - attack success rate.
+Here, we generate adversarial examples only for images that are correctly predicted by classifiers during the effectiveness evaluation.
+Evaluation Results
+Table 4: F1-Score of eight image safety classifiers on the UnsafeBench dataset. * marks VLM-based classifiers.
+Dataset Classifier Hate Harassment Violence Self-Harm Sexual Shocking Illegal Activity Deception Political Health Spam Overall
+LAION-5B (Real-World) Q16 0.641 0.517 0.693 0.421 - 0.630 0.681 0.762 0.271 0.144 - -
+MultiHeaded 0.320 - 0.247 0.000 0.692 0.644 - - 0.209 - - -
+SD_Filter - - - - 0.833 - - - - - - -
+NSFW_Detector - 0.517 - - 0.783 - - - - - - -
+NudeNet - - - - 0.650 - - - - - - -
+LLaVA* 0.227 0.167 0.413 0.451 0.674 0.714 0.498 0.376 0.116 0.333 0.059 0.426
+InstructBLIP* 0.351 0.394 0.606 0.275 0.796 0.467 0.722 0.653 0.444 0.379 0.380 0.520
+GPT-4V* 0.556 0.706 0.774 0.557 0.866 0.724 0.897 0.827 0.605 0.405 0.718 0.712
+Lexica (AI-Generated) Q16 0.336 0.416 0.612 0.521 - 0.836 0.497 0.384 0.597 0.615 - -
+MultiHeaded 0.225 - 0.533 - 0.815 0.780 - - 0.744 - - -
+SD_Filter - - - - 0.727 - - - - - - -
+NSFW_Detector - 0.259 - - 0.678 - - - - - - -
+NudeNet - - - - 0.596 - - - - - - -
+LLaVA* 0.169 0.224 0.632 0.663 0.866 0.826 0.534 0.179 0.340 0.665 0.045 0.634
+InstructBLIP* 0.178 0.332 0.629 0.383 0.757 0.795 0.665 0.302 0.785 0.663 0.519 0.589
+GPT-4V* 0.254 0.635 0.712 0.613 0.827 0.875 0.701 0.422 0.909 0.621 0.171 0.707
+Overall Q16 0.533 0.475 0.648 0.483 - 0.784 0.571 0.652 0.482 0.498 - -
+MultiHeaded 0.292 - 0.426 - 0.757 0.749 - - 0.600 - - -
+SD_Filter - - - - 0.785 - - - - - - -
+NSFW_Detector - 0.449 - - 0.738 - - - - - - -
+NudeNet - - - - 0.624 - - - - - - -
+LLaVA* 0.210 0.189 0.550 0.590 0.780 0.800 0.519 0.332 0.266 0.575 0.054 0.549
+InstructBLIP* 0.270 0.368 0.615 0.333 0.777 0.697 0.687 0.506 0.660 0.545 0.490 0.559
+GPT-4V* 0.423 0.681 0.738 0.590 0.847 0.839 0.780 0.673 0.780 0.492 0.537 0.709
+Effectiveness
+Table 4 shows the effectiveness of eight image classifiers on the UnsafeBench dataset.
+This table is consistent with the aligning result shown in Table 3.
+If a classifier cannot identify a specific unsafe category according to Table 3, we fill the corresponding cell with a “-.”
+Conventional vs.
+VLM-Based Classifiers.
+The top-performing image safety classifier is the commercial VLM-based classifier, GPT-4V.
+It achieves the highest overall F1-Score, surpassing LLaVA and InstructBLIP, as well as other conventional classifiers in most unsafe categories, except for Hate and Deception.
+GPT-4V shows exceptional effectiveness in detecting Sexual (0.847), Shocking (0.839), Political (0.780), Illegal Activity (0.780), and Violence content (0.738).
+Among conventional classifiers, Q16 stands out by identifying the broadest spectrum of unsafe content, with an F1-Score from 0.475 to 0.784.
+However, Q16 does not support the detection of Sexual content and requires improvements in Harassment, Self-Harm, and Political categories.
+In contrast, MultiHeaded, SD_Filter, NSFW_Detector, and NudeNet can detect sexual images with an F1-Score of 0.624-0.785.
+Although GPT-4V shows superior performance on our annotated dataset, the wide application is constrained by its commercial nature, mainly due to the financial cost and slow inference speed.
+It costs approximately \$20 and 50 minutes to classify 1K images.
+Additionally, we observe that zero-shot VLMs’ performances significantly rely on prompt designing.
+For example, GPT-4V’s overall F1-Score achieves 0.68 and 0.71 when we classify images with the second and third prompts (shown in Table 10 in the Appendix).
+However, the score drops to 0.61 on the first prompt.
+To obtain reliable results from VLMs like GPT-4V, we typically query the VLM with the same image using multiple prompts and determine the label based on a majority vote.
+This approach further increases the cost, making GPT-4V impractical for application on large-scale image datasets.
+Current research [59, 35, 48, 73] prefers smaller classifiers over larger VLMs due to their faster inference speed.
+A common practice is combining Q16 and a classifier that detects sexually explicit content like NudeNet to detect generally unsafe images [59, 35, 48, 73], e.g., regarding the image is unsafe if either Q16 or NudeNet classifies it as unsafe.
+According to Table 4, this strategy can indeed cover many unsafe categories with a relatively good performance.
+We calculate the overall F1-Score of Q16 combined with NudeNet across the unsafe categories they can cover, and the score is 0.665.
+This performance needs to be further enhanced.
+Imbalanced Effectiveness Across Different Categories.
+To understand which unsafe categories are more effectively detected, we count the number of classifiers capable of identifying each category and record the average F1-Score in Figure 3.
+We find that the Sexual and Shocking categories have higher average F1-Scores, close to 0.8, compared to other unsafe categories.
+They are also covered by more (5-7) classifiers.
+In contrast, the remaining unsafe categories, particularly Hate, Harassment, and Self-Harm, have lower average F1-Scores (below 0.5).
+This discrepancy highlights the imbalanced effectiveness across different types of unsafe content.
+Among 11 unsafe categories, we particularly focus on the Hate category, given that hateful images can proliferate across Web communities as memes and are often used in coordinated hate campaigns [72, 36].
+For instance, the anti-Semitic symbol, Happy Merchant [5], widely spreads on platforms like 4chan and mainly targets the Jewish community.
+Due to the harmful nature, OpenAI ranks the Hate category as the highest priority in its content policy [17].
+However, among the five classifiers that can detect hateful content, the highest F1-Score is 0.533, achieved by Q16.
+The second highest is 0.423, achieved by GPT-4V.
+To understand the underlying reason behind such low F1-Scores, we zoom into the misclassified examples from the Hate category for both Q16 and GPT-4V.
+We calculate the false positive and false negative rates for both classifiers.
+For Q16, the false positive rate is 0.166 and the false negative rate is 0.306.
+In comparison, for GPT-4V, the false positive rate is 0.425, and the false negative rate is 0.144.
+This result suggests that both classifiers have a considerable likelihood to misclassify hateful images: either failing to identify truly hateful images or incorrectly predicting safe images as hateful.
+For model providers and platform moderators, the false negative rate may be more important as it indicates the likelihood of hateful images evading the detection of models’ safeguards.
+In our manual review of common false negatives, we observe that 58 images containing hateful content are misclassified as safe by both Q16 and GPT-4V.
+Notably, some Neo-Nazi and anti-Semitic symbols, such as the tattooed swastika [13] and the Happy Merchant meme [5], manage to evade detection.
+We show these examples in Figure 12 in the Appendix.
+Figure 3: Average F1-Score and number of classifiers for each unsafe category.
+Figure 4: Average F1-Score of classifiers on real-world and AI-generated images.
+(a) False Negatives (Misclassify Unsafe as Safe)
+(b) False Positives (Misclassify Safe as Unsafe)
+Figure 5: Image clusters from the Sexual category that are misclassified by SD_Filter, NSFW_Detector, and NudeNet.
+We annotate each central image with its cluster ID and cluster size.
+We blur sexual images for censoring purposes.
+Real-World vs.
+AI-Generated.
+Real-world images from LAION-5B and AI-generated ones from Lexica have different distributions.
+To compare the performance of tested classifiers on two groups of images, we calculate the average F1-Score of each classifier on real-world and AI-generated images in Figure 4.
+We find that SD_Filter, NSFW_Detector, NudeNet, and GPT-4V have better performance in identifying real-world unsafe images compared to AI-generated ones.
+For example, the F1-Score of SD_Filter decreases from 0.833 on real-world images to 0.727 on AI-generated images.
+Meanwhile, MultiHeaded, LLaVA, and InstructBLIP have higher F1-Scores on AI-generated images than real-world ones.
+The different behaviors of classifiers on two sets of images are multifaceted.
+A key factor is the distribution shift between training and testing images, including nuanced differences in semantics, image styles, etc.
+Recall the training images of the classifiers provided in Table 1, classifiers like NudeNet and NSFW_Detector are trained on real-world NSFW images [16], as indicated by [15] and [14].
+Moreover, MultiHeaded is trained only on unsafe images by text-to-image models, thus demonstrating better performance on AI-generated images.
+Effectiveness Degradation of Classifiers Trained on Real-World Images Only.
+Next, we explore in detail why certain classifiers experience degraded performance on AI-generated images, and what characteristics of AI-generated images contribute to it.
+We select two groups of classifiers that exhibit degraded performances on AI-generated images compared to real-world images.
+The first group consists of SD_Filter, NFSW_Detector, and NudeNet.
+These classifiers achieve F1-Scores ranging from 0.650 to 0.833 in categorizing real-world images within the Sexual category.
+However, the scores drop to 0.596-0.727 when applied to AI-generated images based on Table 4.
+The second group, including Q16 and GPT-4V, recognizes real-world violent images with F1-Scores of 0.693 to 0.774 but drops to 0.612 to 0.712 for AI-generated violent images.
+To investigate the reason behind this performance degradation on AI-generated images, we again examine the misclassified examples, including false negatives and false positives from both LAION-5B and Lexica.
+To characterize images from two sources, we perform KMeans clustering and group these misclassified examples into $K$ clusters, with images from the same cluster sharing similar semantics.
+Specifically, we utilized the CLIP image encoder to generate image embeddings that capture semantic information.
+We then apply KMeans clustering with $K=4$ for better visualization purposes.
+Finally, we create four clusters for each group of images and retrieve the nearest image to each cluster centroid.
+We display these central images, representing their respective clusters in Figure 5 and Figure 14 in the Appendix, each figure corresponding to one group of classifiers.
+(a) Strengths of Artistic Representation
+(b) Grid Sizes
+Figure 6: The original real-world image and its AI-generated variations applying the artistic style and grid layout.
+The original image is unsafe and correctly predicted by Q16.
+Text in red indicates that image variations are misclassified as safe.
+Figure 7: Percentage of correct predictions by NudeNet, NSFW_Detector, and Q16 for images with various strengths of artistic representation and grid sizes.
+Distinct Characteristics in AI-Generated Images.
+Comparing the misclassified images between real-world and AI-generated ones, we observe two distinct characteristics prevalent in AI-generated images: artistic representation and grid layout.
+First, images in an artistic representation, [^5] It is an open question whether unsafe content in artistic representation remains unsafe.
+In this study, we adhere to the criteria from the unsafe image taxonomy and evaluate image safety based on its content and intention, regardless of the representation form.
+despite showing explicitly unsafe content, can often escape detection of classifiers.
+For example, among false negatives shown in 5(a), AI-generated images within clusters 2 and 4 show nudity in an artistic style, yet they are misclassified as safe.
+In contrast, real-world images depicting realistic nudity are rarely misclassified as safe, given that they also widely exist in our annotated dataset.
+Similarly, among the misclassified images from the Violence category (see 14(a) in the Appendix), AI-generated images within clusters 1, 2, and 3 are also misclassified, which portray artistic representations of violence, e.g., fighting scenes.
+Second, images in a grid layout from AI-generated ones are frequently misclassified by classifiers.
+These images are composites made of smaller images, commonly four or nine, arranged in a grid layout.
+Examples include cluster 1, 3 in 5(a), and cluster 3 in 14(a).
+Based on these findings, we hypothesize that these distinctive characteristics of AI-generated images, like artistic representation and grid layout, contribute to the reduced effectiveness, particularly for models trained on real-world images, e.g., NSFW_Detector, NudeNet, and Q16.
+Case Study on Artistic Representation and Grid Layout.
+To test our hypothesis, we conduct a case study where we generate image variations with different levels of artistic representation and grid sizes to observe if classifiers can consistently make accurate predictions.
+To this end, we randomly sample 20 realistic unsafe images from LAION-5B in both Sexual and Violence categories.
+These images are accurately identified as either sexual content by NSFW_Detector and NudeNet or as violent content by Q16, and these three classifiers have been verified to be trained on real-world images only.
+For each unsafe image, we then create artistic versions using the shape-preserving image editing function of a text-to-image model, Stable Diffusion [57].
+Specifically, we input the image into the model and use the prompt “in the style of oil painting” to generate artistic images.
+By adjusting the strength parameter, we control the degree of the artistic style applied to the original image.
+For each image, we produce six artistic versions with the strength parameter ranging from 0.1 to 0.6.
+Additionally, we generate image variations with different grid sizes by reconstructing an image with $2\times 2$, $3\times 3$, and $4\times 4$ grids, with each smaller image serving as a part of the grid.
+Figure 6 shows image variations with an example from the Violence category.
+In total, we collect 40 original images and their 360 variations, including 240 ($40\times 6$) artistic and 120 ($40\times 3$) grid variations.
+To guarantee that the variations in the artistic representation retain the same content as the original images, we conduct a manual annotation to filter out any variations where the content has been altered.
+We use the remaining 17 original images and their 153 variations to test NSFW_Detector, NudeNet, and Q16.
+We present the percentage of correct predictions in Figure 7.
+The result indicates that with the increasing strength of artistic representation and grid sizes, the classifiers’ effectiveness generally decreases.
+This confirms the impact of these characteristics found in AI-generated images on the performance of classifiers trained on real-world images only.
+Table 5: Robustness of image safety classifier against Gaussian perturbation and three types of adversarial perturbations.
+We report the mean robust accuracy (RA) and the standard deviation based on results from three sampling times.
+All perturbations are bounded by a fixed $\epsilon=0.01$ and obtained within the same number of iterations (1 for GN/FGSM; 100 for PGD/DeepFool).
+Dataset Model GN FGSM PGD DeepFool Average
+LAION-5B Q16 1.000 $\pm$0.000 0.497 $\pm$0.003 0.002 $\pm$0.002 0.170 $\pm$0.011 0.417 $\pm$0.002
+MultiHeaded 1.000 $\pm$0.000 0.707 $\pm$0.025 0.001 $\pm$0.002 0.321 $\pm$0.040 0.507 $\pm$0.016
+SD_Filter 1.000 $\pm$0.000 0.603 $\pm$0.001 $0$$\pm$0.000 0.047 $\pm$0.011 0.413 $\pm$0.003
+NSFW_Detector 0.999 $\pm$0.001 0.603 $\pm$0.005 0.001 $\pm$0.001 0.423 $\pm$0.020 0.507 $\pm$0.006
+NudeNet 1.000 $\pm$0.000 0.105 $\pm$0.013 $0$$\pm$0.000 0.092 $\pm$0.009 0.299 $\pm$0.006
+LLaVA* 0.945 $\pm$0.011 0.784 $\pm$0.020 0.712 $\pm$0.005 0.458 $\pm$0.010 0.725 $\pm$0.006
+InstructBLIP* 0.995 $\pm$0.002 0.621 $\pm$0.003 0.472 $\pm$0.013 0.163 $\pm$0.006 0.563 $\pm$0.006
+Lexica Q16 0.999 $\pm$0.001 0.313 $\pm$0.016 0.001 $\pm$0.001 0.115 $\pm$0.016 0.357 $\pm$0.008
+MultiHeaded 0.999 $\pm$0.001 0.474 $\pm$0.008 0.001 $\pm$0.001 0.157 $\pm$0.004 0.408 $\pm$0.001
+SD_Filter 0.999 $\pm$0.001 0.353 $\pm$0.005 $0$$\pm$0.000 0.069 $\pm$0.005 0.355 $\pm$0.001
+NSFW_Detector 0.999 $\pm$0.001 0.451 $\pm$0.007 $0$$\pm$0.000 0.325 $\pm$0.009 0.444 $\pm$0.004
+NudeNet 1.000 $\pm$0.000 0.076 $\pm$0.000 0.003 $\pm$0.000 0.082 $\pm$0.000 0.290 $\pm$0.000
+LLaVA* 0.959 $\pm$0.016 0.648 $\pm$0.009 0.587 $\pm$0.015 0.469 $\pm$0.023 0.666 $\pm$0.010
+InstructBLIP* 0.999 $\pm$0.001 0.648 $\pm$0.017 0.565 $\pm$0.001 0.324 $\pm$0.015 0.634 $\pm$0.007
+Take-Aways.
+We assess the effectiveness of eight image safety classifiers on real-world and AI-generated images across 11 unsafe categories.
+Our findings reveal several insights.
+First, of all the classifiers evaluated, the commercial model GPT-4V stands out as the most effective in identifying a broad spectrum of unsafe content.
+However, due to its associated time and financial costs, it is impractical to moderate large-scale image datasets.
+Currently, there is a lack of an open-source image safety classifier that can comprehensively and effectively identify unsafe content.
+Second, the effectiveness varies significantly across 11 unsafe categories.
+Images from the Sexual and Shocking categories are detected more effectively, while categories such as Hate require further improvement.
+For instance, we observe that Neo-Nazi and anti-Semitic symbols can sometimes evade detection by both Q16 and GPT-4V.
+Finally, we find certain classifiers trained on real-world images experience performance degradation on AI-generated images.
+Notably, AI-generated images often exhibit unique characteristics, such as artistic representation of unsafe content and unsafe images in a grid layout.
+Robustness
+Results.
+We randomly sample 500 images three times that are correctly classified by each classifier and create adversarial examples.
+Table 5 lists the RA of seven open-source classifiers for four types of perturbations.
+VLM-based classifiers show the highest robustness.
+They achieve RAs between 0.563-0.725 on LAION-5B images and 0.634-0.666 on Lexica images, higher than any conventional classifiers.
+Meanwhile, conventional classifiers are more vulnerable to adversarial attacks, with RAs of 0.299-0.507 on LAION-5B images and 0.290-0.444 on Lexica images.
+Among them, NudeNet shows the lowest RA (0.290-0.299), revealing that it is the most vulnerable classifier to adversarial attacks.
+Relying on Pre-Trained Foundational Models Enhances Robustness Than Training Smaller Classifiers From Scratch.
+We review the model architecture, training paradigm, and training dataset for each classifier (see Section 2).
+As the least robust classifier, NudeNet is an Xception-based classifier and is trained on 160K labeled images using fully supervised learning.
+Other conventional classifiers are built on CLIP, which is pre-trained on 400 million image-text pairs and then fine-tuned on their own labeled datasets using linear probing or prompt learning.
+They present higher robustness compared to training a small classifier using supervised learning, i.e., NudeNet.
+VLM-based classifiers present the highest robustness and also rely on large pre-trained models.
+For example, LLaVA utilizes CLIP as the image feature extractor and LLaMA [65] as the LLM to reason over the input image and generate output.
+For developing future classifiers, adapting large pre-trained foundation models can yield a more resilient system.
+Figure 8: Confidence scores of classifiers’ predictions over different groups of images.
+The median value is denoted next to the dashed line.
+AI-generated images tend to have lower confidence scores than real-world images.
+Classifiers Are More Susceptible to Adversarial Attacks using AI-Generated Images Than Real-World Images.
+We find that most classifiers, except for InstructBLIP, exhibit lower RAs on AI-generated images compared to their real-world counterparts.
+Their average RAs show a consistent decrease from the range of 0.299–0.725 to 0.290–0.666, with a maximum drop of 10%.
+To investigate the reason behind this, we calculate the maximum probability as the confidence score of four conventional classifiers when classifying different groups of images.
+As shown in Figure 8, successful adversarial examples (denoted in blue) generally have lower confidence scores.
+This indicates that these examples are closer to the decision boundary and can be more easily perturbed to cross over this boundary compared to non-adversarial examples.
+When comparing AI-generated and real-world images, AI-generated images tend to have lower confidence scores in the distribution.
+This potentially contributes to the higher robustness susceptibility for these classifiers.
+Interestingly, we also find these classifiers only show evident RA decreases on adversarial perturbations (especially FGSM), and not with Gaussian perturbations.
+The difference in creating the two types of perturbations is that adversarial perturbations are designed to maximize the training loss, i.e., cross-entropy loss, whereas Gaussian perturbations are not designed to do so.
+We then analyze classifiers’ cross-entropy loss and their loss change due to the addition of adversarial perturbations, using FGSM as an example.
+In Figure 15 (in the Appendix), we observe a general trend of higher loss values and loss increase in AI-generated images than real-world ones.
+This implies that classifiers tend to be more sensitive to the perturbations in AI-generated images, leading to a higher loss increase even with the same amount of perturbation.
+This also explains why, even if the median confidence scores and loss values are close between AI-generated images and real-world images, such as the NSFW_Detector, the AI image group still makes the classifier more vulnerable to adversarial attacks, because they are more prone to crossing the decision boundary.
+PerspectiveVision
+Motivation & Overview
+The findings from UnsafeBench point out two shortcomings in open-source image safety classifiers, including zero-shot VLMs.
+First, there is a lack of an open-source image moderation tool that effectively identifies the wide spectrum of unsafe images, particularly in several neglected categories such as hateful content and self-harm.
+Second, these classifiers show both degraded effectiveness and robustness in detecting AI-generated unsafe images due to distributions that differ from real-world images.
+With AI-generated content becoming increasingly common in our daily lives, there is an urgent need to develop a comprehensive image moderation tool that accounts for these AI products.
+The UnsafeBench dataset serves as a good starting point, as it covers 11 unsafe categories and AI-generated content.
+Using this dataset, we aim to build a comprehensive moderation tool, PerspectiveVision, that can effectively identify a wide range of unsafe images from the real world and AI.
+PerspectiveVision is inspired by Google’s Perspective API, a benchmark tool for detecting toxic text in the NLP domain.
+The API assesses text toxicity across various attributes, such as severe toxicity, hate, and profanity.
+Similarly, PerspectiveVision is designed to evaluate the safety of a given image and further identify the specific category if the image is deemed unsafe.
+Users can customize the scope of unsafe images by selecting multiple categories to suit different definitions of unsafe content.
+The high-level overview is demonstrated in Figure 9.
+Figure 9: High-level overview of PerspectiveVision.
+Table 6: F1-Score of PerspectiveVision models on external evaluation datasets. We use GPT-4V as a baseline for comparison.
+In-Distribution Out-of-Distribution (OOD)
+Classifier UnsafeBench-Test SMID MultiHeaded-D NudeNet-D Self-Harm Violent Behaviors OOD Average
+Linear Probing (CLIP) 0.859 0.365 0.654 0.954 0.801 0.740 0.752
+Prompt Learning (CLIP) 0.687 0.623 0.500 0.984 0.971 0.931 0.802
+LoRA Fine-Tuning (LLaVA) 0.844 0.557 0.675 0.986 0.948 0.908 0.836
+GPT-4V 0.717 0.532 0.577 0.991 0.924 0.921 0.830
+Methodology
+Figure 10: An example of LLaVA fine-tuning dataset.
+The main idea is to build a unified model with VLMs like LLaVA instead of training classifiers separately for different categories.
+To this end, we fine-tune a VLM, i.e., LLaVA, to identify unsafe images by generating a response that indicates both the safety status and the specific unsafe category.
+Since fine-tuning the entire LLaVA is computationally intensive, we use Low-Rank Adaptation (LoRA) [40] while training.
+Training Dataset Construction.
+The training dataset is based on our annotated images.
+The dataset contains triplet elements: image, prompt, and target output.
+We adopt a similar prompt template as LLaMA-Guard [41] (an LLM-based safeguard), where we define the unsafe image taxonomy with specific unsafe categories and output format.
+Figure 10 demonstrates an example of the LLaVA fine-tuning dataset.
+To reduce overfitting, for each annotated image, we randomly remove K irrelevant categories and shuffle the order of the rest categories to build the prompt.
+Training Dataset Augmentation.
+To further improve generalizability, we augment the dataset by label flipping and class balancing.
+On top of the basic training dataset, for images that are annotated as unsafe, we randomly sample K irrelevant categories to serve as the unsafe image taxonomy in the prompt.
+We ensure that these images are not unsafe within these categories and then flip the target output to the safe class.
+This helps the model learn the association between user-defined taxonomy in the prompt and the image label.
+Finally, we balance the number of samples between two classes by supplementing the smaller class.
+Baselines
+We employ two common techniques as baselines: linear probing and prompt learning to adapt CLIP to identify unsafe images.
+Linear probing [54] adds a linear probing layer on top of CLIP, allowing for fine-tuning on specific tasks, and prompt learning [31, 39] refers to a process that adapts CLIP for various downstream tasks by optimizing prompts.
+For linear probing, we train 11 linear probing layers, each one identifying images from a specific unsafe category.
+Similarly, we optimize 11 sets of soft prompts using prompt learning.
+PerspectiveVision Evaluation
+Evaluation Datasets.
+We randomly split the UnsafeBench dataset into an 80% training and 20% testing ratio and take the test split as in-distribution evaluation dataset.
+To assess the model’s generalizability, we further collect multiple external datasets mainly from the training data of conventional classifiers.
+We regard them as out-of-distribution datasets: SMID [22], MultiHeaded Dataset [12], NudeNet Dataset [16], Self-Harm Dataset [20, 21], and Violent Behavior Dataset [25].
+We outline these datasets and detailed experimental setup in Section A.4 and Section A.5 in the Appendix.
+Effectiveness Evaluation.
+Table 6 presents the evaluation results of the PerspectiveVision models across UnsafeBench test set and five OOD datasets.
+Here, we benchmark GPT-4V against these datasets to establish a high-performing baseline.
+Among the three methods employed, linear probing (CLIP) and LoRA fine-tuning LLaVA demonstrate a higher fitting level on the UnsafeBench test set, with an overall F1-Score of 0.859 and 0.844.
+Meanwhile, on external datasets, LoRA fine-tuned LLaVA achieves the highest average F1-Score, 0.836.
+Although closely followed by GPT-4V, which has an F1-Score of 0.830, our model is notable for its open-source nature, large-scale detection, and fine-grained classification of different unsafe categories.
+Particularly, it surpasses other models in identifying unsafe images from the MultiHeaded dataset, which contains 800 AI-generated images by different text-to-image models.
+This improvement can be attributed to the inclusion of a large number of AI-generated images from UnsafeBench in its training data.
+Table 7: Robust accuracy of PerspectiveVision.
+Dataset GN FGSM PGD DeepFool Average
+UnsafeBench-Test 0.986 0.960 0.926 0.936 0.952
+MultiHeaded-D 0.966 0.912 0.886 0.884 0.912
+Robustness Evaluation.
+We also test the robust accuracy of LoRA fine-tuned LLaVA against Gaussian and adversarial perturbations.
+Following the same setup in robustness evaluation, we randomly extract 500 correctly classified images from the UnsafeBench test set and an OOD AI-generated dataset, MultiHeaded-D, and craft adversarial examples with the same perturbation budget.
+We find that LoRA fine-tuning significantly improves robustness in this image safety classification task.
+While zero-shot LLaVA shows an average RA of 0.666-0.725 (see Table 5) under four perturbations, fine-tuned LLaVA increases the average RA to 0.952 on UnsafeBench test set and 0.912 on OOD AI-generated images, in Table 7.
+Take-Aways.
+We build a comprehensive image classifier, PerspectiveVision, which identifies unsafe images across 11 categories by fine-tuning LLaVA on our UnsafeBench dataset.
+The inclusion of diverse AI-generated images not only enhances PerspectiveVision’s effectiveness, particularly with AI-generated images, but also significantly improves its robustness under the same strength of adversarial attacks.
+Related Work
+Visual Content Moderation.
+Moderating visual content is a critical task for both the research community and platform moderators.
+To mitigate the proliferation of unsafe online images, researchers propose various solutions.
+For detecting sexual and pornographic images, NudeNet [15] and NSFW_Detector [14] are developed, which are trained on real-world NSFW images.
+To identify a wider range of unsafe content, Chramowski et al.
+[60] build Q16 using the prompt learning technique on a dataset containing morally negative/positive images.
+Additionally, other researchers focus on detecting specific subsets of unsafe images, such as hateful memes [42, 58, 52, 36] and violent protest images [69].
+On the commercial front, platform moderators are also actively engaged in proposing moderation solutions.
+Google’s SafeSearch detection API [3] is capable of identifying unsafe content across five categories: adult, spoof, medical, violence, and racy.
+Similarly, Microsoft provides an image moderation API [10] that specifically evaluates adult and racy content.
+Despite the variety of these solutions, employing open-source classifiers is a common practice within the research community to identify unsafe images.
+However, their performances on real-world images are under-explored, largely due to the absence of large labeled datasets.
+In our study, we first construct a comprehensive dataset encompassing a broad spectrum of unsafe content.
+We then thoroughly analyze their performances, including the effectiveness across different unsafe categories and the robustness against adversarial examples.
+Counteracting AI-Generated Unsafe Images.
+Since text-to-image models like Stable Diffusion gained popularity in 2022, concerns have been raised regarding their risks of generating realistic unsafe images.
+Plenty of studies [59, 52, 70, 71, 27] focus on assessing these models’ risks.
+Schramowski et al.
+[59] take the first step in estimating the probability of Stable Diffusion in generating unsafe images when providing harmful prompts.
+Qu et al.
+[52] adopt a similar approach and find that text-to-image models are prone to generate sexually explicit, violent, disturbing, hateful, and political images.
+Other researchers investigate the proactive generation of unsafe images from text-to-image models through various attacks, such as data poisoning attacks [70] and adversarial examples [71, 27].
+To mitigate the risks, another line of research focuses on enhancing safety measures.
+For example, Schramowski et al.
+[59] propose safe latent diffusion, which steers the generated images away from a list of unsafe concepts during the generation process.
+Guo et al.
+[38] takes the first step in using VLMs and chain-of-thought to identify unsafe images from user-generated content in games.
+All the above works rely on existing image safety classifiers or VLMs to identify the unsafe images generated by text-to-image models.
+However, since these classifiers are mostly trained on real-world images, it is unclear how effectively they generalize to AI-generated images.
+Our benchmarking framework, UnsafeBench, investigates their ability to generalize to AI-generated unsafe images and explores the distinct characteristics of those images.
+Conclusion & Discussion
+We establish UnsafeBench, a benchmarking framework that comprehensively evaluates the effectiveness and robustness of image safety classifiers on both real-world and AI-generated images.
+We construct a large image safety dataset of 10K manually annotated images and evaluate five conventional classifiers and three VLMs.
+Our analysis reveals the drawbacks of these classifiers and the unique challenges posed by AI-generated images to both their effectiveness and robustness.
+We therefore introduce an improved image moderation tool, PerspectiveVision, that accounts for both real-world and AI-generated images.
+Our work has the following implications for the research community and platform moderators.
+Importance of UnsafeBench Dataset.
+We provide the first large-scale image safety dataset with 11 categories of unsafe content and two sources.
+Annotated images in this dataset are equipped with three “labels”: 1) safe/unsafe, 2) specific category, and 3) sourced from the real world or AI.
+These attributes greatly improve the dataset’s utility, enabling future studies to investigate unsafe images in a multifaceted manner.
+We hope the dataset serves as a solid foundation for future image safety research.
+Challenges Posed by AI-Generated Images.
+In the era of generative AI, the influence of AI-generated content like unsafe imagery and deepfakes can be complex and multifaceted to existing models.
+Our analysis reveals that they do not only reduce the effectiveness of these models but also render them more vulnerable to adversarial attacks than real-world images.
+This is because, although current AI-generated images are realistic, there remains a distribution shift from real-world images in image quality statistics (noise level, entropy, etc.), styles, and layouts.
+For platform moderators and model developers, it is essential to include AI-generated unsafe images in the training data to mitigate these negative impacts.
+Utility of PerspectiveVision.
+PerspectiveVision is built to address the drawbacks of existing image safety classifiers, including imbalanced performance across categories, state-of-the-art model (GPT-4V) being closed-source, and degraded performances on AI-generated images.
+It is also the first attempt to build an image moderation tool that covers both the wide spectrum of unsafe content and unsafe images from different sources.
+The research community and platform moderators can directly apply it for various moderation tasks, such as reporting online unsafe images and evaluating the risks of new text-to-image models.
+Limitations.
+Our work has limitations.
+First, the UnsafeBench dataset is annotated by only three authors, with the final label determined by majority vote.
+We did not rely on crowdsourcing workers for two reasons: 1) annotation requires expert knowledge in the image safety domain, such as identifying hate symbols, which cannot be guaranteed and would require specific training for crowdsourcing workers; 2) due to ethical considerations, we aim to minimize the number of annotators exposed to such unsafe content and prevent them from exposing it to third parties.
+Second, images sourced from Lexica are mostly generated by a single text-to-image model, Stable Diffusion.
+We will expand our dataset with images generated by other text-to-image models and routinely update it.
+Third, we only evaluate the generalizability of PerspectiveVision on a limited range of unsafe categories due to the lack of labeled datasets.
+We will continue expanding the evaluation with more datasets and improving the model in future work.
+Appendix A Appendix
+Details of Conventional Classifiers
 Q16 [60].
@@
 It has been extensively applied in detecting sexually explicit images  [59, 68, 48, 35, 73] and is often combined with Q16 for identifying general unsafe content [59, 35, 48, 73].
-Large Visual Language Models (VLMs)
-Large visual language models have achieved extraordinary capabilities in understanding visual and text content.
-Given an image and a text instruction, these models can read the image and generate responses following the instruction.
-Recent studies [41, 59] show that VLMs can be used to detect user-generated unsafe images [41] and hateful memes [59].
+Details of Large VLMs
 LLaVA [45].
@@
 In this study, we use the gpt-4-vision-preview checkpoint.
-Overview of UnsafeBench
-We build UnsafeBench, a benchmarking framework that comprehensively evaluates the performance of image safety classifiers.
-We show an overview of UnsafeBench in Figure 1.
-Due to the absence of a comprehensive labeled dataset in the image safety domain, we first construct the UnsafeBench dataset, which contains 10K images with human annotations.
-We then collect the existing image safety classifiers, as conventional classifiers, and three VLMs capable of classifying unsafe images, as VLM-based classifiers.
-We identify the range of unsafe content covered by these classifiers and align them with our unsafe image taxonomy, i.e., 11 unsafe categories.
-Finally, we evaluate the effectiveness and robustness of these classifiers; effectiveness measures how accurately the classifiers can identify unsafe images, while robustness reflects their ability to maintain accuracy when faced with perturbed images.
-Figure 1: Overview of UnsafeBench.
-Dataset Construction
-Image Collection.
-We rely on open large-scale multimodal datasets, LAION-5B [65] and Lexica [10], to collect real-world and AI-generated unsafe images.
-LAION-5B is currently the largest public image-text dataset in the world [8], containing 5.85 billion image-text pairs collected from webpages [65].
-We regard LAION-5B as the source to collect real-world images.
-[^10] It is possible that webpages also contain AI-generated images.
-However, LAION-5B was collected in March, 2022 [8], while text-to-image generative models like Stable Diffusion and Midjourney gained popularity after July 2022 [25, 13].
-Thus, we believe the majority of the images in LAION-5B are real-world images.
-Lexica [10] is an AI-generated image gallery that allows users to create and disclose images generated by models like Stable Diffusion [60].
-Both datasets function as a search engine and enable users to find the most relevant images based on a textual description.
-Inspired by this, we use unsafe keywords to query these datasets and collect potentially unsafe images.
-To collect unsafe keywords, we split the definitions of 11 unsafe categories (see Section 2) as initial unsafe keywords.
-Additionally, we employ a large language model, Vicuna-33b, [^11] https://huggingface.co/lmsys/vicuna-33b-v1.3.
-to augment the definition of each unsafe category and provide 10 specific examples.
-We show the generated examples in Table 8 in the Appendix.
-Combining the initial unsafe keywords and those generated with Vicuna-33b, we obtain 158 unsafe keywords, covering 11 unsafe categories (each containing 12-16 keywords).
-We then query LAION-5B and Lexica using these keywords and collect the most relevant images.
-Specifically, for each query, Lexica returns 50 images by default.
-LAION-5B, on the other hand, returns a flexible number of image URLs.
-However, some URLs are repetitive or inaccessible.
-To ensure a balanced number of images from both datasets, we set the number of returned images from LAION-5B to 100.
-After removing duplicates, we collect a total of 12,932 potentially unsafe images, comprising 5,815 images from LAION-5B and 7,117 from Lexica.
-Image Annotation.
-We perform a human annotation to determine if these collected images are truly unsafe.
-Three authors of this paper serve as annotators and manually annotate them independently.
-We require these annotators to first read the definition from each unsafe category as the criterion in determining whether an image is safe, unsafe, or not applicable (N/A).
-Note that N/A mainly represents noise images that cannot be identified as safe or unsafe, e.g., a blurry image or one with unidentifiable texts.
-The annotation process undergoes two rounds.
-In the first round, two annotators independently assign a label of safe, unsafe, or N/A to each image.
-For images where two annotators disagree, we further introduce the third annotator to provide additional labels.
-The final label for each image is then determined based on the majority vote among these labels.
-In the second round, annotators revisit images annotated as unsafe and determine the unsafe category.
-If an image displays a mix of unsafe elements, we choose the predominant category it violates.
-Through two rounds of annotation, we ensure that each image is labeled as safe, unsafe, or N/A, and if it is unsafe, it can be classified into a specific unsafe category.
-To evaluate the reliability of our annotation, we calculate the Fleiss’ Kappa score [36, 35], a statistical measure for assessing the agreement among annotators.
-The score is 0.47, indicating a moderate level of agreement among annotators [36, 35].
-We understand such a Fleiss’ Kappa score is not perfect, as the annotators’ perceptions of unsafe images may still vary under the same definition of our unsafe image taxonomy.
-That explains why it is necessary to introduce the third annotator in the first round to improve the reliability of our annotation.
-Table 2: Statistics of the annotated images.
-Dataset # All # Safe # Unsafe # N/A Fleiss’ Kappa
-LAION-5B 5,815 3,228 1,832 755 0.49
-Lexica 7,117 2,870 2,216 2,031 0.46
-All 12,932 6,098 4,048 2,786 0.47
-Table 3: Aligning the unsafe content covered by image safety classifiers with 11 unsafe categories.
-Classifier Hate Harassment Violence Self-Harm Sexual Shocking Illegal Activity Deception Political Health Spam
-Conventional Classifiers Q16
-MultiHeaded
-SD_Filter
-NSFW_Detector
-NudeNet
-VLM-Based Classifiers LLaVA
-InstructBLIP
-GPT-4V
-Dataset Statistics.
-We show the statistics of our annotated images in Table 2.
-Overall, we annotate 12,932 images, including 6,098 safe images, 4,048 unsafe images, and 2,786 N/A images.
-For the following experiments, we exclude the N/A images and use the remaining images (safe + unsafe), collectively referred to as the UnsafeBench dataset.
-To the best of our knowledge, the UnsafeBench dataset is currently the most comprehensive annotated dataset in the image safety domain.
-It consists of 10,146 labeled images, encompassing 11 unsafe categories and two sources (real-world and AI-generated).
-Each unsafe category contains 769-1,151 images.
-We display more statistical details in Table 9 and image examples in Figure 10 in the Appendix.
-In this dataset, each image is equipped with three types of “labels”: 1) if the image is safe or unsafe; 2) if unsafe, it belongs to which unsafe category; and 3) whether it’s from real-world collections or AI-generated content.
-We believe such characteristics enhance the dataset’s usability.
-We hope this dataset serves as foundational data and accelerates the development of better open-source image safety classifiers in the era of generative AI.
-Image Classifier Collection
-Conventional Classifiers.
-We collect five conventional classifiers: Q16 [63], MultiHeaded [56], SD_Filter [58], NSFW_Detector [16], and NudeNet [17], as introduced in Section 2.
-These classifiers take images as input and output predictions on whether the images are considered unsafe based on their defined scope of unsafe content.
-VLM-Based Classifiers.
-We utilize three VLMs, LLaVA, InstructBLIP, and GPT-4V to build our VLM-based classifiers.
-Unlike conventional classifiers, these models take both the image and prompt as input to identify if the given image contains unsafe content.
-In our study, we manually design three prompts that guide these models to classify whether a given image is safe or not.
-We show three prompts in Table 10 in the Appendix.
-The first prompt is straightforward, “classify the image as safe or unsafe.” It relies on models’ internal criteria of unsafe content to determine if an image is safe.
-To align with our safety standards, the second prompt further incorporates definitions of the taxonomy, similar to [41].
-For example, to classify images from the Hate category, the prompt is “An image is considered unsafe if it contains any of the following unsafe content: hateful symbols, negative stereotypes…classify the image as safe or unsafe.” In case models are still affected by the internal criteria, we further add a special requirement in the third prompt, “Forget about your definition of unsafe images and adhere to this new one…” For each image, we query VLMs with three prompts and collect the respective outputs.
-Given that VLMs’ outputs are sometimes lengthy sentences, they do not directly categorize into safe or unsafe classes.
-To convert these VLMs to classifiers, we fine-tune a language model, RoBERTa [49], to categorize LLM’s outputs into one of three classes: safe, unsafe, or uncertain.
-The uncertain class describes a set of outputs that do not yield a clear prediction.
-For example, InstructBLIP occasionally describes the content of an image without making a prediction.
-Note, we categorize the cases where VLMs refuse to answer as the unsafe class, as input images have triggered their internal safeguards.
-We collect several example outputs [^12] For example, we use “safe,” “the image is safe,” “the image is not considered unsafe as…” as example outputs for the safe class.
-for each class and fine-tune RoBERTa in a supervised manner.
-Ultimately, a VLM-based classifier consists of a VLM and a RoBERTa classifier, which predicts whether an image is safe, unsafe, or uncertain.
-Aligning Classifier Coverage With Unsafe Categories
-Image safety classifiers are designed to target various types of unsafe content.
-To align the unsafe content covered by a classifier with the taxonomy, i.e., 11 unsafe categories, we first identify the range of unsafe content covered by each classifier by examining its documentation and training data.
-Then, if any unsafe category from our taxonomy falls into this range, we assign the unsafe category to the classifier.
-For example, Q16 [63] detects “morally negative” images.
-We examine the definition of “morally negative” images in [32] and find that it covers a list of unsafe concepts, such as harm, inequality, degradation, etc.
-By mapping these unsafe concepts to our taxonomy, we align the ambiguous definition of “morally negative” with specific unsafe categories.
-For VLMs, as we adopt the unsafe image taxonomy from OpenAI, we assume that GPT-4V covers all 11 categories.
-We also infer that LLaVA and InstructBLIP are capable of handling these categories since they are fine-tuned on the instruction dataset generated by GPT-4V [48, 33].
-We show the aligning result in Table 3.
-Table 4: F1-Score of eight image safety classifiers on the UnsafeBench dataset. * marks VLM-based classifiers.
-Dataset Classifier Hate Harassment Violence Self-Harm Sexual Shocking Illegal Activity Deception Political Health Spam Overall
-Real-World Images (LAION-5B) Q16 0.641 0.517 0.693 0.421 - 0.630 0.681 0.762 0.271 0.144 - -
-MultiHeaded 0.320 - 0.247 0.000 0.692 0.644 - - 0.209 - - -
-SD_Filter - - - - 0.833 - - - - - - -
-NSFW_Detector - 0.517 - - 0.783 - - - - - - -
-NudeNet - - - - 0.650 - - - - - - -
-LLaVA* 0.227 0.167 0.413 0.451 0.674 0.714 0.498 0.376 0.116 0.333 0.059 0.426
-InstructBLIP* 0.351 0.394 0.606 0.275 0.796 0.467 0.722 0.653 0.444 0.379 0.380 0.520
-GPT-4V* 0.556 0.706 0.774 0.557 0.866 0.724 0.897 0.827 0.605 0.405 0.718 0.712
-AI-Generated Images (Lexica) Q16 0.336 0.416 0.612 0.521 - 0.836 0.497 0.384 0.597 0.615 - -
-MultiHeaded 0.225 - 0.533 - 0.815 0.780 - - 0.744 - - -
-SD_Filter - - - - 0.727 - - - - - - -
-NSFW_Detector - 0.259 - - 0.678 - - - - - - -
-NudeNet - - - - 0.596 - - - - - - -
-LLaVA* 0.169 0.224 0.632 0.663 0.866 0.826 0.534 0.179 0.340 0.665 0.045 0.634
-InstructBLIP* 0.178 0.332 0.629 0.383 0.757 0.795 0.665 0.302 0.785 0.663 0.519 0.589
-GPT-4V* 0.254 0.635 0.712 0.613 0.827 0.875 0.701 0.422 0.909 0.621 0.171 0.707
-Overall Q16 0.533 0.475 0.648 0.483 - 0.784 0.571 0.652 0.482 0.498 - -
-MultiHeaded 0.292 - 0.426 - 0.757 0.749 - - 0.600 - - -
-SD_Filter - - - - 0.785 - - - - - - -
-NSFW_Detector - 0.449 - - 0.738 - - - - - - -
-NudeNet - - - - 0.624 - - - - - - -
-LLaVA* 0.210 0.189 0.550 0.590 0.780 0.800 0.519 0.332 0.266 0.575 0.054 0.549
-InstructBLIP* 0.270 0.368 0.615 0.333 0.777 0.697 0.687 0.506 0.660 0.545 0.490 0.559
-GPT-4V* 0.423 0.681 0.738 0.590 0.847 0.839 0.780 0.673 0.780 0.492 0.537 0.709
-Evaluation Methodology
-Effectiveness.
-We evaluate the effectiveness of eight image safety classifiers (five conventional classifiers and three VLM-based classifiers).
-For conventional classifiers, we feed images from each unsafe category and obtain binary predictions (safe/unsafe).
-Regarding VLM-based classifiers, we input both images with the designed prompts to gather outputs, which are then classified into safe/unsafe/uncertain classes by the fine-tuned RoBERTa.
-Uncertain predictions account for about 0.1%.
-However, since conventional classifiers directly predict images as safe or unsafe, to ensure a fair comparison, we randomly assign a prediction of safe or unsafe to VLM’s uncertain predictions.
-Recall that we design three prompts for VLMs, resulting in three separate predictions for each image.
-The final prediction is determined by a majority vote.
-Evaluation Metric for Effectiveness.
-With the predictions in place, we calculate the F1-Score to evaluate the effectiveness of image safety classifiers.
-We choose F1-Score as it accounts for both false positives and false negatives, providing a balanced measure of the classifier’s Precision and Recall.
-Robustness.
-Robustness reflects the classifier’s ability to maintain its performance beyond training examples.
-We evaluate the robustness of classifiers using images perturbed with both random noise (Gaussian noise) and adversarial noise (i.e., adversarial examples) [52, 50].
-Adversarial examples ($x_{adv}$) are original inputs ($x$) with optimized perturbations ($\Delta x$), which maximizes the loss ($L(\theta,x+\Delta x,y)$) and fools the model into making incorrect predictions.
-The p-norm of the perturbation vector is limited by $\epsilon$ to ensure the perturbation is imperceptible.
-$\begin{gathered}x_{adv}:=\underset{\Delta x}{\text{argmax}}\,L(\theta,x+\Delta x,y),\\ \quad\left\|\Delta x\right\|_{p}<\epsilon.\end{gathered}$ (1)
-We craft adversarial examples using three gradient-based adversarial algorithms, FGSM [40], PGD [50], and DeepFool [53], with details introduced in Section A.1 in the Appendix.
-These algorithms optimize perturbations using gradients of the classifiers’ loss with respect to the input image.
-For conventional classifiers, we directly optimize the perturbations using three types of adversarial algorithms.
-However, for VLM-based classifiers (VLM+RoBERTa), the standard optimization approach is not applicable because we lose continuous gradient flow after VLMs generate word tokens.
-To address this challenge, we approximate the gradients by instructing the VLMs to generate a target output ($y_{tar}$), which is contrary to the image label.
-For example, if the VLM initially classifies a safe image with the output “This image is safe,” we compute the gradients when the target output is “unsafe.” Different from conventional classifiers, where we maximize the loss between the classifier’s prediction and the original label ($y$), here, we minimize the loss of a VLM between its prediction and the defined target output ($y_{tar}$).
-We craft adversarial examples with FGSM and PGD for open-source VLM-based classifiers (LLaVA and InstructBLIP).
-Note, we exclude DeepFool as it requires the gradients of logits, which are not directly applicable in this context.
-$\begin{gathered}x_{adv}:=\underset{\Delta x}{\text{argmin}}\,L(\theta,x+\Delta x,y_{tar}),\\ \quad\left\|\Delta x\right\|_{p}<\epsilon.\end{gathered}$ (2)
-Evaluation Metric for Robustness.
-We calculate the Robust Accuracy (RA) to evaluate the robustness of image safety classifiers.
-RA is the percentage of perturbed images that have been correctly predicted by classifiers out of all perturbed images.
-Here, we generate perturbed images only for those that are correctly predicted by classifiers during the effectiveness evaluation.
-Furthermore, we calculate the Mean Square Error (MSE) between original images and adversarial examples to reveal the magnitude of perturbations required to mislead classifiers into making incorrect predictions.
-Evaluation Results
-Effectiveness
-Table 4 shows the effectiveness of eight image classifiers on the UnsafeBench dataset.
-This table is consistent with the aligning result shown in Table 3.
-If a classifier cannot identify a specific unsafe category according to Table 3, we fill the corresponding cell with a “-.”
-Classifier-Level.
-The top-performing image safety classifier is the commercial VLM-based classifier, GPT-4V.
-It achieves the highest overall F1-Score, surpassing LLaVA and InstructBLIP, as well as other conventional classifiers in most unsafe categories, except for Hate and Deception.
-GPT-4V shows exceptional effectiveness in detecting Sexual (0.847), Shocking (0.839), Political (0.780), Illegal Activity (0.780), and Violence content (0.738).
-Among conventional classifiers, Q16 stands out by identifying the broadest spectrum of unsafe content, with an F1-Score from 0.475 to 0.784.
-However, Q16 does not support the detection of Sexual content and requires improvements in Harassment, Self-Harm, and Political categories.
-In contrast, MultiHeaded, SD_Filter, NSFW_Detector, and NudeNet can detect sexual images with an F1-Score of 0.624-0.785.
-Although GPT-4V shows superior performance on our annotated dataset, the wide application is constrained by its commercial nature, mainly due to the financial cost and slow inference speed.
-It costs approximately \$20 and 50 minutes to classify 1K images.
-Additionally, we observe that zero-shot VLMs’ performances significantly rely on prompt designing.
-For example, GPT-4V’s overall F1-Score achieves 0.68 and 0.71 when we classify images with the second and third prompts (shown in Table 10 in the Appendix).
-However, the score drops to 0.61 on the first prompt.
-To obtain reliable results from VLMs like GPT-4V, we typically query the VLM with the same image using multiple prompts and determine the label based on a majority vote.
-This approach further increases the cost, making GPT-4V impractical for application on large-scale image datasets.
-Current research [62, 38, 51, 76] prefers smaller classifiers over larger VLMs due to their faster inference speed.
-A common practice is combining Q16 and a classifier that detects sexually explicit content like NudeNet to detect generally unsafe images [62, 38, 51, 76], e.g., regarding the image is unsafe if either Q16 or NudeNet classifies it as unsafe.
-According to Table 4, this strategy can indeed cover many unsafe categories with a relatively good performance.
-We calculate the overall F1-Score of Q16 combined with NudeNet across the unsafe categories they can cover, and the score is 0.665.
-This performance needs to be further enhanced.
-Different Unsafe Categories.
-To understand which unsafe categories are more effectively detected, we count the number of classifiers capable of identifying each category and record the average F1-Score in Figure 2.
-We find that the Sexual and Shocking categories have higher average F1-Scores, close to 0.8, compared to other unsafe categories.
-They are also covered by more classifiers, with 5-7 classifiers capable of identifying such content.
-In contrast, the remaining unsafe categories, particularly Spam, Hate, Harassment, and Self-Harm, have lower average F1-Scores (below 0.5).
-This discrepancy highlights the imbalanced moderation across different types of unsafe content.
-Figure 2: Average F1-Score and number of classifiers for each unsafe category.
-Among 11 unsafe categories, we particularly focus on the Hate category, given that hateful images can proliferate across Web communities as memes and are often used in coordinated hate campaigns [75, 39].
-For instance, the anti-Semitic symbol, Happy Merchant [6], widely spreads on platforms like 4chan and mainly targets the Jewish community.
-Due to the harmful nature, OpenAI ranks the Hate category as the highest priority in its content policy [19].
-However, among the five classifiers that can detect hateful content, the highest F1-Score is 0.533, achieved by Q16.
-The second highest is 0.423, achieved by GPT-4V.
-To understand the underlying reason behind such low F1-Scores, we zoom into the misclassified examples from the Hate category for both Q16 and GPT-4V.
-We calculate the false positive and false negative rates for both classifiers.
-For Q16, the false positive rate is 0.166 and the false negative rate is 0.306.
-In comparison, for GPT-4V, the false positive rate is 0.425, and the false negative rate is 0.144.
-This result suggests that both classifiers have a considerable likelihood to misclassify hateful images: either failing to identify truly hateful images or incorrectly predicting safe images as hateful.
-For model providers and platform moderators, the false negative rate may be more important as it indicates the likelihood of hateful images evading the detection of models’ safeguards.
-In our manual review of common false negatives, we observe that 58 images containing hateful content are misclassified as safe by both Q16 and GPT-4V.
-Notably, some Neo-Nazi and anti-Semitic symbols, such as the tattooed swastika [15] and the Happy Merchant meme [6], manage to evade detection.
-We show these examples in Figure 12 in the Appendix.
-Figure 3: Average F1-Score of classifiers on real-world and AI-generated images.
-(a) False Negatives (Misclassify Unsafe as Safe)
-(b) False Positives (Misclassify Safe as Unsafe)
-Figure 4: Image clusters from the Sexual category that are misclassified by SD_Filter, NSFW_Detector, and NudeNet.
-We annotate each central image with its cluster ID and cluster size.
-We blur sexual images for censoring purposes.
-Real-World Versus AI-Generated.
-Real-world images from LAION-5B and AI-generated ones from Lexica have different distributions (see 11(a) in the Appendix).
-To compare the performance of tested classifiers on two groups of images, we calculate the average F1-Score of each classifier on real-world and AI-generated images in Figure 3.
-We find that SD_Filter, NSFW_Detector, NudeNet, and GPT-4V have better performance in identifying real-world unsafe images compared to AI-generated ones.
-For example, the F1-Score of SD_Filter decreases from 0.833 on real-world images to 0.727 on AI-generated images.
-Meanwhile, MultiHeaded, LLaVA, and InstructBLIP have higher F1-Scores on AI-generated images than real-world ones.
-The different behaviors of classifiers on two sets of images are multifaceted.
-A key factor is the distribution shift between training and testing images, including nuanced differences in semantics, image styles, etc.
-After examining the training images of the classifiers (if available), we find that classifiers like NudeNet and NSFW_Detector are trained on real-world NSFW images [18], as indicated by [17] and [16].
-Meanwhile, the training data for SD_Filter and GPT-4V has not been disclosed.
-For classifiers that are more effective on AI-generated images, i.e., MultiHeaded, we find that it is trained on images generated by text-to-image models.
-Next, we explore in detail why certain classifiers experience degraded performance on AI-generated images, and what characteristics of AI-generated images contribute to it.
-We select two groups of classifiers that exhibit degraded performances on AI-generated images compared to real-world images.
-The first group consists of SD_Filter, NFSW_Detector, and NudeNet.
-These classifiers achieve F1-Scores ranging from 0.650 to 0.833 in categorizing real-world images within the Sexual category.
-However, the scores drop to 0.596-0.727 when applied to AI-generated images based on Table 4.
-The second group, including Q16 and GPT-4V, recognizes real-world violent images with F1-Scores of 0.693 to 0.774 but drops to 0.612 to 0.712 for AI-generated violent images.
-To investigate the reason behind this performance degradation on AI-generated images, we again examine the misclassified examples, including false negatives and false positives from both LAION-5B and Lexica.
-To characterize images from these two sources, we perform KMeans clustering and group these misclassified examples into $K$ clusters, with images from the same cluster sharing similar semantics.
-Specifically, we utilized the CLIP image encoder to generate image embeddings that capture semantic information.
-We then apply KMeans clustering with $K$ ranging from 2 to 15 and employ the Elbow method [^13] https://www.scikit-yb.org/en/latest/api/cluster/elbow.html.
-to determine the optimal number of clusters.
-This method shows that the optimal number of clusters is 5-8 depending on different classifiers.
-Finally, we create five clusters and retrieve the nearest image to each cluster centroid.
-We display these central images, representing their respective clusters in Figure 4 and Figure 13 in the Appendix, each figure corresponding to one group of classifiers.
-(a) Strengths of Artistic Representation
-(b) Grid Sizes
-Figure 5: The original real-world image and its AI-generated variations applying the artistic style and grid layout.
-The original image is unsafe and correctly predicted by Q16.
-Text in red indicates that image variations are misclassified as safe.
-Figure 6: Percentage of correct predictions by NudeNet, NSFW_Detector, and Q16 for images with various strengths of artistic representation and grid sizes.
-Comparing the misclassified images between real-world and AI-generated ones, we observe two distinct characteristics prevalent in AI-generated images: artistic representation and grid layout.
-First, images in an artistic representation, [^14] It is an open question whether unsafe content in artistic representation remains unsafe.
-In this study, we adhere to the criteria from the unsafe image taxonomy and evaluate image safety based on its content and intention, regardless of the representation form.
-despite showing explicitly unsafe content, can often escape detection of classifiers.
-For example, among false negatives shown in 4(a), AI-generated images within clusters 1, 2, 4, and 5 showcase nudity in an artistic style, and yet they are misclassified as safe.
-In contrast, real-world images depicting realistic nudity are rarely misclassified as safe, given that they also widely exist in our annotated dataset.
-Similarly, among the misclassified images from the Violence category (see 13(a) in the Appendix), AI-generated images within clusters 1, 2, and 4 are also misclassified, which portray artistic representations of violence, such as fighting scenes.
-Second, images in a grid layout from AI-generated ones are frequently misclassified by classifiers.
-These images are composites made of multiple smaller images, commonly four or nine, arranged in a grid layout.
-Examples include cluster 2, 3, 5 in 4(a), cluster 5 in 4(b), and cluster 2, 4 in 13(a).
-Based on these findings, we hypothesize that these distinctive characteristics of AI-generated images, like artistic representation and grid layout, contribute to the performance degradation, particularly for models trained on real-world images, e.g., NSFW_Detector, NudeNet, and Q16.
-To test this hypothesis, we conduct a case study where we generate image variations with different levels of artistic representation and grid sizes to observe if classifiers can consistently make accurate predictions.
-To this end, we randomly select 20 realistic unsafe images from LAION-5B in both Sexual and Violence categories.
-These images are accurately identified as either sexual content by NSFW_Detector and NudeNet or as violent content by Q16, and these three classifiers have been verified to be trained on real-world images only.
-For each unsafe image, we then create artistic versions using the shape-preserving image editing function of a text-to-image model, Stable Diffusion [60].
-Specifically, we input the image into the model and use the prompt “in the style of oil painting” to generate artistic images.
-By adjusting the strength parameter, we control the degree of the artistic style applied to the original image.
-For each image, we produce six artistic versions with the strength parameter ranging from 0.1 to 0.6.
-Additionally, we generate image variations with different grid sizes by reconstructing an image with $2\times 2$, $3\times 3$, and $4\times 4$ grids, with each smaller image serving as a part of the grid.
-Figure 5 shows image variations with an example from the Violence category.
-In total, we collect 40 original images and their 360 variations, including 240 ($40\times 6$) artistic and 120 ($40\times 3$) grid variations.
-To guarantee that the variations in the artistic representation retain the same content as the original images, we conducted a manual annotation to filter out any variations where the content has been altered.
-We use the remaining 17 original images and their 153 variations to test NSFW_Detector, NudeNet, and Q16.
-We present the percentage of correct predictions in Figure 6.
-The result indicates that with the increasing strength of artistic representation and grid sizes, the classifiers’ performances generally decrease.
-This confirms the impact of these characteristics found in AI-generated images on the performance of classifiers trained on real-world images only.
-Table 5: Robustness of image safety classifier against Gaussian perturbation and three types of adversarial perturbations. We report the MSE (average square error) and RA (robust accuracy).
-GN FGSM PGD DeepFool Average
-Model MSE ($\uparrow$) RA ($\uparrow$) MSE ($\uparrow$) RA ($\uparrow$) MSE ($\uparrow$) RA ($\uparrow$) MSE ($\uparrow$) RA ($\uparrow$) MSE ($\uparrow$) RA ($\uparrow$)
-Q16 <1e-3 0.972 <1e-3 0.402 <1e-3 0.004 <1e-3 0.160 <1e-3 0.385
-MultiHeaded <1e-3 0.988 <1e-3 0.596 <1e-3 0.012 <1e-3 0.218 <1e-3 0.454
-SD_Filter <1e-3 1.000 <1e-3 0.552 <1e-3 0.000 <1e-3 0.876 <1e-3 0.607
-NSFW_Detector <1e-3 0.976 <1e-3 0.496 <1e-3 0.012 <1e-3 0.372 <1e-3 0.464
-NudeNet <1e-3 1.000 <1e-3 0.086 <1e-3 0.004 <1e-3 0.084 <1e-3 0.293
-LLaVA* <1e-3 0.996 1.365 0.710 0.849 0.180 - - 0.738 0.629
-InstructBLIP* <1e-3 0.996 1.417 0.624 0.881 0.204 - - 0.766 0.608
-Take-Aways.
-We assess the effectiveness of eight image safety classifiers on real-world and AI-generated images across 11 unsafe categories.
-Our findings reveal several insights.
-First, of all the classifiers evaluated, the commercial model GPT-4V stands out as the most effective in identifying a broad spectrum of unsafe content.
-However, due to its associated time and financial costs, it is impractical to moderate large-scale image datasets.
-Currently, there is a lack of an open-source image safety classifier that can comprehensively and effectively identify unsafe content.
-Second, the effectiveness varies significantly across 11 unsafe categories.
-Images from the Sexual and Shocking categories are detected more effectively, while categories such as Hate require further improvement.
-For instance, we observe that Neo-Nazi and anti-Semitic symbols can sometimes evade detection by both Q16 and GPT-4V.
-Finally, we find certain classifiers trained on real-world images experience performance degradation on AI-generated images.
-The underlying reason is the distribution shift between these two types of images.
-Notably, AI-generated images often exhibit unique characteristics, such as artistic representation of unsafe content and the composition of unsafe images in a grid layout.
-These distinct characteristics may pose challenges to classifiers trained only on real-world images, leading to the observed performance degradation.
-Robustness
-Setup.
-Generating adversarial examples for the entire UnsafeBench dataset can be time-consuming, because it requires performing an independent optimization process for each image.
-To address this issue, we randomly sample 500 images that each classifier correctly predicted to evaluate their robustness.
-We then generate perturbed images by adding Gaussian noise and creating optimized perturbations using three types of adversarial algorithms: FGSM, PGD, and DeepFool.
-We employ the $L_{\infty}$ norm to constrain perturbations and set $\epsilon$ to 0.01 for all experiments.
-We demonstrate the perturbed images and different types of perturbations in Figure 7.
-With these perturbed images, we measure the MSE and RA (metrics are introduced in Section 3.4) of seven open-source classifiers.
-Note that we exclude GPT-4V from this evaluation because crafting adversarial examples requires access to the model’s internal parameters, which is not available for GPT-4V.
-Figure 7: Perturbed images with Gaussian perturbations and three types of adversarial perturbations.
-Adversarial examples generated with FGSM, PGD, and DeepFool cause Q16 to misclassify these unsafe images as safe.
-Results.
-Table 5 demonstrates the MSE and RA of seven open-source classifiers across four types of perturbations.
-VLM-based classifiers show the highest robustness.
-They achieve RAs between 0.629 and 0.608, higher than any conventional classifiers.
-This suggests that a significant percentage of perturbed images can still be correctly classified by VLM-based classifiers.
-Moreover, their MSEs range from 0.738 to 0.766, far exceeding the conventional classifiers, whose MSEs are less than $31e-$ 3.
-This indicates that it requires significant perturbations to deceive VLM-based classifiers.
-Meanwhile, Q16, MultiHeaded, SD_Filter, and NSFW_Detector present a moderate level of robustness, with RAs ranging from 0.385 to 0.607.
-However, NudeNet shows the lowest RA, i.e., 0.293, revealing that it is the most vulnerable classifier to adversarial perturbations.
-Even imperceptible perturbations with MSE less than $31e-$ 3 can deceive NudeNet in most cases involving adversarial examples.
-To explore different levels of robustness, we review the model architecture, training paradigm, and training dataset for each classifier (see Section 2).
-As the least robust classifier, NudeNet is an Xception-based classifier and is trained on 160K labeled images using fully supervised learning.
-Other conventional classifiers are built on CLIP, which is pre-trained on 400 million image-text pairs and then fine-tuned on their own labeled datasets using linear probing or prompt learning.
-They present higher robustness compared to training a small classifier using supervised learning, i.e., NudeNet.
-VLM-based classifiers present the highest robustness and also rely on large pre-trained models.
-For example, LLaVA utilizes CLIP as the image feature extractor and LLaMA [68] as the LLM to reason over the input image and generate output.
-To conclude, we find that under the attack of random perturbed and adversarial examples, relying on large pre-trained foundational models enhances robustness compared to training new, smaller classifiers from scratch.
-For developing future classifiers, adapting large pre-trained foundation models can yield a more resilient system.
-PerspectiveVision
-Motivation & Overview
-The findings from UnsafeBench point out two shortcomings in open-source image safety classifiers, including zero-shot VLMs.
-First, their effectiveness needs to be improved, particularly in several neglected categories such as Hate and Self-harm.
-Second, these classifiers face challenges in detecting AI-generated unsafe images due to distributions that differ from real-world images.
-With AI-generated content becoming increasingly common in our daily lives, there is an urgent need to develop a comprehensive image moderation tool that accounts for these AI products.
-The UnsafeBench dataset serves as a good starting point, as it covers 11 unsafe categories and AI-generated content.
-Using this dataset, we take the first step towards building a comprehensive moderation tool, PerspectiveVision, that can effectively identify a wide range of unsafe images from the real world and AI.
-PerspectiveVision is inspired by Google’s Perspective API, a benchmark tool for detecting toxic text in the NLP domain.
-The API assesses text toxicity across various attributes, such as severe toxicity, hate, and profanity.
-Similarly, PerspectiveVision evaluates the safety of a given image and further identifies the specific category if the image is deemed unsafe.
-Users can customize the scope of unsafe images by selecting multiple categories to suit different definitions of unsafe content.
-The high-level overview is demonstrated in Figure 8.
-Figure 8: High-level overview of PerspectiveVision.
-Table 6: F1-Score of PerspectiveVision models on the UnsafeBench test set.
-Hate Harassment Violence Self-Harm Sexual Shocking Illegal Activity Deception Political Health Spam Overall
-Linear Probing (CLIP) 0.776 0.736 0.767 0.586 0.924 0.894 0.877 0.841 0.907 0.903 0.899 0.859
-Prompt Learning (CLIP) 0.437 0.557 0.679 0.337 0.890 0.758 0.765 0.596 0.829 0.635 0.654 0.687
-LoRA Fine-Tuning (LLaVA) 0.686 0.755 0.760 0.743 0.913 0.861 0.889 0.780 0.878 0.909 0.855 0.844
-Table 7: F1-Score of PerspectiveVision models on external evaluation datasets. We use GPT-4V as a baseline for comparison.
-SMID MultiHeaded-D NudeNet-D Hate Symbols Self-Harm Violent Behaviors Overall
-Linear Probing (CLIP) 0.365 0.654 0.954 0.598 0.801 0.740 0.723
-Prompt Learning (CLIP) 0.623 0.500 0.984 0.849 0.971 0.931 0.810
-LoRA Fine-Tuning (LLaVA) 0.557 0.675 0.986 0.593 0.948 0.908 0.794
-GPT-4V 0.532 0.577 0.991 0.786 0.924 0.921 0.818
-Methodology
-The general idea is to leverage the knowledge of pre-trained models, such as CLIP or VLMs, and adapt them to identify unsafe images under the guidance of annotated images.
-We elaborate on specific methodologies employed with each pre-trained model as follows.
-Adapting CLIP for image classification tasks is a popular choice.
-Among the five conventional classifiers tested in UnsafeBench, four of them (Q16, MultiHeaded, NSFW_Detector, SD_Filter) utilize CLIP for image feature extraction.
-Similar to these classifiers, we employ linear probing and prompt learning to adapt CLIP to identify unsafe images.
-Linear probing [57] is a technique that adds a linear probing layer on top of a pre-trained model, allowing for fine-tuning on specific tasks, and prompt learning [34, 42] refers to a process that adapts a pre-trained model for various downstream tasks by optimizing prompts.
-For linear probing, we train 11 linear probing layers, each one identifying images from a specific unsafe category.
-Similarly, we optimize 11 sets of soft prompts using prompt learning.
-In the inference phase, according to the user’s defined scope of unsafe content, only the linear probing layers or soft prompts associated with the selected unsafe categories will be activated to determine if the image is safe or not.
-Different from classifiers using linear probing and prompt learning, VLMs are generative models that output more information rather than simply logit probabilities.
-If we fine-tune the VLM to generate a response that indicates both the safety status and the specific unsafe category, we can build a unified model with VLMs like LLaVA instead of training classifiers separately for different categories.
-To this end, we employ supervised fine-tuning to adapt LLaVA for identifying unsafe images.
-We first build the fine-tuning dataset based on our annotated images.
-The dataset contains triplet elements: image, prompt, and target output.
-We adopt a similar prompt template as LLaMA-Guard [44] (an LLM-based safeguard), where we define the unsafe image taxonomy with specific unsafe categories and output format.
-Figure 9 demonstrates an example of the LLaVA fine-tuning dataset.
-Since fine-tuning the entire LLaVA is computationally intensive, we resort to the Low-Rank Adaptation (LoRA) technique [43].
-LoRA is a parameter-efficient fine-tuning technique that introduces a small set of additional trainable parameters to the model.
-It updates only these extra parameters while keeping the original model parameters frozen.
-With our fine-tuning dataset and the LoRA technique, we train a unified model that simultaneously predicts the safety status and identifies the specific unsafe category.
-The probability value is extracted from the logits associated with the next token prediction.
-Figure 9: An example of LLaVA fine-tuning dataset.
-“<image>” is the visual token, representing the provided image.
-“Human” presents a prompt that describes the image safety classification task including the unsafe image taxonomy and the required output format.
-“Assistant” provides the target output.
-PerspectiveVision Evaluation
-Datasets.
-We randomly split the UnsafeBench dataset into an 80% training and 20% testing ratio.
-To assess the model’s generalizability, we gather multiple external datasets for further evaluation: SMID [24], MultiHeaded Dataset [14], NudeNet Dataset [18], Hate Symbols [2], Self-Harm Dataset [22, 23], and Violent Behavior Dataset [28].
-We outline these datasets and detailed experimental setup in Section A.2 and Section A.3 in the Appendix.
-Evaluation on UnsafeBench Test Set.
-Table 6 shows the evaluation performance of PerspectiveVision models on the UnsafeBench test set.
-Among the three methods employed, linear probing (CLIP) and LoRA fine-tuning LLaVA demonstrate a higher fitting level on the UnsafeBench test set, with an overall F1-Score of 0.859 and 0.844.
-Meanwhile, prompt learning (CLIP) yields a lower F1-Score of 0.687.
-We observe that these models perform particularly well in identifying images within the Sexual, Political, Spam, and Shocking categories, where the F1-Scores are high, close to or exceeding 0.90.
-Evaluation on External Datasets.
-Table 7 presents the evaluation results of the PerspectiveVision models across six external datasets.
-Here, we benchmark GPT-4V against these datasets to establish a high-performing baseline.
-GPT-4V demonstrates an overall F1-Score of 0.818, particularly standing out when classifying NudeNet images and Hate symbols.
-Prompt learning (CLIP) delivers comparable performance, with an overall F1-Score of 0.810.
-It shows a strong generalization capability in identifying morally negative, hate symbols, self-harm, and violent images.
-Following closely is the LoRA fine-tuned LLaVA, which achieves an F1-Score of 0.794.
-It surpasses other models in identifying AI-generated unsafe images from the MultiHeaded dataset.
-To conclude, prompt learning (CLIP) and LoRA fine-tuned LLaVA demonstrate better generalizability on external evaluation datasets, serving as the primary backbone models for PerspectiveVision.
-Take-Aways.
-We build a comprehensive image classifier, PerspectiveVision, which identifies unsafe images across 11 unsafe categories.
-Similar to Google’s Perspective API, users can customize their definition of unsafe images by selecting from various categories.
-Based on our evaluation, PerspectiveVision performs comparably to zero-shot GPT-4V in identifying unsafe images, with its best backbone achieving an F1-Score of 0.810 on external evaluation datasets.
-We hope it serves as a baseline and will make it publicly available.
-Related Work
-Visual Content Moderation.
-Moderating visual content is a critical task for both the research community and platform moderators.
-To mitigate the proliferation of unsafe online images, researchers propose various solutions.
-For detecting sexual and pornographic images, NudeNet [17] and NSFW_Detector [16] are developed, which are trained on real-world NSFW images.
-To identify a wider range of unsafe content, Chramowski et al.
-[63] build Q16 using the prompt learning technique on a dataset containing morally negative/positive images.
-Additionally, other researchers focus on detecting specific subsets of unsafe images, such as hateful memes [45, 61, 55, 39] and violent protest images [72].
-On the commercial front, platform moderators are also actively engaged in proposing moderation solutions.
-Google’s SafeSearch detection API [4] is capable of identifying unsafe content across five categories: adult, spoof, medical, violence, and racy.
-Similarly, Microsoft provides an image moderation API [11] that specifically evaluates adult and racy content.
-Despite the variety of these solutions, employing open-source classifiers is a common practice within the research community to identify unsafe images.
-However, their performances on real-world images are underexplored, largely due to the absence of large labeled datasets.
-In our study, we first construct a comprehensive dataset encompassing a broad spectrum of unsafe content.
-We then thoroughly analyze their performances, including the effectiveness across different unsafe categories and the robustness against adversarial examples.
-Counteracting AI-Generated Unsafe Images.
-Since text-to-image models like Stable Diffusion gained popularity in 2022, concerns have been raised regarding their risks of generating realistic unsafe images.
-Plenty of studies [62, 55, 73, 74, 30] focus on assessing these models’ risks.
-Schramowski et al.
-[62] take the first step in estimating the probability of Stable Diffusion in generating unsafe images when providing harmful prompts.
-Qu et al.
-[55] adopt a similar approach and find that text-to-image models are prone to generate sexually explicit, violent, disturbing, hateful, and political images.
-Other researchers investigate the proactive generation of unsafe images from text-to-image models through various attacks, such as data poisoning attacks [73] and adversarial examples [74, 30].
-To mitigate the risks, another line of research focuses on enhancing safety measures.
-For example, Schramowski et al.
-[62] propose safe latent diffusion, which steers the generated images away from a list of unsafe concepts during the generation process.
-Guo et al.
-[41] takes the first step in using VLMs to identify unsafe images from user-generated content games.
-All the above works rely on existing image safety classifiers or VLMs to identify the unsafe images generated by text-to-image models.
-However, since these classifiers are mostly trained on real-world images, it is unclear how effectively they generalize to AI-generated images.
-Our benchmarking framework, UnsafeBench, investigates their ability to generalize to AI-generated unsafe images and explores the distinct characteristics of those images.
-Conclusion & Discussion
-We establish UnsafeBench, a benchmarking framework that comprehensively evaluates the performance of image safety classifiers on both real-world and AI-generated images.
-We construct a large image safety dataset of 10K manually annotated images and evaluate five conventional classifiers and three VLMs.
-Our analysis reveals that current image safety classifiers are constrained by imbalanced performance across different types of unsafe content and potential performance degradation on AI-generated images.
-We therefore introduce a comprehensive image moderation tool, PerspectiveVision, to evaluate the safety of a given image.
-Our work has the following implications for the research community, platform moderators, and model providers.
-Importance of UnsafeBench.
-One important contribution of UnsafeBench is the comprehensive dataset of 10K images, which are manually annotated as safe or unsafe.
-Besides these annotations, they can also be categorized into 11 unsafe categories or two sources (real-world or AI-generated).
-These attributes largely improve the dataset’s utility, enabling studying unsafe images in a multifaceted manner.
-We hope the dataset serves as a solid foundation for future image safety research.
-Another important contribution is the performance comparison among different image safety classifiers, which assists future studies in selecting the most suitable classifier according to the target unsafe content.
-Enhancing Effectiveness on AI-Generated Images.
-Our analysis reveals that there exists a performance degradation when identifying unsafe AI-generated images, especially for classifiers trained merely on real-world images.
-This degradation is largely due to the distribution shift between two groups of images.
-Specifically, AI-generated images present distinct characteristics, such as artistic representations of unsafe content and grid layout.
-For platform moderators and model developers, it is necessary to account for AI-generated unsafe images in the era of generative AI.
-Extending the training dataset of image safety classifiers with AI-generated examples enhances generalizability.
-At the same time, we hope the UnsafeBench dataset aids future research to uncover more characteristics of AI-generated unsafe images.
-Utility of PerspectiveVision.
-The results from UnsafeBench suggest that, currently, there is a lack of an open-source image safety classifier capable of effectively detecting a wide range of unsafe content.
-Inspired by Google’s Perspective API, which evaluates the toxicity of the given text from multiple perspectives, we establish a similar tool in the image safety domain, PerspectiveVision, as the first attempt.
-Similar to Perspective API, our tool enables users to customize the scope of unsafe content.
-The research community and platform moderators can directly apply it for various moderation tasks, such as reporting online unsafe images and evaluating the risks of new text-to-image models.
-Limitations.
-Our work has limitations.
-First, the UnsafeBench dataset is biased towards unsafe content, as it is curated by retrieving images most related to unsafe keywords.
-Second, images sourced from Lexica are mostly generated by a single text-to-image model, Stable Diffusion.
-We will expand our dataset with images generated by other text-to-image models and routinely update it.
-Third, we only evaluate the generalizability of PerspectiveVision on a limited range of unsafe categories due to the lack of labeled datasets.
-We will continue expanding the evaluation with more datasets and improving the model in future work.
-Appendix A Appendix
 Adversarial Examples
@@
 To obtain the perturbation at step $t$, it computes the gradient of the model’s decision function with respect to the input, $\nabla f_{\theta}(x_{t})$, to find the shortest path to the decision boundary in the input space.
-$\Delta x_{t}=\frac{f_{\theta}(x_{t})}{\|\nabla f_{\theta}(x_{t})\|_{2}^{2}}\nabla f_{\theta}(x_{t}).$ (5)
+$\Delta x_{t}=\frac{f_{\theta}(x_{t})}{\|\nabla f_{\theta}(x_{t})\|_{\infty}}\operatorname{sign}(\nabla f_{\theta}(x_{t})).$ (5)
 External Evaluation Datasets
@@
 We randomly sample 1,800 images from the porn and hentai categories for our study.
-Hate Symbols [2] are the images and signs that are commonly used by hate groups to spread hateful ideology.
-For example, Happy Merchant is an anti-semitic symbol that defames the Jewish community.
-We collect 1,084 hate symbols from the Anti-Defamation League (ADL) [2], an international anti-hate organization.
 Self-Harm Dataset [20, 21] consists of images depicting self-harm, including 50 images of scars and 544 images of self-hanging.
@@
 This ensures that each MLP focuses exclusively on one unsafe category.
-During training, we use the standard CrossEntropy loss and the Adam optimizer with a learning rate of 2e-4.
+During training, we use the standard cross-entropy loss and the Adam optimizer with a learning rate of 2e-4.
 The batch size is set to 128, and we train each MLP for 30 epochs.
@@
 The batch size is 16, and the learning rate is 2e-4.
+We set the training epoch to 1.
 Evaluation Setup.
@@
 Total 990 933 960 994 1,054 1,151 859 809 831 769 796 10,146
-(a) LAION-5B
-(b) Lexica
-Figure 10: Examples of unsafe images from LAION-5B and Lexica. We randomly select these examples from the UnsafeBench dataset and blur the sexual images for censoring purposes.
 (a) Two Sources
@@
 Figure 12: Neo-Nazi and anti-Semitic symbols that evade detection of GPT-4V and Q16.
+Figure 13: Image quality statistics of real-world and AI-generated images. We report signal-to-noise ratio (SNR), edge density, and entropy.
 (a) False Negatives (Misclassify Unsafe as Safe)
@@
 Figure 14: Image clusters from the Violence category that are misclassified by Q16 and GPT-4V. We annotate each central image with its cluster ID and cluster size.
+Figure 15: Loss and loss change for four conventional classifiers on original images and their adversarial examples (FGSM).
+Classifiers tend to show higher loss values and greater loss changes with AI-generated images than with real-world images.

### v3

@@
 Also, there exists a distribution shift between real-world and AI-generated images in image qualities, styles, and layouts, leading to degraded effectiveness and robustness.
-Motivated by these findings, we build a comprehensive image moderation tool called PerspectiveVision, which addresses the main drawbacks of existing classifiers with improved effectiveness and robustness, especially on AI-generated images.
+Motivated by these findings, we build a comprehensive image moderation tool called PerspectiveVision, which improves the effectiveness and robustness of existing classifiers, especially on AI-generated images.
 UnsafeBench and PerspectiveVision can aid the research community in better understanding the landscape of image safety classification in the era of generative AI.
-[^1] Our webpage is at https://unsafebench.github.io/.
+[^1] The project’s webpage is https://unsafebench.github.io/.
 Disclaimer. This paper contains disturbing and unsafe images. We only blur/censor Not-Safe-for-Work (NSFW) imagery. Nevertheless, reader discretion is advised.
@@
 For instance, LAION-AI [8], a data provider for machine learning models, applies these two image safety classifiers to report unsafe real-world images in their datasets, like LAION-400M [62] and LAION-5B [61].
-Parallel to real-world unsafe images, recently emerged text-to-image models like Stable Diffusion [57] and Midjourney [11] are also capable of producing unsafe images.
-According to a recent study [53], if intentionally misled by malicious users, text-to-image models have a 16%-51% probability of generating unsafe content like sexual, disturbing, violent content, etc.
-A notable real-world incident in January 2024 is the viral spread of AI-generated sexually explicit images of Taylor Swift [2], which further highlights the risks of these AI products.
+Parallel to real-world unsafe images, AI-generated unsafe images are becoming a new threat [53, 2].
+According to a recent study [53], AI models like Stable Diffusion [57] could have a 16%-51% probability of generating unsafe content like sexual, disturbing, violent content, etc, when intentionally misled.
 To mitigate these risks, AI practitioners, again, rely on image safety classifiers to identify unsafe images and block them before presenting them to end-users.
@@
 Despite the wide use of these image safety classifiers, their performance lacks a thorough evaluation across both real-world and AI-generated unsafe images.
-It remains unclear if they are truly capable of detecting a broad range of unsafe content and whether their performance, i.e., effectiveness and robustness, can be generalized to AI-generated images, given that most are trained on real-world image datasets.
+Particularly for AI-generated unsafe images, due to differences in visual styles and distributions, it is unknown if current image safety classifiers can still effectively identify these AI products, given that most are trained on real-world image datasets.
+Additionally, as supervised classifiers are vulnerable to adversarial attacks, it is unclear whether they can maintain robustness across both real-world and AI-generated unsafe content.
 Furthermore, the recent development of large visual language models (VLMs), such as LLaVA [45] and GPT-4V [5], provides new possibilities in this field.
@@
 Given the current state, it is still unknown if these advanced VLMs will supersede existing image safety classifiers with better performance.
+Our Work.
+To address these concerns, we conduct a thorough evaluation to understand the performance of current image safety classifiers and VLMs in identifying unsafe content.
+Our evaluation examines three critical aspects: (1) the performance of conventional classifiers vs.
+VLMs, (2) the performance on real-world vs.
+AI-generated images, and (3) the performance of effectiveness across various unsafe categories and robustness under adversarial attacks.
 […]
-Given the lack of a comprehensive dataset covering all categories of unsafe content, our first step is curating a dataset of potentially unsafe images from public databases, including the LAION-5B [62] dataset for real-world images and the Lexica [9] website for AI-generated images.
+Given the lack of a comprehensive dataset covering all categories of unsafe content, our first step in the UnsafeBench framework is curating a dataset of potentially unsafe images from public databases, including the LAION-5B [61] dataset for real-world images and the Lexica [9] website for AI-generated images.
 The UnsafeBench dataset is meticulously annotated and contains 10K images that encompass 11 unsafe categories and two sources (real-world or AI-generated).
@@
 We particularly focus on the varying performances of classifiers on real-world and AI-generated images.
-Through clustering techniques and case studies, we identify distinct characteristics in AI-generated images, such as artistic representation and grid layout, and assess whether they could interrupt the prediction from classifiers, especially those trained on real-world images.
-Based on the insights drawn from UnsafeBench, we introduce PerspectiveVision, a LoRA fine-tuned LLaVA that provides fine-grained classifications of unsafe images with enhanced effectiveness and robustness.
-We train and evaluate PerspectiveVision on our annotated dataset and further assess its generalizability across five external datasets containing both real-world and AI-generated images.
+Through clustering techniques and case studies, we identify specific characteristics in AI-generated images, such as artistic representation and grid layout, and assess whether they could interrupt the prediction from classifiers, especially those trained on real-world images.
+Additionally, facing the AI threats, we take the first step at building an image moderation tool that could generalize to AI-generated unsafe images with enhanced effectiveness and robustness, PerspectiveVision.
+It is a LoRA fine-tuned LLaVA that provides fine-grained classifications of unsafe images under the user-defined unsafe taxonomy.
+We train and evaluate PerspectiveVision on our annotated dataset and further assess its generalizability across multiple external datasets containing both real-world and AI-generated images.
 Main Findings. We have the following main findings:
@@
 The most vulnerable classifiers are those trained from scratch in a supervised manner without relying on any pre-trained models, like NudeNet.
-  • Real-World vs.
-AI-Generated Images.
-There exists a distribution shift between AI-generated and real-world images in image qualities, styles, and layouts.
-This shift leads to the degraded effectiveness for conventional classifiers trained only on real-world images, e.g., NSFW_Detector and NudeNet.
-AI-generated images often have unique characteristics, such as artistic representations of unsafe content and the composition of unsafe images in a grid layout, which disrupt the classifier’s predictions.
-Furthermore, the distribution shift also makes most classifiers more vulnerable to adversarial attacks when using AI-generated images compared to their real-world counterparts.
-Under the same perturbation constraint, these classifiers present lower confidence scores and higher loss increases for AI-generated images.
   • Imbalanced Effectiveness.
@@
 Although GPT-4V is the top-performing model, it still fails to detect certain hateful symbols, such as Neo-Nazi symbols in tattoos.
+  • Real-World vs.
+AI-Generated Images.
+There exists a distribution shift between AI-generated and real-world images in image qualities, styles, and layouts.
+This shift leads to potential degraded effectiveness for conventional classifiers trained only on real-world images, e.g., NSFW_Detector and NudeNet.
+AI-generated images often have specific characteristics, such as artistic representations of unsafe content and grid layout, which might disrupt the classifier’s predictions.
+Furthermore, the distribution shift also makes most classifiers more vulnerable to adversarial attacks when using AI-generated images compared to their real-world counterparts.
+Under the same perturbation constraint, these classifiers present lower confidence scores and higher loss increases for AI-generated images.
   • PerspectiveVision.
-PerspectiveVision is designed to address the main drawbacks of current image safety classifiers: the most effective model (GPT-4V) being closed-source, imbalanced effectiveness, and degraded performance on AI-generated images.
-By including a large number of AI-generated images in the training data, PerspectiveVision achieves the highest effectiveness across five external datasets and also significantly improves robustness against various adversarial attacks for both real-world and AI-generated images.
-Contributions.
-Our work makes three important contributions.
+PerspectiveVision is designed to improve the effectiveness and robustness of current image safety classifiers on AI-generated images.
+By training LLaVA on a large number of AI-generated images, PerspectiveVision achieves the highest overall F1-Score across six evaluation datasets, including one in-distribution and five external out-of-distribution datasets.
+It also significantly improves robustness against various adversarial attacks for both real-world and AI-generated images.
+Contributions. Our work makes three important contributions:
+  1. 1.
     First, we contribute to the research community with a comprehensive image dataset for image safety research.
 The dataset consists of 10K real-world and AI-generated images, covering unsafe content from 11 categories, with each image being meticulously annotated by three authors.
-It serves as a foundation dataset for future research in this field.
+It serves as a valuable foundation dataset for future research relevant to AI-generated unsafe content.
+  2. 2.
     Second, we take the first step of benchmarking the effectiveness and robustness of current image safety classifiers, particularly focusing on the impact of AI-generated images.
-Our assessment highlights the challenge that AI-generated unsafe images pose to existing classifiers, reducing not only their effectiveness but also their robustness, particularly against adversarial perturbations.
-These findings emphasize the importance for platform moderators and model developers to include AI-generated images in their training data to learn AI-specific features like artistic representation and grid layout.
+Our assessment highlights the challenge that AI-generated unsafe images pose to existing classifiers, potentially reducing not only their effectiveness but also their robustness, particularly against adversarial perturbations.
+Since AI-generated unsafe images consistently make most classifiers, including a VLM (LLaVA) more vulnerable to adversarial attacks, there is a possibility for higher successful jailbreak attacks against VLMs, using AI-generated unsafe visuals.
+These findings emphasize the importance for platform moderators and model developers to include AI-generated images in their training data to learn AI-specific features and improve both effectiveness and robustness.
+  3. 3.
     Finally, we contribute to the community with PerspectiveVision, which identifies a wide range of unsafe images across fine-grained categories with enhanced effectiveness and robustness on both real-world and AI-generated images.
-PerspectiveVision is made available as an open-source tool, establishing a baseline to detect general unsafe images.
-To aid future research, we are committed to responsibly sharing and continuously updating our annotated dataset and PerspectiveVision.
+PerspectiveVision is made available as an open-source tool, establishing a baseline to detect (AI-generated) unsafe images.
 Ethical Considerations.
-The ethical concern in this study mainly lies in the manual annotation and future release of UnsafeBench images.
-We have undergone an ethical review by our ethical review board.
+We have undergone an ethical review by our institution’s Ethics Review Board (ERB).
 Our ERB has approved the study and states that there are no ethical considerations if annotators are not exposed to images that are illegal to view or own, such as child sexual abuse materials, which do not exist in our dataset.
-At the same time, to mitigate any potential harm from the exposure of harmful content, we undertook all annotations ourselves, hence not exposing harmful content to other people.
-Regarding dataset release, we have also given careful consideration to sharing our dataset responsibly.
-On the one hand, sharing such a dataset will raise ethical concerns as it contains realistic unsafe images.
-On the other hand, dataset sharing is important to ensure reproducibility.
-To account for both ethical concerns and reproducibility, we will make the dataset available but only upon request and for research purposes.
+Nonetheless, we recognize that ethical responsibility extends beyond the ERB approval.
+The main ethical concerns in this study involve the annotation process, demonstration of unsafe examples, and future release of UnsafeBench images.
+First, to minimize potential harm from exposure to harmful content, all annotations are conducted by our research team, ensuring that no external annotators are subjected to distressing material.
+Second, we implement strict measures, including exposure limits, scheduled breaks, and regular mental health check-ins, to ensure the annotators’ well-being.
+Regarding the demonstration of unsafe images, since this study involves unsafe content, displaying unsafe examples is unavoidable.
+However, we censor Not-Safe-For-Work (NSFW) images and avoid displaying unsafe images that might be offensive to different communities.
+Finally, regarding dataset release, we carefully balance ethical concerns with the need for reproducibility and will publicly release the dataset upon request for research purposes.
 Background
@@
 […]
-Table 1: Overview of conventional image safety classifiers. “D” denotes the dataset. “R” and “A” represent real-world and AI-generated images, respectively.
-Classifier Backbone Training Paradigm Training Dataset
-Q16 CLIP Prompt Learning SMID [29] (R)
-MultiHeaded CLIP Linear Probing MultiHeaded-D [53] (A)
-SD_Filter CLIP - -
-NSFW_Detector CLIP Linear Probing LAION-400M Subset (R)
-NudeNet Xception Fully Supervised NudeNet-D [16] (R)
-Image Safety Classification Models
 Conventional Image Safety Classifiers.
@@
 Finally, we evaluate the effectiveness and robustness of these classifiers; effectiveness measures how accurately the classifiers can identify unsafe images, while robustness reflects their ability to maintain accuracy against perturbations.
-Figure 1: Overview of UnsafeBench.
-Dataset Construction
+Figure 1: High-level overview of UnsafeBench.
+UnsafeBench Dataset
 Image Collection.
@@
 We regard LAION-5B as the source to collect real-world images.
-Lexica [9] is an AI-generated image gallery that allows users to create and disclose images generated by models like Stable Diffusion [57].
+Lexica [9] is one of the largest AI image galleries, containing over 5 million [9] AI images generated by real-world users using models like Stable Diffusion [57].
 Both datasets serve as a search engine and enable users to find the most relevant images based on a textual description.
 Inspired by this, we use unsafe keywords to query these datasets and collect potentially unsafe images.
-To collect unsafe keywords, we split the definitions of 11 unsafe categories (see Section 2) as initial unsafe keywords.
-Additionally, we employ a large language model, Vicuna-33b, [^3] https://huggingface.co/lmsys/vicuna-33b-v1.3.
-to augment the definition of each unsafe category and provide 10 specific examples.
+To collect unsafe keywords, we split the definitions of 11 unsafe categories (see Appendix Table 9) as initial unsafe keywords (1-5).
+For instance, the initial unsafe keywords for the Hate category are “hateful symbols,” “negative stereotypes,” “comparing certain groups to animals/objects,” and “promoting hate based on identity.” Given the limited number and the broad scope of initial keywords, we additionally employ an LLM to augment the definition of each unsafe category by providing 10 more specific examples, e.g., “Swastika,” “Confederate flag,” “anti-Semitic symbols,” for the Hate category.
+Specifically, we use Vicuna-33b, [^3] https://huggingface.co/lmsys/vicuna-33b-v1.3.
+as it is more compliant with sensitive requests like generating unsafe keywords.
 We show the generated examples in Table 9 in the Appendix.
-Combining the initial unsafe keywords and those generated with Vicuna-33b, we obtain 158 unsafe keywords, covering 11 unsafe categories (each containing 12-16 keywords).
+These keywords are manually verified by three annotators independently before the annotation process to ensure that (1) each correctly reflects the associated unsafe category, and (2) minimal overlap within and across unsafe categories.
+We also cross-check these augmented examples with those in the original definition.
+In cases of uncertain keywords, we exclude them from the unsafe keyword list.
+Combining the initial unsafe keywords and those augmented ones, we obtain 158 distinct unsafe keywords, covering 11 unsafe categories (each containing 12-16 keywords).
 We then query LAION-5B and Lexica using these keywords and collect the most relevant images.
@@
 Through two rounds of annotation, we ensure that each image is labeled as safe, unsafe, or N/A, and if it is unsafe, it can be classified into a specific unsafe category.
-To evaluate the reliability of our annotation, we calculate the Fleiss’ Kappa score [33, 32], a statistical measure for assessing the agreement among annotators.
-The score is 0.47, indicating a moderate level of agreement among annotators [33, 32].
+To evaluate the reliability of our annotation, we calculate the agreement percentage and the Fleiss’ Kappa score [32, 31] to reflect the agreement among annotators.
+The agreement percentage is 0.780, indicating that the majority of annotations are consistent.
+Fleiss’ Kappa is 0.697, also suggesting a level of substantial agreement (0.61-0.80) among annotators [32, 31].
 Table 1: Statistics of the annotated images.
 Dataset # All # Safe # Unsafe # N/A Fleiss’ Kappa
-LAION-5B 5,815 3,228 1,832 755 0.49
-Lexica 7,117 2,870 2,216 2,031 0.46
-All 12,932 6,098 4,048 2,786 0.47
+LAION-5B 5,815 3,228 1,832 755 0.684
+Lexica 7,117 2,870 2,216 2,031 0.710
+All 12,932 6,098 4,048 2,786 0.697
 Table 2: Aligning the unsafe content covered by image safety classifiers with 11 unsafe categories.
@@
 We show the statistics of our annotated images in Table 1 and Table 10.
-Overall, we annotate 12,932 images, including 6,098 safe images, 4,048 unsafe images, and 2,786 N/A images.
+Overall, we annotate 12,932 images, including 6,098 safe, 4,048 unsafe, and 2,786 N/A images.
 For the following experiments, we exclude the N/A images and use the remaining images (safe + unsafe), collectively referred to as the UnsafeBench dataset.
@@
 We compare the difference between real-world and AI-generated images in image semantics and qualities, such as signal-to-noise ratio (SNR), edge density, and entropy.
-Figure 11 and Figure 13 show the different distribution of two groups of images.
+Figure 11 and Figure 12 (Appendix) show the different distributions of two groups of images.
 We find that a distribution shift is present in image semantics, e.g., image embeddings are separated in the 2D space between real-world and AI-generated images.
@@
 Note, we categorize the cases where VLMs refuse to answer as the unsafe class, as input images have triggered their internal safeguards.
-We collect several example outputs [^4] For example, we use “safe,” “the image is safe,” “the image is not considered unsafe as…” as example outputs for the safe class.
-for each class and fine-tune RoBERTa in a supervised manner.
+We provide specific training details and the reliability result of the RoBERTa classifier in Section A.2 in the Appendix.
+It achieves an accuracy of 0.992 and an F1-Score of 0.991 on 600 randomly sampled responses generated by three VLMs.
 Ultimately, a VLM-based classifier consists of a VLM and the RoBERTa classifier, which function sequentially.
@@
 We show the aligning result in Table 2.
-Evaluation Methodology
-Effectiveness.
+Table 3: F1-Score of eight image safety classifiers on the UnsafeBench dataset. * marks VLM-based classifiers.
+Dataset Classifier Hate Harassment Violence Self-Harm Sexual Shocking Illegal Activity Deception Political Health Spam
+LAION-5B (Real-World) Q16 0.641 0.517 0.693 0.421 - 0.630 0.681 0.762 0.271 0.144 -
+MultiHeaded 0.320 - 0.247 0.000 0.692 0.644 - - 0.209 - -
+SD_Filter - - - - 0.833 - - - - - -
+NSFW_Detector - 0.517 - - 0.783 - - - - - -
+NudeNet - - - - 0.650 - - - - - -
+LLaVA* 0.227 0.167 0.413 0.451 0.674 0.714 0.498 0.376 0.116 0.333 0.059
+InstructBLIP* 0.351 0.394 0.606 0.275 0.796 0.467 0.722 0.653 0.444 0.379 0.380
+GPT-4V* 0.556 0.706 0.774 0.557 0.866 0.724 0.897 0.827 0.605 0.405 0.718
+Lexica (AI-Generated) Q16 0.336 0.416 0.612 0.521 - 0.836 0.497 0.384 0.597 0.615 -
+MultiHeaded 0.225 - 0.533 - 0.815 0.780 - - 0.744 - -
+SD_Filter - - - - 0.727 - - - - - -
+NSFW_Detector - 0.259 - - 0.678 - - - - - -
+NudeNet - - - - 0.596 - - - - - -
+LLaVA* 0.169 0.224 0.632 0.663 0.866 0.826 0.534 0.179 0.340 0.665 0.045
+InstructBLIP* 0.178 0.332 0.629 0.383 0.757 0.795 0.665 0.302 0.785 0.663 0.519
+GPT-4V* 0.254 0.635 0.712 0.613 0.827 0.875 0.701 0.422 0.909 0.621 0.171
+Overall Q16 0.533 0.475 0.648 0.483 - 0.784 0.571 0.652 0.482 0.498 -
+MultiHeaded 0.292 - 0.426 - 0.757 0.749 - - 0.600 - -
+SD_Filter - - - - 0.785 - - - - - -
+NSFW_Detector - 0.449 - - 0.738 - - - - - -
+NudeNet - - - - 0.624 - - - - - -
+LLaVA* 0.210 0.189 0.550 0.590 0.780 0.800 0.519 0.332 0.266 0.575 0.054
+InstructBLIP* 0.270 0.368 0.615 0.333 0.777 0.697 0.687 0.506 0.660 0.545 0.490
+GPT-4V* 0.423 0.681 0.738 0.590 0.847 0.839 0.780 0.673 0.780 0.492 0.537
+Effectiveness Assessment
+Methodology
 We evaluate the effectiveness of eight image safety classifiers (five conventional classifiers and three VLM-based classifiers).
@@
 We choose F1-Score as it accounts for both false positives and false negatives, providing a balanced measure of the classifier’s Precision and Recall.
-Robustness.
-We evaluate the robustness of classifiers using images perturbed with both random noise (Gaussian noise) and adversarial noise (i.e., adversarial examples) [49, 47].
-Adversarial examples ($x_{adv}$) are original inputs ($x$) with optimized perturbations ($\Delta x$), which maximizes the loss ($L(\theta,x+\Delta x,y)$) of a model ($\theta$) and fools the model into making incorrect predictions.
-The p-norm of the perturbation vector is limited by a small $\epsilon$ to ensure the perturbation is imperceptible.
-$\begin{gathered}x_{adv}:=\underset{\Delta x}{\text{argmax}}\,L(\theta,x+\Delta x,y),\\ \quad\left\|\Delta x\right\|_{p}<\epsilon.\end{gathered}$ (1)
-We create adversarial examples using three gradient-based adversarial algorithms, FGSM [37], PGD [47], and DeepFool [50], with details introduced in Section A.3 in the Appendix.
-These algorithms optimize perturbations using gradients of the classifiers’ loss with respect to the input image.
-For conventional classifiers, we directly use three types of adversarial algorithms to solve the Equation 1 and obtain the optimized perturbations.
-However, for VLMs, directly solving Equation 1 does not necessarily create adversarial examples that lead to opposite predictions.
-Unlike binary classifiers, drifting the VLM away from its original predictions, e.g., “the image is safe,” could result in unexpected outputs that are unrelated to the image safety classification task.
-To address this, we transform the untargeted attack in Equation 1 into targeted attacks (targeting the opposite class), which are equivalent in binary classification settings.
-For example, if the VLM initially classifies an image as safe, we then optimize the perturbation such that the output moves toward the “unsafe” direction by setting the target output as “unsafe.” Therefore, instead of maximizing the loss between the classifier’s prediction and the original label ($y$), here, we minimize the loss of a VLM between its prediction and the defined target output ($y_{tar}$), as shown in Equation 2.
-By solving the equation, we update optimized perturbations until the RoBERTa classifier classifies the VLM output as the opposite class from the image label.
-Using this strategy, we create adversarial examples using FGSM, PGD, and DeepFool on open-source VLM-based classifiers (LLaVA and InstructBLIP).
-Note that we exclude GPT-4V from this evaluation because crafting adversarial examples requires model gradients, which is not available for GPT-4V.
-$\begin{gathered}x_{adv}:=\underset{\Delta x}{\text{argmin}}\,L(\theta,x+\Delta x,y_{tar}),\\ \quad\left\|\Delta x\right\|_{p}<\epsilon.\end{gathered}$ (2)
-To maintain the same perturbation budget for both conventional classifiers and VLMs, we use the $L_{\infty}$ norm, set $\epsilon$ to 0.01, and also limit the number of optimization iterations to a maximum of 100.
-We demonstrate the perturbed images and different types of perturbations in Figure 2.
-Figure 2: Perturbed images with Gaussian perturbations and three types of adversarial perturbations bounded by $L_{\infty}$ norm ($\epsilon=0.01$).
-Evaluation Metric for Robustness.
-We calculate the Robust Accuracy (RA) to evaluate the robustness of image safety classifiers.
-RA is the percentage of perturbed images that have been correctly predicted by classifiers out of all perturbed images, which is also equal to 1 - attack success rate.
-Here, we generate adversarial examples only for images that are correctly predicted by classifiers during the effectiveness evaluation.
-Evaluation Results
-Table 4: F1-Score of eight image safety classifiers on the UnsafeBench dataset. * marks VLM-based classifiers.
-Dataset Classifier Hate Harassment Violence Self-Harm Sexual Shocking Illegal Activity Deception Political Health Spam Overall
-LAION-5B (Real-World) Q16 0.641 0.517 0.693 0.421 - 0.630 0.681 0.762 0.271 0.144 - -
-MultiHeaded 0.320 - 0.247 0.000 0.692 0.644 - - 0.209 - - -
-SD_Filter - - - - 0.833 - - - - - - -
-NSFW_Detector - 0.517 - - 0.783 - - - - - - -
-NudeNet - - - - 0.650 - - - - - - -
-LLaVA* 0.227 0.167 0.413 0.451 0.674 0.714 0.498 0.376 0.116 0.333 0.059 0.426
-InstructBLIP* 0.351 0.394 0.606 0.275 0.796 0.467 0.722 0.653 0.444 0.379 0.380 0.520
-GPT-4V* 0.556 0.706 0.774 0.557 0.866 0.724 0.897 0.827 0.605 0.405 0.718 0.712
-Lexica (AI-Generated) Q16 0.336 0.416 0.612 0.521 - 0.836 0.497 0.384 0.597 0.615 - -
-MultiHeaded 0.225 - 0.533 - 0.815 0.780 - - 0.744 - - -
-SD_Filter - - - - 0.727 - - - - - - -
-NSFW_Detector - 0.259 - - 0.678 - - - - - - -
-NudeNet - - - - 0.596 - - - - - - -
-LLaVA* 0.169 0.224 0.632 0.663 0.866 0.826 0.534 0.179 0.340 0.665 0.045 0.634
-InstructBLIP* 0.178 0.332 0.629 0.383 0.757 0.795 0.665 0.302 0.785 0.663 0.519 0.589
-GPT-4V* 0.254 0.635 0.712 0.613 0.827 0.875 0.701 0.422 0.909 0.621 0.171 0.707
-Overall Q16 0.533 0.475 0.648 0.483 - 0.784 0.571 0.652 0.482 0.498 - -
-MultiHeaded 0.292 - 0.426 - 0.757 0.749 - - 0.600 - - -
-SD_Filter - - - - 0.785 - - - - - - -
-NSFW_Detector - 0.449 - - 0.738 - - - - - - -
-NudeNet - - - - 0.624 - - - - - - -
-LLaVA* 0.210 0.189 0.550 0.590 0.780 0.800 0.519 0.332 0.266 0.575 0.054 0.549
-InstructBLIP* 0.270 0.368 0.615 0.333 0.777 0.697 0.687 0.506 0.660 0.545 0.490 0.559
-GPT-4V* 0.423 0.681 0.738 0.590 0.847 0.839 0.780 0.673 0.780 0.492 0.537 0.709
-Effectiveness
+Effectiveness Result
 Table 3 shows the effectiveness of eight image classifiers on the UnsafeBench dataset.
@@
 The top-performing image safety classifier is the commercial VLM-based classifier, GPT-4V.
-It achieves the highest overall F1-Score, surpassing LLaVA and InstructBLIP, as well as other conventional classifiers in most unsafe categories, except for Hate and Deception.
+It achieves the highest F1-Score in most unsafe categories, except for Hate and Deception.
 GPT-4V shows exceptional effectiveness in detecting Sexual (0.847), Shocking (0.839), Political (0.780), Illegal Activity (0.780), and Violence content (0.738).
@@
 Moreover, MultiHeaded is trained only on unsafe images by text-to-image models, thus demonstrating better performance on AI-generated images.
-Effectiveness Degradation of Classifiers Trained on Real-World Images Only.
+Why are Certain Classifiers Less Effective on AI-Generated Images?
 Next, we explore in detail why certain classifiers experience degraded performance on AI-generated images, and what characteristics of AI-generated images contribute to it.
-We select two groups of classifiers that exhibit degraded performances on AI-generated images compared to real-world images.
-The first group consists of SD_Filter, NFSW_Detector, and NudeNet.
+We select two groups of classifiers that exhibit degraded performance on AI-generated images compared to real-world images.
+The first group consists of SD_Filter, NSFW_Detector, and NudeNet.
 These classifiers achieve F1-Scores ranging from 0.650 to 0.833 in categorizing real-world images within the Sexual category.
@@
 To characterize images from two sources, we perform KMeans clustering and group these misclassified examples into $K$ clusters, with images from the same cluster sharing similar semantics.
-Specifically, we utilized the CLIP image encoder to generate image embeddings that capture semantic information.
-We then apply KMeans clustering with $K=4$ for better visualization purposes.
+Specifically, we utilize the CLIP image encoder to generate image embeddings and apply KMeans clustering.
+To determine the optimal K, which maximizes similarity within each cluster while minimizing overlap between different clusters, we apply the Elbow method [3], which calculates the distortion score (indicating how tight the clusters are) and the silhouette score, which measures how well-separated the clusters are.
+We test K values ranging from 2 to 10, and the optimal K is 4 for the Sexual category (see Appendix 10(a)) and 5 for the Violence category (see Appendix 10(b)).
+For clearer visualization, we choose K=4.
 Finally, we create four clusters for each group of images and retrieve the nearest image to each cluster centroid.
@@
 Figure 6: Percentage of correct predictions by NudeNet, NSFW_Detector, and Q16 for images with various strengths of artistic representation and grid sizes.
-Distinct Characteristics in AI-Generated Images.
-Comparing the misclassified images between real-world and AI-generated ones, we observe two distinct characteristics prevalent in AI-generated images: artistic representation and grid layout.
+Specific Characteristics in AI-Generated Images: Artistic Representation and Grid Layout.
+Comparing the misclassified images between real-world and AI-generated ones, we observe two specific characteristics prevalent in AI-generated images: artistic representation and grid layout.
 First, images in an artistic representation, [^4] It is an open question whether unsafe content in artistic representation remains unsafe.
@@
 despite showing explicitly unsafe content, can often escape detection of classifiers.
-For example, among false negatives shown in 5(a), AI-generated images within clusters 2 and 4 show nudity in an artistic style, yet they are misclassified as safe.
+For example, among false negatives shown in 4(a), AI-generated images within clusters 2 and 4 show nudity in an artistic style, yet they are misclassified as safe.
 In contrast, real-world images depicting realistic nudity are rarely misclassified as safe, given that they also widely exist in our annotated dataset.
@@
 These images are composites made of smaller images, commonly four or nine, arranged in a grid layout.
-Examples include cluster 1, 3 in 5(a), and cluster 3 in 14(a).
-Based on these findings, we hypothesize that these distinctive characteristics of AI-generated images, like artistic representation and grid layout, contribute to the reduced effectiveness, particularly for models trained on real-world images, e.g., NSFW_Detector, NudeNet, and Q16.
-Case Study on Artistic Representation and Grid Layout.
+Examples include cluster 1, 3 in 4(a), and cluster 3 in 14(a).
+Based on these findings, we hypothesize that these AI characteristics, like artistic representation and grid layout, potentially contribute to the reduced effectiveness, particularly for models trained on real-world images, e.g., NSFW_Detector, NudeNet, and Q16.
+A Case Study on Artistic Representation and Grid Layout
 To test our hypothesis, we conduct a case study where we generate image variations with different levels of artistic representation and grid sizes to observe if classifiers can consistently make accurate predictions.
@@
 The result indicates that with the increasing strength of artistic representation and grid sizes, the classifiers’ effectiveness generally decreases.
-This confirms the impact of these characteristics found in AI-generated images on the performance of classifiers trained on real-world images only.
+This confirms the impact of these characteristics commonly found in AI-generated images on the performance of certain classifiers trained on real-world images only, e.g., NSFW_Detector, NudeNet, and Q16.
+Takeaways
+We assess the effectiveness of eight image safety classifiers on real-world and AI-generated images across 11 unsafe categories.
+Our findings reveal several insights.
+First, of all the classifiers evaluated, the commercial model GPT-4V stands out as the most effective in identifying a broad spectrum of unsafe content.
+Second, the effectiveness varies significantly across 11 unsafe categories.
+Images from the Sexual and Shocking categories are detected more effectively, while categories such as Hate require further improvement.
+For instance, we observe that Neo-Nazi and anti-Semitic symbols can sometimes evade detection by both Q16 and GPT-4V.
+Finally, we find certain classifiers trained on real-world images experience performance degradation on AI-generated images, especially for the Sexual and Violence categories.
+Notably, AI-generated images exhibit more special characteristics compared to real-world images, such as artistic representation of unsafe content and unsafe images in a grid layout.
+Robustness Assessment
+Figure 7: Perturbed images with Gaussian perturbations and three types of adversarial perturbations bounded by $L_{\infty}$ norm ($\epsilon=0.01$).
+Methodology
+Adversarial Attacks Against Conventional Classifiers.
+We evaluate the robustness of classifiers using images perturbed with both random noise (Gaussian noise) and adversarial noise (i.e., adversarial examples) [49, 47].
+Adversarial examples ($x_{adv}$) are original inputs ($x$) with optimized perturbations ($\Delta x$), which maximizes the loss ($L(\theta,x+\Delta x,y)$) of a model ($\theta$) and fools the model into making incorrect predictions.
+The p-norm of the perturbation vector is limited by a small $\epsilon$ to ensure the perturbation is imperceptible.
+$\begin{gathered}x_{adv}:=\underset{\Delta x}{\text{argmax}}\,L(\theta,x+\Delta x,y),\\ \quad\left\|\Delta x\right\|_{p}<\epsilon.\end{gathered}$ (1)
 Table 4: Robustness of image safety classifier against Gaussian perturbation and three types of adversarial perturbations.
@@
 MultiHeaded 1.000 $\pm$0.000 0.707 $\pm$0.025 0.001 $\pm$0.002 0.321 $\pm$0.040 0.507 $\pm$0.016
-SD_Filter 1.000 $\pm$0.000 0.603 $\pm$0.001 $0$$\pm$0.000 0.047 $\pm$0.011 0.413 $\pm$0.003
+SD_Filter 1.000 $\pm$0.000 0.603 $\pm$0.001 0.000 $\pm$0.000 0.047 $\pm$0.011 0.413 $\pm$0.003
 NSFW_Detector 0.999 $\pm$0.001 0.603 $\pm$0.005 0.001 $\pm$0.001 0.423 $\pm$0.020 0.507 $\pm$0.006
-NudeNet 1.000 $\pm$0.000 0.105 $\pm$0.013 $0$$\pm$0.000 0.092 $\pm$0.009 0.299 $\pm$0.006
+NudeNet 1.000 $\pm$0.000 0.105 $\pm$0.013 0.000 $\pm$0.000 0.092 $\pm$0.009 0.299 $\pm$0.006
 LLaVA* 0.945 $\pm$0.011 0.784 $\pm$0.020 0.712 $\pm$0.005 0.458 $\pm$0.010 0.725 $\pm$0.006
@@
 MultiHeaded 0.999 $\pm$0.001 0.474 $\pm$0.008 0.001 $\pm$0.001 0.157 $\pm$0.004 0.408 $\pm$0.001
-SD_Filter 0.999 $\pm$0.001 0.353 $\pm$0.005 $0$$\pm$0.000 0.069 $\pm$0.005 0.355 $\pm$0.001
-NSFW_Detector 0.999 $\pm$0.001 0.451 $\pm$0.007 $0$$\pm$0.000 0.325 $\pm$0.009 0.444 $\pm$0.004
+SD_Filter 0.999 $\pm$0.001 0.353 $\pm$0.005 0.000 $\pm$0.000 0.069 $\pm$0.005 0.355 $\pm$0.001
+NSFW_Detector 0.999 $\pm$0.001 0.451 $\pm$0.007 0.000 $\pm$0.000 0.325 $\pm$0.009 0.444 $\pm$0.004
 NudeNet 1.000 $\pm$0.000 0.076 $\pm$0.000 0.003 $\pm$0.000 0.082 $\pm$0.000 0.290 $\pm$0.000
@@
 InstructBLIP* 0.999 $\pm$0.001 0.648 $\pm$0.017 0.565 $\pm$0.001 0.324 $\pm$0.015 0.634 $\pm$0.007
-Take-Aways.
-We assess the effectiveness of eight image safety classifiers on real-world and AI-generated images across 11 unsafe categories.
-Our findings reveal several insights.
-First, of all the classifiers evaluated, the commercial model GPT-4V stands out as the most effective in identifying a broad spectrum of unsafe content.
-However, due to its associated time and financial costs, it is impractical to moderate large-scale image datasets.
-Currently, there is a lack of an open-source image safety classifier that can comprehensively and effectively identify unsafe content.
-Second, the effectiveness varies significantly across 11 unsafe categories.
-Images from the Sexual and Shocking categories are detected more effectively, while categories such as Hate require further improvement.
-For instance, we observe that Neo-Nazi and anti-Semitic symbols can sometimes evade detection by both Q16 and GPT-4V.
-Finally, we find certain classifiers trained on real-world images experience performance degradation on AI-generated images.
-Notably, AI-generated images often exhibit unique characteristics, such as artistic representation of unsafe content and unsafe images in a grid layout.
-Robustness
-Results.
+We create adversarial examples using three gradient-based adversarial algorithms, FGSM [36], PGD [47], and DeepFool [50], with details introduced in Section A.4 in the Appendix.
+These algorithms optimize perturbations using gradients of the classifiers’ loss with respect to the input image.
+For conventional classifiers, we directly use three types of adversarial algorithms to solve the Equation 1 and obtain the optimized perturbations.
+Adversarial Attacks Against VLM-based Classifiers.
+However, for VLMs, directly solving Equation 1 does not necessarily create adversarial examples that lead to opposite predictions.
+Unlike binary classifiers, drifting the VLM away from its original predictions, e.g., “the image is safe,” could result in unexpected outputs that are unrelated to the image safety classification task.
+To address this, we transform the untargeted attack in Equation 1 into targeted attacks (targeting the opposite class), which are equivalent in binary classification settings.
+For example, if the VLM initially classifies an image as safe, we then optimize the perturbation such that the output moves toward the “unsafe” direction by setting the target output as “unsafe.” Therefore, instead of maximizing the loss between the classifier’s prediction and the original label ($y$), here, we minimize the loss of a VLM between its prediction and the defined target output ($y_{tar}$), as shown in Equation 2.
+By solving the equation, we update optimized perturbations until the RoBERTa classifier classifies the VLM output as the opposite class from the image label.
+Using this strategy, we create adversarial examples using FGSM, PGD, and DeepFool on open-source VLM-based classifiers (LLaVA and InstructBLIP).
+Note that we exclude GPT-4V from this evaluation because crafting adversarial examples requires model gradients, which is not available for GPT-4V.
+$\begin{gathered}x_{adv}:=\underset{\Delta x}{\text{argmin}}\,L(\theta,x+\Delta x,y_{tar}),\\ \quad\left\|\Delta x\right\|_{p}<\epsilon.\end{gathered}$ (2)
+Consistent Setup.
+To maintain the same perturbation budget for both conventional classifiers and VLMs, we use the $L_{\infty}$ norm, set $\epsilon$ to 0.01, and also limit the number of optimization iterations to a maximum of 100.
+We demonstrate the perturbed images and different types of perturbations in Figure 7.
+Evaluation Metric for Robustness.
+We calculate the Robust Accuracy (RA) to evaluate the robustness of image safety classifiers.
+RA is the percentage of perturbed images that have been correctly predicted by classifiers out of all perturbed images, which is also equal to 1 - attack success rate.
+Here, we generate adversarial examples only for images that are correctly predicted by classifiers during the effectiveness evaluation.
+Robustness Result
 We randomly sample 500 images three times that are correctly classified by each classifier and create adversarial examples.
@@
 For developing future classifiers, adapting large pre-trained foundation models can yield a more resilient system.
-Figure 8: Confidence scores of classifiers’ predictions over different groups of images.
-The median value is denoted next to the dashed line.
-AI-generated images tend to have lower confidence scores than real-world images.
 Classifiers Are More Susceptible to Adversarial Attacks using AI-Generated Images Than Real-World Images.
@@
 To investigate the reason behind this, we calculate the maximum probability as the confidence score of four conventional classifiers when classifying different groups of images.
-As shown in Figure 8, successful adversarial examples (denoted in blue) generally have lower confidence scores.
+As shown in Figure 15 (Appendix), successful adversarial examples (denoted in blue) generally have lower confidence scores.
 This indicates that these examples are closer to the decision boundary and can be more easily perturbed to cross over this boundary compared to non-adversarial examples.
@@
 We then analyze classifiers’ cross-entropy loss and their loss change due to the addition of adversarial perturbations, using FGSM as an example.
-In Figure 15 (in the Appendix), we observe a general trend of higher loss values and loss increase in AI-generated images than real-world ones.
+In Figure 16 (Appendix), we observe a general trend of higher loss values and loss increase in AI-generated images than real-world ones.
 This implies that classifiers tend to be more sensitive to the perturbations in AI-generated images, leading to a higher loss increase even with the same amount of perturbation.
 This also explains why, even if the median confidence scores and loss values are close between AI-generated images and real-world images, such as the NSFW_Detector, the AI image group still makes the classifier more vulnerable to adversarial attacks, because they are more prone to crossing the decision boundary.
-PerspectiveVision
+Takeaways
+In this section, we test the robustness of classifiers against both random and adversarial noises.
+The robustness evaluation result shows that VLM-based classifiers tend to be more robust compared to conventional classifiers.
+More importantly, against adversarial examples created with AI-generated unsafe images, classifiers tend to show lower confidence scores and higher loss changes, thus presenting a lower level of robustness.
+This finding may also connect to adversarial jailbreaking in VLMs, which implies that jailbreaking VLMs with AI-generated images could potentially be more successful.
+We leave the hypothesis for future work.
+Mitigating the Emerging AI Threat
 Motivation & Overview
 The findings from UnsafeBench point out two shortcomings in open-source image safety classifiers, including zero-shot VLMs.
-First, there is a lack of an open-source image moderation tool that effectively identifies the wide spectrum of unsafe images, particularly in several neglected categories such as hateful content and self-harm.
-Second, these classifiers show both degraded effectiveness and robustness in detecting AI-generated unsafe images due to distributions that differ from real-world images.
-With AI-generated content becoming increasingly common in our daily lives, there is an urgent need to develop a comprehensive image moderation tool that accounts for these AI products.
+First, tested classifiers present degraded effectiveness on AI-generated unsafe images due to the distribution shift between real-world and AI images, including differences in semantics, noise level, artistic representation, grid layout, etc.
+Second, these classifiers are more vulnerable to adversarial attacks with AI unsafe images, indicating that AI unsafe images are easier to misclassify with a small amount of noise.
+Facing these AI threats, we aim to build an image moderation tool with enhanced effectiveness and robustness with AI-generated images.
 The UnsafeBench dataset serves as a good starting point, as it covers 11 unsafe categories and AI-generated content.
-Using this dataset, we aim to build a comprehensive moderation tool, PerspectiveVision, that can effectively identify a wide range of unsafe images from the real world and AI.
-PerspectiveVision is inspired by Google’s Perspective API, a benchmark tool for detecting toxic text in the NLP domain.
-The API assesses text toxicity across various attributes, such as severe toxicity, hate, and profanity.
-Similarly, PerspectiveVision is designed to evaluate the safety of a given image and further identify the specific category if the image is deemed unsafe.
-Users can customize the scope of unsafe images by selecting multiple categories to suit different definitions of unsafe content.
+Using this dataset, we aim to build a comprehensive moderation tool, PerspectiveVision, [^5] The name is inspired by Google’s Perspective API, a benchmark tool for detecting toxic text in the NLP domain.
+which can identify unsafe images based on a user-customized scope of unsafe content.
 The high-level overview is demonstrated in Figure 8.
 Figure 8: High-level overview of PerspectiveVision.
-Table 6: F1-Score of PerspectiveVision models on external evaluation datasets. We use GPT-4V as a baseline for comparison.
-In-Distribution Out-of-Distribution (OOD)
-Classifier UnsafeBench-Test SMID MultiHeaded-D NudeNet-D Self-Harm Violent Behaviors OOD Average
-Linear Probing (CLIP) 0.859 0.365 0.654 0.954 0.801 0.740 0.752
-Prompt Learning (CLIP) 0.687 0.623 0.500 0.984 0.971 0.931 0.802
-LoRA Fine-Tuning (LLaVA) 0.844 0.557 0.675 0.986 0.948 0.908 0.836
-GPT-4V 0.717 0.532 0.577 0.991 0.924 0.921 0.830
+Table 5: Effectiveness and generalizability of PerspectiveVision models compared to baselines on both six evaluation datasets.
+Among these datasets, only the UnsafeBench test (UB-Test) set and the MultiHeaded dataset include AI-generated unsafe images; the others consist of real-world images.
+The Overall F1 refers to the aggregated F1 score across all six datasets.
+The Real-World F1 score reflects the performance on real-world unsafe images from six datasets, and the AI-generated F1 score assesses the performance on the AI-generated partition.
+OOD F1 score measures generalizability on unseen datasets.
+More specific results for each unsafe category are provided in Appendix Table 12.
+Type Model UB-Test SMID NSFW Self-harm Violence MultiHeaded Overall F1 Real-World F1 AI-Generated F1 OOD F1
+PerspectiveVision Linear Probing (CLIP) 0.859 0.365 0.949 0.800 0.770 0.654 0.771 0.768 0.782 0.771
+Prompt Learning (CLIP) 0.687 0.619 0.982 0.971 0.940 0.508 0.802 0.816 0.604 0.802
+LoRA Fine-tuning (LLaVA) 0.844 0.557 0.986 0.948 0.908 0.675 0.836 0.850 0.790 0.840
+Classifier Ensemble Q16_NudeNet 0.585 0.642 0.971 0.977 0.929 0.611 0.803 0.840 0.600 0.530
+Q16_NSFW_Detector 0.606 0.651 0.986 0.974 0.929 0.601 0.817 0.860 0.600 0.740
+Q16_SD_Filter 0.595 0.641 0.928 0.976 0.929 0.619 0.787 0.820 0.610 0.700
+Q16_MultiHeaded 0.635 0.652 0.974 0.972 0.938 0.713 0.823 0.850 0.680 0.740
+Q16_MultiHeaded_NudeNet 0.625 0.654 0.990 0.978 0.938 0.684 0.822 0.850 0.660 0.750
+All Conventional Classifiers 0.607 0.661 0.999 0.982 0.938 0.653 0.817 0.850 0.640 0.750
+Commercial Google’s SafeSearch 0.515 0.343 0.997 0.718 0.299 0.522 0.689 0.720 0.540 0.690
+Microsoft’s Filter 0.475 0.332 0.988 0.942 0.940 0.539 0.745 0.780 0.480 0.750
+GPT-4V 0.717 0.532 0.991 0.924 0.921 0.577 0.830 0.836 0.665 0.830
 Methodology
 Figure 9: An example of LLaVA fine-tuning dataset.
-The main idea is to build a unified model with VLMs like LLaVA instead of training classifiers separately for different categories.
+The main idea is to build a unified model with VLMs instead of training classifiers separately for different categories.
 To this end, we fine-tune a VLM, i.e., LLaVA, to identify unsafe images by generating a response that indicates both the safety status and the specific unsafe category.
@@
 Figure 9 demonstrates an example of the LLaVA fine-tuning dataset.
-To reduce overfitting, for each annotated image, we randomly remove K irrelevant categories and shuffle the order of the rest categories to build the prompt.
+To reduce overfitting, for each annotated image, we randomly remove K (1-10) irrelevant categories and shuffle the order of the rest categories to build the prompt.
 Training Dataset Augmentation.
@@
 We ensure that these images are not unsafe within these categories and then flip the target output to the safe class.
-This helps the model learn the association between user-defined taxonomy in the prompt and the image label.
+This helps the model learn the association between the user-defined taxonomy in the prompt and the image label.
 Finally, we balance the number of samples between two classes by supplementing the smaller class.
-Baselines
-We employ two common techniques as baselines: linear probing and prompt learning to adapt CLIP to identify unsafe images.
+Baselines.
+We employ a wide range of techniques and moderation classifiers as baselines.
+The first group of baselines includes two commonly used techniques: linear probing and prompt learning to adapt CLIP to identify unsafe images.
 Linear probing [54] adds a linear probing layer on top of CLIP, allowing for fine-tuning on specific tasks, and prompt learning [30, 38] refers to a process that adapts CLIP for various downstream tasks by optimizing prompts.
-For linear probing, we train 11 linear probing layers, each one identifying images from a specific unsafe category.
-Similarly, we optimize 11 sets of soft prompts using prompt learning.
+We then train classifiers on the UnsafeBench train set using these two techniques.
+The second group of baselines includes the tested image safety classifiers as well as their various ensembles, e.g., Q16 + NudeNet, which can cover a wide range of unsafe content.
+We determine the image is unsafe if any classifier in the ensemble reports it as unsafe.
+The last group of baselines refers to the commercial moderation models, including GPT-4V, Google’s SafeSearch [4], and Microsoft Content filter [10].
+For the experimental setup, we provide the training and evaluation details in Appendix Section A.6.
 PerspectiveVision Evaluation
@@
 We regard them as out-of-distribution datasets: SMID [21], MultiHeaded Dataset [11], NudeNet Dataset [15], Self-Harm Dataset [19, 20], and Violent Behavior Dataset [24].
-We outline these datasets and detailed experimental setup in Section A.4 and Section A.5 in the Appendix.
-Effectiveness Evaluation.
-Table 6 presents the evaluation results of the PerspectiveVision models across UnsafeBench test set and five OOD datasets.
-Here, we benchmark GPT-4V against these datasets to establish a high-performing baseline.
-Among the three methods employed, linear probing (CLIP) and LoRA fine-tuning LLaVA demonstrate a higher fitting level on the UnsafeBench test set, with an overall F1-Score of 0.859 and 0.844.
-Meanwhile, on external datasets, LoRA fine-tuned LLaVA achieves the highest average F1-Score, 0.836.
-Although closely followed by GPT-4V, which has an F1-Score of 0.830, our model is notable for its open-source nature, large-scale detection, and fine-grained classification of different unsafe categories.
-Particularly, it surpasses other models in identifying unsafe images from the MultiHeaded dataset, which contains 800 AI-generated images by different text-to-image models.
+We outline these datasets in Section A.5 in the Appendix.
+Effectiveness & Generalizability.
+Table 5 presents the evaluation results of the PerspectiveVision models across six datasets.
+Note that several of these evaluation datasets also serve as training data for tested baselines.
+For example, Q16 is trained on SMID, NudeNet is trained on the NSFW dataset, and the MultiHeaded classifier is trained on the MultiHeaded dataset.
+Therefore, directly comparing model performance on each individual dataset may not be fair.
+To address this, we focus on the aggregated performance across all datasets.
+Among all evaluated models, the fine-tuned LLaVA achieves the highest overall F1 score of 0.836, across six evaluation datasets.
+Notably, it also obtains the highest F1 score on AI-generated images, reaching 0.796.
 This improvement can be attributed to the inclusion of a large number of AI-generated images from UnsafeBench in its training data.
+To evaluate generalizability, we calculate the out-of-distribution (OOD) F1 score, i.e., performance on unseen datasets.
+For example, for PerspectiveVision, all datasets excluding the UnsafeBench test set are treated as unseen datasets; for Q16 and NudeNet, datasets other than SMID and NSFW are considered OOD.
+As shown in Table 5, the fine-tuned LLaVA achieves the highest OOD F1 score, indicating strong generalization capability.
+To conclude, the fine-tuned LLaVA presents the highest effectiveness and generalizability on six evaluation datasets, serving as the top-performing checkpoint in PerspectiveVision models.
 Table 6: Robust accuracy of PerspectiveVision.
@@
 MultiHeaded-D 0.966 0.912 0.886 0.884 0.912
-Robustness Evaluation.
+Robustness.
 We also test the robust accuracy of LoRA fine-tuned LLaVA against Gaussian and adversarial perturbations.
@@
 We find that LoRA fine-tuning significantly improves robustness in this image safety classification task.
-While zero-shot LLaVA shows an average RA of 0.666-0.725 (see Table 5) under four perturbations, fine-tuned LLaVA increases the average RA to 0.952 on UnsafeBench test set and 0.912 on OOD AI-generated images, in Table 7.
-Take-Aways.
+While zero-shot LLaVA shows an average RA of 0.666-0.725 (see Table 4) under four perturbations, fine-tuned LLaVA increases the average RA to 0.952 on the UnsafeBench test set and 0.912 on OOD AI-generated images, in Table 6.
+Takeaways.
 We build a comprehensive image classifier, PerspectiveVision, which identifies unsafe images across 11 categories by fine-tuning LLaVA on our UnsafeBench dataset.
@@
 For detecting sexual and pornographic images, NudeNet [14] and NSFW_Detector [13] are developed, which are trained on real-world NSFW images.
-To identify a wider range of unsafe content, Chramowski et al.
+To identify a wider range of unsafe content, Schramowski et al.
 [60] build Q16 using the prompt learning technique on a dataset containing morally negative/positive images.
@@
 Since text-to-image models like Stable Diffusion gained popularity in 2022, concerns have been raised regarding their risks of generating realistic unsafe images.
-Plenty of studies [59, 52, 70, 71, 27] focus on assessing these models’ risks.
+Plenty of studies [59, 52, 67, 68, 26, 39] focus on assessing these models’ risks.
 Schramowski et al.
@@
 However, since these classifiers are mostly trained on real-world images, it is unclear how effectively they generalize to AI-generated images.
-Our benchmarking framework, UnsafeBench, investigates their ability to generalize to AI-generated unsafe images and explores the distinct characteristics of those images.
-Conclusion & Discussion
+Our benchmarking framework, UnsafeBench, investigates their ability to generalize to AI-generated unsafe images and explores AI specific characteristics.
+Conclusion
 We establish UnsafeBench, a benchmarking framework that comprehensively evaluates the effectiveness and robustness of image safety classifiers on both real-world and AI-generated images.
 We construct a large image safety dataset of 10K manually annotated images and evaluate five conventional classifiers and three VLMs.
-Our analysis reveals the drawbacks of these classifiers and the unique challenges posed by AI-generated images to both their effectiveness and robustness.
-We therefore introduce an improved image moderation tool, PerspectiveVision, that accounts for both real-world and AI-generated images.
-Our work has the following implications for the research community and platform moderators.
-Importance of UnsafeBench Dataset.
-We provide the first large-scale image safety dataset with 11 categories of unsafe content and two sources.
-Annotated images in this dataset are equipped with three “labels”: 1) safe/unsafe, 2) specific category, and 3) sourced from the real world or AI.
-These attributes greatly improve the dataset’s utility, enabling future studies to investigate unsafe images in a multifaceted manner.
-We hope the dataset serves as a solid foundation for future image safety research.
-Challenges Posed by AI-Generated Images.
-In the era of generative AI, the influence of AI-generated content like unsafe imagery and deepfakes can be complex and multifaceted to existing models.
-Our analysis reveals that they do not only reduce the effectiveness of these models but also render them more vulnerable to adversarial attacks than real-world images.
-This is because, although current AI-generated images are realistic, there remains a distribution shift from real-world images in image quality statistics (noise level, entropy, etc.), styles, and layouts.
-For platform moderators and model developers, it is essential to include AI-generated unsafe images in the training data to mitigate these negative impacts.
-Utility of PerspectiveVision.
-PerspectiveVision is built to address the drawbacks of existing image safety classifiers, including imbalanced performance across categories, state-of-the-art model (GPT-4V) being closed-source, and degraded performances on AI-generated images.
-It is also the first attempt to build an image moderation tool that covers both the wide spectrum of unsafe content and unsafe images from different sources.
-The research community and platform moderators can directly apply it for various moderation tasks, such as reporting online unsafe images and evaluating the risks of new text-to-image models.
+Our evaluation reveals how AI-generated unsafe images pose a challenge to existing classifiers in terms of both effectiveness and robustness.
+We further introduce the image moderation tool, PerspectiveVision, to capture the distribution shift from AI images in image quality statistics (noise level, etc.), styles, and layouts.
 Limitations.
 Our work has limitations.
-First, the UnsafeBench dataset is annotated by only three authors, with the final label determined by majority vote.
-We did not rely on crowdsourcing workers for two reasons: 1) annotation requires expert knowledge in the image safety domain, such as identifying hate symbols, which cannot be guaranteed and would require specific training for crowdsourcing workers; 2) due to ethical considerations, we aim to minimize the number of annotators exposed to such unsafe content and prevent them from exposing it to third parties.
-Second, images sourced from Lexica are mostly generated by a single text-to-image model, Stable Diffusion.
+First, the UnsafeBench images (both safe and unsafe images) are collected using the same set of unsafe keywords, which makes them more challenging to classify for image safety classifiers than using irrelevant safe keywords (e.g., cats, dogs).
+However, we intentionally designed this curation process to collect challenging borderline examples, such that the classifiers are expected to distinguish between truly unsafe content and similar but acceptable content.
+Second, the UnsafeBench dataset is annotated by three experts in the research team.
+We did not rely on crowdsourcing workers for two reasons: 1) annotation requires expert knowledge in the image safety domain; and 2) due to ethical considerations, we aimed to prevent unsafe content from being exposed to third parties.
+Third, images sourced from Lexica are mostly generated by a single text-to-image model, Stable Diffusion.
 We will expand our dataset with images generated by other text-to-image models and routinely update it.
-Third, we only evaluate the generalizability of PerspectiveVision on a limited range of unsafe categories due to the lack of labeled datasets.
-We will continue expanding the evaluation with more datasets and improving the model in future work.
+Acknowledgements
+We thank all anonymous reviewers for their constructive suggestions.
+This work is partially funded by the European Health and Digital Executive Agency (HADEA) within the project “Understanding the individual host response against Hepatitis D Virus to develop a personalized approach for the management of hepatitis D” (DSolve, grant agreement number 101057917) and the BMBF with the project “Repräsentative, synthetische Gesundheitsdaten mit starken Privatsphärengarantien” (PriSyn, 16KISAO29K).
 Appendix A Appendix
 Details of Conventional Classifiers
+Table 7: Overview of conventional image safety classifiers. “D” denotes the dataset. “R” and “A” represent real-world and AI-generated images, respectively.
+Classifier Backbone Training Paradigm Training Dataset
+Q16 CLIP Prompt Learning SMID [28] (R)
+MultiHeaded CLIP Linear Probing MultiHeaded-D [53] (A)
+SD_Filter CLIP - -
+NSFW_Detector CLIP Linear Probing LAION-400M Subset (R)
+NudeNet Xception Fully Supervised NudeNet-D [15] (R)
 Q16 [60].
@@
 [^6] https://github.com/ml-research/Q16.
-It first extracts the image embedding with the CLIP [54] image encoder and calculates the cosine similarity between the image embedding and two soft prompts (optimized text embeddings), one representing the positive class and the other being the negative class.
-Q16 then assigns the image to the more relevant class.
-To obtain the optimal soft prompts, the Q16 developer starts with two normal prompts, e.g., “This image is positive/negative,” and iteratively updates their text embeddings to maximize the cosine similarity values on a labeled image dataset, SMID [29].
-The Socio-Moral Image Database (SMID) includes 2,941 morally positive and negative images, covering concepts including harm, inequality, degradation, deception, etc.
-Q16 has been widely used in evaluating the safety/appropriateness of images [60, 62, 59, 44, 73, 35, 28, 66, 27], including identifying unsafe images sourced from LAION-5B [62] and text-to-image models [59, 44, 73, 35, 48], as well as testing the reliability of safety mechanism built in text-to-image models [28, 66, 27].
+It is trained on a labeled image dataset, SMID [28], which includes 2,941 morally positive and negative images, covering concepts including harm, inequality, degradation, deception, etc.
 MultiHeaded Safety Classifier (MultiHeaded) [53].
@@
 The classifier is trained on 800 labeled images generated by text-to-image models like Stable Diffusion [53].
-It has been employed to estimate the cleanliness (percentage of unsafe images) of large public datasets [53] such as LAION-400M [63], LAION-2B [6], LAION-aesthetics [61], and YFCC15M [64].
 Stable Diffusion Filter (SD_Filter) [55].
@@
 [^8] https://huggingface.co/CompVis/stable-diffusion-safety-checker.
-It relies on the CLIP image encoder to assess image embeddings, then measures the distance between these embeddings and text embeddings of 20 sensitive concepts [23], such as “sexual,” “nude,” and “sex.” During the image generation process, the filter compares the embedding distance between the generated image with a set of pre-defined thresholds.
-If the image is too close to any sensitive concepts, the filter will predict it as unsafe and block it from users.
+It relies on the CLIP image encoder to assess image embeddings, then measures the distance between these embeddings and text embeddings of 20 sensitive concepts [22], such as “sexual,” “nude,” and “sex.
 NSFW_Detector [13].
@@
 It mainly detects sexual and nudity images.
-LAION-AI [8] uses this model to assess the NSFW score for images in their large-scale datasets including LAION-400M [63] and LAION-5B [7].
 NudeNet [14].
 NudeNet is a lightweight nudity detection tool and consists of a detector that detects the sexually explicit areas in images and a classifier that directly classifies them as explicit images or not.
-In this study, we use the binary NudeNet classification model.
+We use the binary NudeNet classification model.
 [^10] https://pypi.org/project/nudenet/.
-It is a classifier with Xception as the backbone, trained on roughly 160K automatically-labeled images [16].
-It has been extensively applied in detecting sexually explicit images  [59, 68, 48, 35, 73] and is often combined with Q16 for identifying general unsafe content [59, 35, 48, 73].
+Details of the RoBERTa Classifier
+Table 8: Reliability of the RoBERTa classifier.
+VLM Accuracy Precision Recall F1-Score
+LLaVA 1.000 1.000 1.000 1.000
+InstructBLIP 0.980 0.981 0.980 0.978
+GPT-4V 0.995 0.995 0.995 0.995
+Overall 0.992 0.992 0.992 0.991
+To convert VLM-generated responses into standard classes, we fine-tune a language model, RoBERTa [46], to categorize LLM outputs into one of three classes: safe, unsafe, or uncertain.
+Since most responses already contain indicative class terms, only a few consist of lengthy sentences.
+To this end, we annotate a small set of 10–20 unique responses per class for each VLM and train the classifier for 10 epochs.
+For example, for the safe class, we use outputs such as “safe,” “the image is safe,” “the image is not considered unsafe as…” as representative examples and fine-tune RoBERTa in a supervised manner.
+To validate the reliability of this classifier, we randomly sample 600 VLM-generated responses (200 from each VLM) and manually classify them into safe, unsafe, and uncertain classes.
+We then compare this ground truth with the predictions from the RoBERTa classifier and compute the metrics shown in Table 8.
+The evaluation confirms the high reliability of the classifier.
 Details of Large VLMs
@@
 It comprehends the image with the CLIP image encoder and understands the user’s prompts with a large language model, Vicuna [23].
-LLaVA is trained on the LAION-CC-SBU dataset and instruction dataset generated by GPT4V [45].
 We use the llava-v1.5-7b checkpoint.
@@
 Adversarial Examples
-Adversarial examples are an important tool to evaluate a classifier’s robustness.
-We introduce three widely used gradient-based methods to create adversarial examples, FGSM [37], PGD [47], and DeepFool [50].
 FGSM (Fast Gradient Sign Method) [36].
@@
 To obtain the adversarial example at time $t+1$, it first applies perturbations on the adversarial examples at time $t$ with the step size of $\alpha$.
-It then projects them back to the $\epsilon$-ball space (to constrain the perturbations).
+It then projects them back to the $\epsilon$-ball space.
 $x^{t+1}=\Pi_{x+S}\left(x^{t}+\alpha\cdot\text{sign}(\nabla_{x}L(\theta,x^{t},y))\right).$ (4)
@@
 Like the Self-Harm Dataset, this dataset was also obtained from Roboflow.
-Experimental Setup
-Model Architectures.
-We utilize the pre-trained CLIP (ViT-L-14) model to extract image features.
-For linear probing, we use an MLP of three layers, containing 768-384-2 neurons, respectively.
-For prompt learning, we set the length of soft prompts to eight, indicating that each soft prompt consists of eight vectors.
-We start with two initial prompts, i.e., “This image is about something safe” and “This image is about something unsafe in [category],” where “[category]” will be replaced with the actual category name.
-Implementation Details. We train all the models on the UnsafeBench training set. We use Pytorch and NVIDIA A100 GPUs for all experiments.
-  • Linear Probing.
-We independently train 11 MLPs, each designated to identify one of the 11 unsafe categories.
-We load the entire set of training images to train each MLP.
-In this case, images classified as unsafe in other categories are still treated as safe.
-This ensures that each MLP focuses exclusively on one unsafe category.
-During training, we use the standard cross-entropy loss and the Adam optimizer with a learning rate of 2e-4.
-The batch size is set to 128, and we train each MLP for 30 epochs.
-  • Prompt Learning.
-Similarly, we optimize 11 sets of soft prompts using the same data loading method as in linear probing.
-We still use the CrossEntropy loss and the Adam optimizer with a learning rate of 2e-4.
-We train each set of soft prompts for 80 epochs with a batch size of 128.
-  • LoRA Fine-Tuning.
-We build the instruction dataset using the prompt template shown in Figure 10.
-Specifically, we construct a pair of instructional data points for each unsafe image, which we call the positive example and the negative example.
-In the positive example, given the unsafe image, we first identify the unsafe category it violates and then randomly append K (from 1 to 10) other unsafe categories to create a customized image safety taxonomy.
-The safety label is still recorded as unsafe.
-Meanwhile, in the negative example, we randomly select K unsafe categories that the given unsafe image does not violate, and then we record the safety label as safe.
-This approach ensures that VLMs learn the safety of images rather than simply memorizing the output format.
-We then fine-tune LLaVA on this dataset with its original training loss (CrossEntropy loss) and optimizer (AdamW), following the setup in [45].
-The batch size is 16, and the learning rate is 2e-4.
-We set the training epoch to 1.
-Evaluation Setup.
-We set the customized scope of unsafe images to all 11 unsafe categories.
-This means that if the given image violates any of these 11 unsafe categories, it will be classified as unsafe.
-The F1-Score is calculated based on binary predictions, safe/unsafe, without considering the specific unsafe category.
 Table 9: Unsafe image taxonomy. In the definition of each unsafe category, we denote the OpenAI content policy in bold and append the specific examples generated by Vicuna-33b.
@@
 11 Spam Unsolicited bulk content. For example, Viagra, Cialis, online pharmacy, Levitra, online casino, poker, buy Tramadol, Xanax, Ambien, pop-up ads, etc.
+Experimental Setup
+Model Architectures.
+We utilize the pre-trained CLIP (ViT-L-14) model to extract image features.
+For linear probing, we use an MLP of three layers, containing 768-384-2 neurons, respectively.
+For prompt learning, we set the length of soft prompts to eight, indicating that each soft prompt consists of eight vectors.
+We start with two initial prompts, i.e., “This image is about something safe” and “This image is about something unsafe in [category],” where “[category]” will be replaced with the actual category name.
+(a) The Sexual Category
+(b) The Violence Category
+Figure 10: Optimal K for the Sexual and Violence category.
+Implementation Details. We train all the models on the UnsafeBench training set. We use Pytorch and NVIDIA A100 GPUs for all experiments.
+  • Linear Probing.
+We independently train 11 MLPs, each designated to identify one of the 11 unsafe categories.
+We load the entire set of training images to train each MLP.
+In this case, images classified as unsafe in other categories are still treated as safe.
+This ensures that each MLP focuses exclusively on one unsafe category.
+During training, we use the standard cross-entropy loss and the Adam optimizer with a learning rate of 2e-4.
+The batch size is set to 128, and we train each MLP for 30 epochs.
+  • Prompt Learning.
+Similarly, we optimize 11 sets of soft prompts using the same data loading method as in linear probing.
+We still use the CrossEntropy loss and the Adam optimizer with a learning rate of 2e-4.
+We train each set of soft prompts for 80 epochs with a batch size of 128.
+  • LoRA Fine-Tuning.
+We build the instruction dataset using the prompt template shown in Figure 9.
+Specifically, we construct a pair of instructional data points for each unsafe image, which we call the positive example and the negative example.
+We then fine-tune LLaVA on this dataset with its original training loss (CrossEntropy loss) and optimizer (AdamW), following the setup in [45].
+The batch size is 16, and the learning rate is 2e-4.
+We set the training epoch to 1.
+Evaluation Setup.
+We set the customized scope of unsafe images to all 11 unsafe categories.
+This means that if the given image violates any of these 11 unsafe categories, it will be classified as unsafe.
+The F1-Score is calculated based on binary predictions, safe/unsafe, without considering the specific unsafe category.
+Figure 11: Image quality statistics of real-world and AI-generated images. We report signal-to-noise ratio (SNR), edge density, and entropy.
 Table 10: Statistics of the UnsafeBench dataset across 11 unsafe categories and two sources.
@@
 Figure 13: Neo-Nazi and anti-Semitic symbols that evade detection of GPT-4V and Q16.
-Figure 13: Image quality statistics of real-world and AI-generated images. We report signal-to-noise ratio (SNR), edge density, and entropy.
 (a) False Negatives (Misclassify Unsafe as Safe)
@@
 Figure 14: Image clusters from the Violence category that are misclassified by Q16 and GPT-4V. We annotate each central image with its cluster ID and cluster size.
+Figure 15: Confidence scores of classifiers’ predictions over different groups of images.
+The median value is denoted next to the dashed line.
+AI-generated images tend to have lower confidence scores than real-world images.
 Figure 16: Loss and loss change for four conventional classifiers on original images and their adversarial examples (FGSM).
 Classifiers tend to show higher loss values and greater loss changes with AI-generated images than with real-world images.
+Table 12: F1-Score of PerspectiveVision models and baselines on the UnsafeBench test set.
+Type Model Hate Harassment Violence Self-harm Sexual Shocking Illegal Activity Deception Political Health Spam Overall
+PerspectiveVision Linear Probing (CLIP) 0.776 0.736 0.767 0.586 0.924 0.894 0.877 0.841 0.907 0.903 0.899 0.859
+Prompt Learning (CLIP) 0.437 0.557 0.679 0.337 0.890 0.758 0.765 0.596 0.829 0.635 0.654 0.687
+LoRA Fine-tuning (LLaVA) 0.686 0.755 0.760 0.743 0.913 0.861 0.889 0.780 0.878 0.909 0.855 0.844
+Classifier Ensemble Q16_NudeNet 0.543 0.454 0.619 0.368 0.704 0.735 0.591 0.679 0.515 0.538 0.219 0.585
+Q16_NSFW_Detector 0.571 0.503 0.646 0.400 0.761 0.765 0.576 0.693 0.504 0.500 0.147 0.606
+Q16_SD_Filter 0.473 0.430 0.610 0.396 0.849 0.710 0.566 0.648 0.507 0.449 0.205 0.595
+Q16_MultiHeaded 0.550 0.495 0.626 0.421 0.788 0.773 0.580 0.660 0.713 0.535 0.216 0.635
+Q16_MultiHeaded_NudeNet 0.524 0.446 0.608 0.359 0.826 0.738 0.588 0.649 0.730 0.566 0.247 0.625
+All Conventional Classifiers 0.444 0.469 0.589 0.331 0.861 0.700 0.571 0.621 0.723 0.512 0.260 0.607
+Commercial Google’s SafeSearch 0.095 0.423 0.519 0.308 0.826 0.623 0.307 0.200 0.589 0.442 0.279 0.515
+Microsoft’s Filter 0.341 0.310 0.642 0.304 0.854 0.435 0.298 0.172 0.125 0.299 0.127 0.475
+GPT-4V 0.439 0.626 0.747 0.500 0.886 0.819 0.848 0.636 0.780 0.500 0.549 0.717